### v1

# Libra: Leveraging Temporal Images for Biomedical Radiology Analysis

Abstract
Radiology report generation (RRG) is a challenging task, as it requires a thorough understanding of medical images, integration of multiple temporal inputs, and accurate report generation. Effective interpretation of medical images, such as chest X-rays (CXRs), demands sophisticated visual-language reasoning to map visual findings to structured reports. Recent studies have shown that multimodal large language models (MLLMs) can acquire multimodal capabilities by aligning with pre-trained vision encoders. However, current approaches predominantly focus on single-image analysis or utilise rule-based symbolic processing to handle multiple images, thereby overlooking the essential temporal information derived from comparing current images with prior ones. To overcome this critical limitation, we introduce Libra, a temporal-aware MLLM tailored for CXR report generation using temporal images. Libra integrates a radiology-specific image encoder with a MLLM and utilises a novel Temporal Alignment Connector to capture and synthesise temporal information of images across different time points with unprecedented precision. Extensive experiments show that Libra achieves new state-of-the-art performance among the same parameter scale MLLMs for RRG tasks on the MIMIC-CXR. Specifically, Libra improves the RadCliQ metric by 12.9% and makes substantial gains across all lexical metrics compared to previous models111Our code is open-sourced at https://github.com/X-iZhang/Libra..
1 Introduction
Radiology reports are a key component of biomedical radiology analysis, designed to enhance clarity and efficiency in medical communication. These structured reports are typically divided into distinct sections, such as Findings, Impression, Indication, Technique, Comparison, and History (Ganeshan et al., 2018). Serving as the primary medium for radiologists to convey their findings and conclusions from imaging studies like chest X-rays (CXRs), radiology reports play a crucial role in guiding diagnostic and therapeutic decisions for various diseases (Najjar, 2023). However, generating these reports manually is a complex and time-consuming task. Automating radiology report generation (RRG) presents a valuable opportunity to boost radiologist productivity, improve communication, and alleviate burnout (Zhang et al., 2020b). Despite its potential, RRG is a challenging task due to the intricate nature of medical imaging and the need for precise and accurate documentation.
Although the advancements in paradigm models like LLaVA (Liu et al., 2023) and InstructBLIP (Dai et al., 2023), they still exhibit a significant performance gap when applied to RRG tasks. In particular, their effectiveness diminishes in biomedical contexts due to the substantial differences between biomedical and general image-text pairs (Tu et al., 2023; Saab et al., 2024). Consequently, these models often perform at a surface level, akin to a layperson-level understanding, when dealing with specialised medical imaging tasks. Recent advancements have explored continued pre-training of general-purpose foundation models for medical tasks, but these models still fall short in addressing the complexities of medical image analysis (Li et al., 2023a; Chaves et al., 2024; Park et al., 2024).
Harnessing the full potential of multimodal large language models (MLLMs) for understanding and reasoning over multimodal data in the biomedical domain poses a significant challenge, primarily due to the granularity and specificity required in medical image analysis (Wang et al., 2023a). Preliminary attempts to apply MLLMs to RRG have largely concentrated on generating reports from a single image, often neglecting the crucial temporal relationships between images taken at different time points (Zhang et al., 2024b). The MIMIC-CXR Database (Johnson et al., 2019a) reveals that 67% of patients underwent at least two studies at different time intervals, underscoring the necessity of incorporating temporal context in clinical practice. This highlights a critical gap in existing models, which are not equipped to handle the dynamic progression of medical conditions reflected in sequential imaging studies.
In standard clinical practice, radiologists rely heavily on comparing current imaging results with previous studies to assess temporal changes. However, many models inadvertently introduce spurious references to prior examinations that do not exist during inference, leading to inaccurate reports. Recent advancements222Detailed related work is discussed in Appx. A.1, and our research objectives are explained in Appx. A.2., such as MedVersa (Zhou et al., 2024) and MAIRA-2 (Bannur et al., 2024), can process multiple images simultaneously. These models insert visual tokens from different images at specific points within the textual input, hinging on the LLM to interpret reference information. Additionally, LLaVA-Rad (Chaves et al., 2024) utilised GPT-4V (OpenAI et al., 2024) for radiology report extraction and augmentation, eliminating hallucinations associated with prior image references in the dataset. Despite these efforts, existing models still lack integrated mechanisms for perceiving temporal information within their architectures, which restricts the model’s ability to achieve the temporal awareness required for comprehensive reporting. Meanwhile, existing MLLMs typically utilise visual embeddings from the last or penultimate layer of the image encoder (Chen et al., 2023a; Zhang et al., 2024a), which can only capture the global characteristics of the image. However, RRG tasks require not only identifying findings but also capturing subtle details such as severity, extent, and the progression of findings (Sloan et al., 2024). Such fine-grained details are difficult to represent using embeddings from any single layer of the image encoder (Jiang et al., 2024). Therefore, for the high-granularity demands of RRG tasks, relying solely on representation from a single layer limits the quality of report generation. To tackle these limitations, we aim to enhance the temporal awareness of MLLMs for RRG task by tackling two main challenges:
It is non-trivial to design effective structures in MLLMs to handle prior study citations across various time points in the RRG task.
The scarcity of effective connectors in MLLMs capable of handling the high-granularity requirements of downstream tasks.
To overcome these challenges, we propose Libra (Leveraging Temporal Images for Biomedical Radiology Analysis), a novel framework designed to incorporate temporal change information into the RRG task. Libra employs a pre-trained visual transformer, RAD-DINO (Pérez-García et al., 2024), as the image encoder to generate robust image features, which are then refined using a new adapter specifically designed for the temporal awareness, before being fed into the medical large language model (LLM), Meditron (Chen et al., 2023b). Through a two-stage training strategy, Libra demonstrates the potential of MLLMs for specialised radiology applications. Our modular approach integrates state-of-the-art open-source pre-trained models for image and text while using a temporal-aware adapter to align these modalities with the text embedding space. Extensive experiments show that Libra substantially outperforms existing MLLMs in RRG tasks, demonstrating superior cross-modal understanding and reasoning capabilities. Specifically, our contributions include:
We present Libra, the temporal-aware MLLM capable of capturing temporal changes and overcoming the challenge of handling prior study citations, setting a new state-of-the-art performance in RRG tasks on the MIMIC-CXR dataset among MLLMs of the same parameter scale.
We designed the Temporal Alignment Connector (TAC) with two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image feature embedding from the encoder, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies. This enhances Libra’s ability to capture and utilise temporal information in the RRG task.
2 Libra: Being Aware of Temporal Changes in Chest X-rays Report Generation
2.1 Model architecture
Libra model follows the common architecture of MLLMs, such as LLaVA, which includes an image encoder, a text decoder (i.e., an LLM pre-trained based on text-only data) and an adapter module for mapping visual representations into text space. In this work, we utilise a frozen biomedical image encoder, RAD-DINO (Pérez-García et al., 2024), a visual transformer extensively pre-trained on medical scans using the DINOv2 image-only self-supervised learning approach (Oquab et al., 2024). The LLM is deployed by Meditron-7B (Chen et al., 2023b), which builds on Llama-2 (Touvron et al., 2023) and further pre-trained on a specialised medical corpus. However, for the adapter module, we specifically designed a Temporal Alignment Connector (TAC) to capture and integrate temporal information across two images from different time points, as shown in Figure 1. The TAC bridges the image encoder and LLM. It consists of two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image representations, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies. This design enables Libra to effectively manage temporal data and enhances its ability to generate accurate and coherent radiology reports.
2.2 Temporal Alignment Connector (TAC): New Adapter Design
To address the challenges of integrating temporal information and aligning high-granularity image features in RRG tasks, we developed the TAC module. It bridges the gap between visual features extracted from multiple temporal snapshots of a patient’s CXRs and the language model. The TAC consists of two sub-modules: the LFE, which is to extract the layerwise representation of images, and the TFM, which is to capture the temporal-aware representation between two images.
2.2.1 Layerwise Feature Extractor (LFE)
To leverage abundant feature information in a pre-trained image encoder, we extract all image patch token features from each hidden layer for a given input image. Specifically, RAD-DINO has 12 hidden layers and processes 518 × 518 images into 14 × 14 patches, generating 1,369 patch token sequences per hidden layer for each image. We do not use the token. Following this, we obtain patch embeddings of the same dimension from each layer, denoted as , where is the number of patch tokens and is the embedding dimension of the image encoder. Then, we combine them as , where is the number of hidden layers. Drawing from VGG (Simonyan & Zisserman, 2015), we employ a progressive compression strategy to reduce the layer dimension, ensuring that the LFE captures the most relevant features. Initially, we utilise Squeeze-and-Excitation (SE) Networks (Hu et al., 2019), which construct informative features by integrating both spatial and channel-wise information within local receptive fields at each layer. The SE block is applied to obtain calibrated feature representations, using GELU (Hendrycks & Gimpel, 2023) as the activation function. Next, we employ a specialised pointwise convolution module to align the feature spaces across different layers, using a depthwise 2D convolution with filters and stride of , without bias. The dimensions are gradually reduced to better align with the text in the LLM. The compressed features are represented as , where is the original layer number and is the layer number after compression. Following the size-reduction pattern of convolutional layers in VGG, the image representations are compressed according to 333Since RAD-DINO has 12 hidden layers, the prime factorisation chain provides the factors as .. Through three stages of progressive compression, we obtain the final representation of image patch features. We define the function to project image features into the same dimension:
where , . The LFE independently processes each image to produce a unified representation for temporal alignment, as (a) in Figure 1.
2.2.2 Temporal Fusion Module (TFM)
The fusion module is inspired by the transformer decoder. It utilises prior images as auxiliary information to weight the current image, generating a refined, temporal-aware representation suitable for the LLM’s latent space. The TFM uses an attention mechanism to dynamically learn relationships between image pairs, enabling the model to adapt to various temporal changes.
Prior Image Prefix Bias: The dataset contains samples with and without a prior image. The patch token features of the current and previous images processed by the LFE are denoted as and . If no true prior image exists, we set , effectively using the current image as a dummy prior. To differentiate this case, we introduce a trainable bias . Following the attention mechanism’s scaling techniques for adjusting hidden space degrees of freedom with a chi-square distribution (Vaswani, 2017; Wang et al., 2018a), we apply an exponent of to the cosine similarity, where is the hidden dimension of the LLM, to obtain a weighted bias addition to :
This nonlinear scaling amplifies higher similarity values, effectively modulating the influence of prior image features. When no true prior image is available, the high similarity score ensures that the effect of the dummy prior is adequately represented. This adjustment prevents samples with a dummy prior image from undergoing redundant rounds of parallel multi-head self-attention during subsequent propagation through the transformer blocks, as shown in Figure 1.
Transformer Block: Our transformer block includes a multi-head cross-attention sub-layer, multi-head self-attention sub-layers, and two multi-layer perceptron (MLP) sub-layers. As shown in Figure 1 (b). The paired are processed with layer normalization and residual connections:
where is a simple neural network composed of two fully connected layers with GELU as the activation function. After that, the features are processed through , a straightforward neural network consisting of four fully connected layers with the same activation function, but with hidden dimensions matching those of the LLM. We define our fusion module function as .
Subsequently, becomes the input sequence that LLM can comprehend, where is the number of patch tokens and is the hidden dimension of LLM. The final output is a refined feature representation that encapsulates the temporal evolution of the patient’s condition, making it suitable for the language model to generate accurate and contextually aware radiology reports.
2.3 Prompt Design
To enhance Libra’s ability to perceive temporal changes and integrate medical information in the RRG task, we designed a comprehensive prompting strategy comprising a system prompt and detailed clinical prompts, as shown in Figure 1. The system prompt is crafted to enable the LLM to recognise temporal variations: “The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes.” Following this, additional sections of the report, such as {Indication}, {History}, {Comparison}, and {Technique}, are incorporated into the clinical instructions (e.g., in Appx. B.4). The clinical prompt to Libra is: “Provide a detailed description of the findings in the radiology image. Following clinical context: .” If no clinical instructions are available, the prompt defaults to: “Provide a detailed description of the findings in the radiology image.” After tokenizing and embedding the prompt and the answer, we insert the image patch tokens at the specified location, typically between the system prompt and the clinical prompts.
2.4 Temporal-aware Training
In this study, we focus on frontal-view images, either posterior-anterior (PA) or anterior-posterior (AP), and the Findings sections of radiology reports, as they contain the most direct clinical observations. We adopt a two-stage training strategy inspired by recent advancements in MLLM fine-tuning techniques (McKinzie et al., 2024). This proposed approach aims to progressively teach the model essential skills, including visual feature alignment and temporal information extraction. Libra’s training process involves two stages: temporal feature alignment and downstream task fine-tuning.
In the first stage, the visual encoder and LLM weights are frozen, and the TAC is trained with Findings and Impression generation, along with CXR-related VQA tasks to extract high-quality image representations and capture temporal changes. In the second stage, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM on the Findings section generation task, keeping the visual encoder and connector weights frozen. Unlike traditional full fine-tuning, LoRA achieves comparable performance with significantly lower training costs. The detailed training configuration, including learning rate schedules and model parameters, is provided in Appx. B.1.
3 Experiments
3.1 Task and Dataset
Task Description We focus on generating the Findings section of radiology reports for a set of frontal CXRs to ensure fair comparison with previous work. The Findings section contains descriptions by radiologists of both normal and abnormal findings. Typically, these CXRs are accompanied by additional sections such as Indication and Technique, which provide the rationale for the study, including clinical history or specific requests from the referring physician. While not directly related to diagnostic interpretation, these sections serve as routine records and assist the model in better understanding temporal changes between images across different periods. Therefore, we incorporate clinical instructions about the current image as prompts to guide Libra to complete the RRG task.
The most common CXR is frontal views, either PA or AP. While other views, such as lateral images, are occasionally used for diagnostic purposes, they mainly serve as supplementary tools to aid in interpreting anatomical structures more comprehensively (Islam et al., 2023). Therefore, the frontal view remains the standard perspective for clinical interpretation of CXRs, and our study is consistent with prior work in RRG tasks, including studies by Chaves et al. (2024) and Hyland et al. (2024). For both the current and prior images, we exclusively use a single frontal view.
Dataset Description To train Libra, we utilise the MIMIC-CXR dataset (Johnson et al., 2019b) and its derivative datasets, Medical-Diff-VQA (Hu et al., 2023) and MIMIC-Ext-MIMIC-CXR-VQA (Bae et al., 2023), as shown in the Table 1. All datasets are split according to the official labels to ensure no data leakage. Detailed descriptions of datasets can be found in the Appx. B.2.
| Data source | Task Type | # Samples | % Has Prior | ||||
| Train (%) | Valid (%) | Test (%) | Train | Valid | Test | ||
| MIMIC-CXR | Findings | 162 955 (13.43%) | 1286 (0.88%) | 2461 (2.78%) | 58.43 | 60.11 | 86.03 |
| Impression | 199 548 (16.45%) | 1671 (1.14%) | 2343 (2.64%) | 64.85 | 67.09 | 85.49 | |
| Medical-Diff-VQA | Difference | 131 563 (10.85%) | 16 372 (11.17%) | 16 389 (18.48%) | 100 | 100 | 100 |
| Abnormality | 116 394 (9.59%) | 14 512 (9.90%) | 14 515 (16.37%) | 100 | 100 | 100 | |
| Presence | 124 654 (10.28%) | 15 549 (10.61%) | 15 523 (17.51%) | 100 | 100 | 100 | |
| View | 44 970 (3.71%) | 5696 (3.89%) | 5599 (6.31%) | 100 | 100 | 100 | |
| Location | 67 187 (5.54%) | 8510 (5.81%) | 8496 (9.58%) | 100 | 100 | 100 | |
| Level | 53 728 (4.43%) | 6722 (4.59%) | 6846 (7.72%) | 100 | 100 | 100 | |
| Type | 22 067 (1.82%) | 2709 (1.85%) | 2702 (3.05%) | 100 | 100 | 100 | |
| MIMIC-Ext-MIMIC | Presence | 109 455 (9.02%) | 26 153 (17.84%) | 4566 (5.15%) | 0 | 0 | 0 |
| -CXR-VQA | Anatomy | 37 952 (3.13%) | 10 210 (6.96%) | 1963 (2.21%) | 0 | 0 | 0 |
| Attribute | 49 948 (4.12%) | 13 111 (8.94%) | 2578 (2.91%) | 0 | 0 | 0 | |
| Abnormality | 60 692 (5.00%) | 16 109 (10.99%) | 3199 (3.61%) | 0 | 0 | 0 | |
| Size | 16 000 (1.32%) | 4000 (2.73%) | 705 (0.80%) | 0 | 0 | 0 | |
| Plane | 7992 (0.66%) | 1992 (1.36%) | 386 (0.44%) | 0 | 0 | 0 | |
| Gender | 7992 (0.66%) | 1992 (1.36%) | 396 (0.45%) | 0 | 0 | 0 | |
| Total | Multi-type | 1 213 097 (100%) | 146 594 (100%) | 88 669 (100%) | 64.73 | 49.09 | 83.67 |
Inspired by the multi-stage training techniques commonly employed in recent MLLMs fine-tuning methods (Zhu et al., 2023), we adopt a two-stage training strategy to infuse multimodal capabilities into the model. In the first stage, we leverage a large-scale dataset to train the adapter, primarily focusing on learning the distribution of CXR token embeddings and the general relationships between CXR and text tokens across the MIMIC dataset. The second stage fine-tuning the model on downstream tasks, where the text decoder (i.e., the LLM) is refined to focus on aligning high-granularity information between CXRs and the Findings section of radiology reports.
In addition to Findings section generation, the temporal feature alignment stage also uses the Impression section and VQA tasks. The Impression section summarises diagnoses and proposes further investigations (Babar et al., 2021), aiding the alignment between CXRs and their textual descriptions. We use the same system and clinical prompts as for Findings, replacing ‘Findings’ with ‘Impression’. For VQA, system prompts remain the same, while clinical prompts adapt to medical-specific questions, guiding the model to generate corresponding captions. These fine-grained VQA tasks expand the MLLM’s biomedical vocabulary and improve the alignment of image-text annotations.
3.2 Evaluation Metrics
We evaluate the generated reports using lexical and radiology-specific metrics, following established protocols.. For lexical metrics, we use ROUGE-L (Lin, 2004), BLEU- (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005) and BERT (Devlin et al., 2019). For radiology-specific metrics, we use RadGraph-F1 (Jain et al., 2021), RG (Delbrouck et al., 2022a), F1-CheXpert (Irvin et al., 2019a), CheXbert vector similarity (Yu et al., 2022), as well as RadCliQ version Yu et al. (2022).
Additionally, we introduce a novel temporal entity F1 score (called ) to assess the model’s ability to capture temporal information. Clinical metrics typically emphasise the accuracy of medical findings, prioritising the detection of clinically relevant entities. In contrast, the temporal entity score specifically measures the accuracy of entities related to progression over time described in the report. Detailed descriptions of all metrics and an illustrative analysis are provided in Appx. B.3.
Temporal Entity F1 Following the work of Bannur et al. (2023), we set a reward list that includes common radiology-related keywords associated with temporal changes. Temporal entities are then extracted from both the ground truth and the generated reports . Importantly, no stemming or lemmatization is applied during token processing to preserve the precise description of temporal changes. After extraction, we compute precision and recall , which are subsequently used to calculate the temporal entity score, defined as the harmonic mean of precision and recall (Van Rijsbergen, 1974), also known as the F1 score.
where is a small value, set to a default of , to prevent division by zero (it is also added to the numerator for special cases where no temporal entities are present in the ground truth). The coefficient controls the balance between precision and recall, with a default value of .
3.3 Main Result
Although MIMIC-CXR provides an “official” test split, strict comparisons with prior studies remain challenging due to differences in test set inclusion criteria and pre-processing steps. For instance, Yu et al. (2022) and Jeong et al. (2023) included only one image per study, resulting in a test set of 1,597 samples, while Tanida et al. (2023) followed the Chest ImaGenome split (Wu et al., 2021b). Such variations in test set distributions can significantly impact the reported results (Park et al., 2024). To ensure fairness, we use a widely adopted test set focused on frontal-view CXRs, aligned with previous studies such as MAIRA-1 (Hyland et al., 2024) and LLaVA-Rad (Chaves et al., 2024) 444The test set consists of 2,461 frontal-view image-report samples..
Additionally, the latest concurrent work M4CXR (Park et al., 2024) employs multi-turn chain-of-thought (Wei et al., 2023) prompting to generate reports, which differs from our task setup. Furthermore, we do not compare with the recent work MAIRA-2, as it is a MLLM specifically designed for grounded radiology report generation task, which incorporates lateral views and prior study reports for each subject within the input prompt. Bannur et al. (2024) emphasises a positive transfer between this task and general report generation, which is also beyond the scope of our study 555For comparisons with the latest concurrent and non-LLM-based models can be found in Appx. D..
Taking these considerations into account, we compared our model with state-of-the-art models, including LLaVA-Med (Li et al., 2023a), CheXagent (Chen et al., 2024), GPT-4V (OpenAI et al., 2024), Med-PaLM (Tu et al., 2023), LLaVA-Rad and MAIRA-1. The results are shown in Table 2. Since many of these models are not publicly available, we present their evaluation results as reported in the original sources.
| Metric | LLaVA-Med | CheXagent | GPT-4V | Med-PaLM | LLaVA-Rad | MAIRA-1 | Libra(%) |
| Lexical: | |||||||
| ROUGE-L | 27.6 | 21.5 | 13.2 | 27.5 | 30.6 | 28.9 | 36.7 (19.9%) |
| BLEU-1 | 35.4 | 16.9 | 16.4 | 32.3 | 38.1 | 39.2 | 51.3 (30.9%) |
| BLEU-4 | 14.9 | 4.7 | 17.8 | 11.5 | 15.4 | 14.2 | 24.5 (37.6%) |
| METEOR | 35.3 | – | – | – | – | 33.3 | 48.9 (38.5%) |
| Clinical: | |||||||
| RadGraph-F1 | 19.1 | – | – | 26.7 | – | 24.3 | 32.9 (23.2%) |
| RG | 23.8 | 20.5 | 13.2 | – | 29.4 | 29.6 | 37.6 (27.0%) |
| RadCliQ() | 3.3 | – | – | – | – | 3.1 | 2.7 (12.9%) |
| CheXbert vector | 36.9 | – | – | – | – | 44.0 | 46.9 (6.59%) |
| CheXpert-F1: | |||||||
| Micro-F1-14 | 42.7 | 39.3 | 35.5 | 53.6 | 57.3 | 55.7 | 55.9 (-2.4%) |
| Macro-F1-14 | 26.9 | 24.7 | 20.4 | 39.8 | 39.5 | 38.6 | 40.4 (1.5%) |
| Micro-F1-5 | 43.9 | 41.2 | 25.8 | 57.9 | 57.4 | 56.0 | 60.1 (3.8%) |
| Macro-F1-5 | 36.3 | 34.5 | 19.6 | 51.6 | 47.7 | 47.7 | 53.8 (4.3%) |
From Table 2 we can see that, for all traditional lexical metrics, Libra substantially outperforms the baseline models, achieving scores of 36.7 for ROUGE-L, 51.3 and 24.5 for BLEU-1 and BLEU-4 respectively, and 48.9 for METEOR. In clinical metrics, Libra also showed outstanding performance, substantially improving in RadGraph-F1 and its variant RG, with scores of 32.9 and 37.6, respectively. Notably, Libra achieved the highest score (i.e. 2.7) on the radiologist-aligned RadCliQ metric. For CheXbert-derived metrics, it also secured the top position with a vector embedding matrix score of 46.9. In the CheXpert 5-class subset, Libra excelled with Micro-F1 and Macro-F1 scores of 60.1 and 53.8, respectively. In the CheXpert 14-class subset, Libra achieved the highest Macro-F1 score of 40.4, while its Micro-F1 score of 55.9 was only slightly below the leading model.
In summary, Libra demonstrated outstanding performance across both sets of metrics, with only minor gaps in one specific measure. The remarkable capabilities of Libra can be attributed to its innovative Temporal Alignment Connector which effectively captures temporal changes in the dataset, enabling the generation of highly accurate and clinically relevant reports.
4 Ablation Study
To thoroughly assess the impact of various optimisation components, we conducted a series of ablation experiments, evaluating different modules and dataset expansions within Libra. All ablation models were tested on the MIMIC-CXR official test split for the Findings section generation task, using identical hyperparameter settings during training and inference.
Q1: Does incorporating temporal information bring positive effects to Libra in RRG tasks?
| Metric | Libra | |
| w/ true | w/ dummy(%) | |
| Lexical: | ||
| ROUGE-L | 36.66 | 36.17 (-1.34%) |
| BLEU-1 | 51.25 | 51.20 (-0.10%) |
| BLEU-4 | 24.54 | 24.33 (-0.86%) |
| METEOR | 48.90 | 48.69 (-0.43%) |
| BERTScore | 62.50 | 61.94 (-0.90%) |
| F1 | 35.34 | 32.72 (-7.41%) |
| Clinical: | ||
| RadGraph-F1 | 32.87 | 32.42 (-1.37%) |
| RG | 37.57 | 36.92 (-1.73%) |
| RadCliQ() | 2.72 | 2.76 (-1.47%) |
| CheXbert vector | 46.85 | 46.31 (-1.15%) |
| CheXpert-F1: | ||
| Micro-F1-14 | 55.87 | 55.25 (-1.11%) |
| Macro-F1-14 | 40.38 | 40.15 (-0.57%) |
| Micro-F1-5 | 60.07 | 58.93 (-1.90%) |
| Macro-F1-5 | 53.75 | 52.61 (-2.12%) |
Temporal information is embedded in paired images and referenced in the corresponding radiology reports, reflecting changes over time through references to previous symptoms and descriptions of their progression. As indicated in Table 1, 86% of the test data includes true prior images. To assess whether Libra can effectively perceive and utilise temporal information during inference, we conducted separate evaluations using either true prior images or only dummy prior images. When true prior images were unavailable, the model treated the current image as a dummy prior image. As shown in Table 3, Libra demonstrated substantial enhancements across all metrics when true prior images were used as references. Specifically, the improvement in clinical scores exceeds that of lexical scores, indicating that temporal information is crucial for generating high-quality medical reports, beyond merely improving linguistic fluency. Notably, the F1 score shows the most significant impact, with a difference of 7.41%. This highlights Libra’s capability to effectively leverage temporal changes provided by true prior images, thereby enhancing the quality of the generated Findings section and overall performance in the RRG task.
Q2: Does the Temporal Alignment Connector effectively improve model performance?
| Metric | Libra-1 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
| Lexical: | |||||
| ROUGE-L | 27.56 | 27.33 (-0.85%) | 27.21 (-1.27%) | 27.43 (-0.48%) | 26.17 (-5.04%) |
| BLEU-1 | 34.84 | 34.17 (-1.92%) | 34.21 (-1.82%) | 34.60 (-0.67%) | 33.03 (-5.20%) |
| BLEU-4 | 11.51 | 11.13 (-3.33%) | 11.11 (-3.47%) | 11.43 (-0.73%) | 10.02 (-12.98%) |
| METEOR | 35.50 | 35.06 (-1.24%) | 34.96 (-1.52%) | 35.28 (-0.62%) | 33.98 (-4.28%) |
| BERTScore | 55.87 | 55.60 (-0.49%) | 55.49 (-0.69%) | 55.74 (-0.23%) | 54.63 (-2.22%) |
| F1 | 26.63 | 25.96 (-2.51%) | 26.21 (-1.57%) | 26.58 (-0.18%) | 25.39 (-4.65%) |
| Clinical: | |||||
| RadGraph-F1 | 22.52 | 22.20 (-1.42%) | 22.03 (-2.19%) | 22.35 (-0.74%) | 21.51 (-4.48%) |
| RG | 27.32 | 26.89 (-1.59%) | 26.72 (-2.19%) | 27.09 (-0.84%) | 25.97 (-4.96%) |
| RadCliQ () | 3.10 | 3.12 (-0.65%) | 3.12 (-0.65%) | 3.11 (-0.32%) | 3.15 (-1.61%) |
| CheXbert vector | 42.02 | 41.57 (-1.07%) | 41.37 (-1.54%) | 41.92 (-0.24%) | 40.93 (-2.59%) |
| CheXpert-F1: | |||||
| Micro-F1-14 | 52.48 | 51.74 (-1.42%) | 51.68 (-1.53%) | 52.13 (-0.67%) | 51.13 (-2.57%) |
| Macro-F1-14 | 36.87 | 36.04 (-2.25%) | 36.12 (-2.03%) | 36.14 (-1.97%) | 35.85 (-2.76%) |
| Micro-F1-5 | 56.63 | 55.37 (-2.23%) | 55.79 (-1.49%) | 55.87 (-1.34%) | 54.51 (-3.74%) |
| Macro-F1-5 | 49.33 | 47.76 (-3.18%) | 47.82 (-3.06%) | 47.98 (-2.75%) | 47.22 (-4.28%) |
To evaluate the impact of TAC on Libra’s performance in the RRG task, we initialised a model with the RAD-DINO (Pérez-García et al., 2024) image encoder, TAC, and Meditron (Chen et al., 2023b) as the LLM. A baseline experiment (Libra-1) was conducted by fine-tuning only the TAC for the Findings generation task. As shown in Table 4, we performed ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC. Removing TFM restricted the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023) but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions. Without LFE, the model follows the LLaVA setup, utilising the penultimate layer of the image encoder. The results indicate that removing any TAC submodule leads to a decline in all metrics compared to Libra-1. Removing TFM caused a notable drop in the F1 score (2%), emphasising its role in capturing temporal information. Removing LFE especially decreases RadGraph-related scores, highlighting its importance in extracting detailed image features. The removal of PIPB impacted clinical scores more than lexical scores. Finally, removing the entire TAC results in substantial declines across all metrics, underscoring its critical role in integrating image details and temporal information666For more ablation studies, including the impact of radiology-specific pre-trained models, see Appx. C..
Q3: Are additional Impression and VQA datasets needed during the feature alignment?
| Metric | Stage: 1 | Stage: 2 | ||
| Libra | Libra-f | Libra | Libra-f | |
| Lexical: | ||||
| ROUGE-L | 27.27 | 27.56△ | 36.66 | 35.31▼ |
| BLEU-1 | 41.24 | 34.84▼ | 51.25 | 49.92▼ |
| BLEU-4 | 13.59 | 11.51▼ | 24.54 | 23.05▼ |
| METEOR | 39.44 | 35.50▼ | 48.90 | 47.99▼ |
| BERTScore | 56.00 | 55.87▼ | 62.50 | 61.28▼ |
| F1 | 24.80 | 26.63△ | 35.34 | 33.52▼ |
| Clinical: | ||||
| RadGraph-F1 | 20.45 | 22.52△ | 32.87 | 30.77▼ |
| RG | 25.19 | 27.32△ | 37.27 | 35.44▼ |
| RadCliQ () | 3.31 | 3.10△ | 2.72 | 2.83▼ |
| CheXbert vector | 35.33 | 42.02△ | 46.85 | 45.32▼ |
| CheXpert-F1: | ||||
| Micro-F1-14 | 43.63 | 52.48△ | 55.87 | 54.11▼ |
| Macro-F1-14 | 25.68 | 36.87△ | 40.38 | 37.16▼ |
| Micro-F1-5 | 49.75 | 56.63△ | 60.07 | 58.76▼ |
| Macro-F1-5 | 40.40 | 49.33△ | 53.75 | 51.99▼ |
To evaluate the impact of dataset expansion during the first stage on Libra’s performance, we trained a model (Libra-f) using only the Findings data in the first stage, while Libra incorporated additional Impression and VQA data for alignment, in Table 5. After the first stage, Libra showed superior performance in lexical metrics compared to Libra-f, but a slight decline in clinical metrics. This can be attributed to the inclusion of VQA tasks, which encouraged the model to focus on more fine-grained and grounded information. While VQA emphasises the detailed description of the single symptom, the Findings section generation task requires a comprehensive overview of the CXR, encompassing multiple normal and abnormal findings. This shift in focus also impacted the F1 score, as each finding typically involves its own temporal changes. Additionally, the reduction in the number of generated findings entities led to a decrease in the number of identified temporal entities. In the second stage, fine-tuned on the Findings dataset, Libra achieved the best results across all metrics, indicating that incorporating additional datasets in the first stage enhances Libra’s understanding and reasoning of CXRs.
5 Performance Analysis
We used CXRs from the official test split to generate Findings sections and perform a qualitative analysis of Libra. Figure 2 illustrates a case where no true prior image was available for reference. For example, while the ground truth report only mentioned the presence of “sternal wires,” Libra not only identified their presence but also provided details on the specific type. This demonstrates Libra can deliver more precise clinical information by not only recognizing conditions but also distinguishing the specific types of surgical interventions.
Original
report
The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Sternal wires are intact.
Libra
The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Median sternotomy wires are intact.
For the samples with true prior images, Libra accurately describes the presence and changes in radiological conditions. As shown in the Figure 3, the current image revealed additional findings of pleural effusion and pneumonia compared to the prior image. The model effectively detected these progressive changes, provided detailed descriptions, and suggested potential causes for further investigation. Notably, it explicitly mentioned the comparative information relative to the prior image, offering a clear understanding of the temporal changes.
Current image
Prior image
Original report (current image)
In comparison
with the study of , there is again evidence of a superior mediastinal mass consistent with the known thyroid enlargement that is better seen on the recent CT scan. Continued bibasilar opacification most likely represents atelectasis, though in the appropriate clinical setting, superimposed pneumonia would have to be considered. The cardiac silhouette is within normal limits, though there may be some mild elevation of pulmonary venous pressure. Small pleural effusions are probably present bilaterally.
Libra
As compared to
radiograph, cardiomediastinal contours are stable. New patchy and linear opacities have developed at the lung bases, right greater than left, and may reflect patchy atelectasis, aspiration, or developing infectious pneumonia. Small pleural effusions are present bilaterally.
w/ dummy prior image
Cardiac silhouette is upper limits of normal in size and accompanied by pulmonary vascular congestion and minimal interstitial edema. Patchy bibasilar opacities are present, and could reflect patchy atelectasis, aspiration, or developing infectious pneumonia. Followup radiographs may be helpful in this regard. Small bilateral pleural effusions are also demonstrated.
Swap order
Cardiomediastinal contours are normal. Previously present bibasilar opacities have nearly completely resolved with minimal residual linear opacification remaining at the right lung base. No new areas of consolidation are identified, and there are no pleural effusions.
Original report (prior image)
The heart is normal in size. The mediastinal and hilar contours appear stable. The aortic arch is calcified. There is no definite pleural effusion or pneumothorax. Streaky opacity in the right costophrenic angle suggests minor atelectasis, which has decreased. There is similar mild spinal curvature.
To further assess the model’s robustness, we conducted a more detailed analysis of this case. First, we generated a report using only the dummy prior image as a reference. While Libra successfully detected abnormal findings, it failed to specify changes relative to the previous image, describing only the current state without indicating disease progression. Moreover, it provided less clarity in distinguishing the distribution and severity of abnormalities. This suggests that, without a true prior image for comparison, the model struggles to accurately describe the progression of findings.
Evaluating Temporal Consistency in Libra’s Report Generation To investigate the model’s ability to capture temporal changes, we swapped the order of the images, using the true prior image as the current image and vice versa. Interestingly, the generated report indicated an improvement in the patient’s condition, with most abnormalities resolved and no new findings detected, which is contrary to the ground truth of the original current image report. Moreover, the report generated from the swapped images closely resembled the original report of the true prior image, as shown at the bottom of Figure 3, which described a stable or improving condition for the patient. This striking observation demonstrates that Libra can effectively utilise both current and prior studies to generate accurate reports for assessing changes, simulating the conditions of standard clinical practice.
6 Conclusion
In this study, we presented Libra, a temporal-aware multimodal large language model specifically designed for chest X-ray report generation tasks. Trained exclusively on the open-access MIMIC-CXR dataset (Johnson et al., 2019b), Libra utilises a two-stage training framework that leverages a radiology-specific pre-trained image encoder and language model, connected through a specially designed Temporal Alignment Connector to bridge visual and textual modalities. By identifying observations from current CXR and referencing prior scan, Libra generates radiology reports with substantial improvements across all metrics compared to existing models of the same parameter scale. Qualitative analysis in the case study highlights Libra’s ability to effectively utilise the temporal relationships between current and prior images, addressing the challenge of hallucinations related to prior study references in RRG tasks. Libra sets a new paradigm for MLLMs in multimodal medical AI research. In the future, we aim to further enhance Libra’s clinical applicability and accuracy, paving the way for the development of general-purpose AI models in the medical domain. For a detailed discussion of Libra’s limitations and future work, please refer to Appx. A.3.
Appendix A Appendix A
A.1 Related Work
Radiology Report Generation Several efforts have been made to develop models tailored for radiology report generation (RRG), driven by the need to address the long-tail distribution of observations in chest X-rays and provide fine-grained descriptions of findings. This task has gained prominence as an important objective for the designed system (Wang et al., 2018b).
Such systems inherently require a language generation component. Initially, recurrent neural networks were used (Liu et al., 2019), but these have recently been replaced by Transformer architectures (Miura et al., 2021b; Chen et al., 2022), including large language models (LLMs) such as PaLM (Chowdhery et al., 2022) and Vicuna-7B (Chiang et al., 2023).
To ensure clinical accuracy in generated text, some studies have moved beyond standard language modelling losses, instead employing reinforcement learning (RL) to optimise for rewards that prioritise “clinically relevant” features, such as the inclusion of specific findings (Liu et al., 2019; Irvin et al., 2019b) or maintaining logical consistency (Miura et al., 2021a; Delbrouck et al., 2022a). However, these methods often depend on external models like CheXbert (Smit et al., 2020) or RadGraph (Jain et al., 2021) to extract clinical entities, adding complexity to the optimisation process.
With the advancement of LLMs, an increasing number of studies have demonstrated that plain auto-regressive language modelling can achieve substantial results. However, we acknowledge that the benefits of more complex training objectives or RL-based approaches may be complementary. Consequently, research on leveraging temporal information in RRG tasks can be broadly categorised into two types: LLM-based methods and non-LLM-based methods.
non-LLM-based Model These architectures are typically transformer encoder-decoder models and their variants. They are often composed of multiple modules that require separate training. Meanwhile, they handle single and double image inputs by symbolically differentiating tasks and employing two distinct neural network architectures: one designed for single-image inputs and another for double-image inputs.
Serra et al. (2023) utilises symbolic alignment in its Longitudinal Projection Module and employs a separately trained BERT-based (Devlin et al., 2019) text generator. RECAP (Hou et al., 2023a) implements a transformer encoder-decoder with symbolic task differentiation and two-stage training, beginning with classification tasks followed by report generation. TiBiX (Sanjeev et al., 2024) adopts a transformer model with causal attention layers and incorporates learnable padding tokens to handle cases without prior images. BioViL-T (Bannur et al., 2023) is a self-supervised vision-language training framework that features a CNN–Transformer hybrid multi-image encoder trained jointly with a BERT-based text model.
On one hand, the difference in model parameter sizes, and on the other, as LLM-based models generally outperform other types of models in the RRG task, papers on non-LLM-based models or those using small language models typically do not compare their methods with LLM-based approaches. At the same time, we conducted comparisons and discussions to reaffirm this point, as detailed in Appx. D.2.
LLM-based Model Current LLM-based models have achieved notable success in the RRG task, largely benefiting from visual instruction tuning (Liu et al., 2023). Structurally, these models (Li et al., 2023a; Chaves et al., 2024; Hyland et al., 2024; Zhou et al., 2024; Park et al., 2024) typically consist of an image encoder and an adapter that bridges the connection to the LLM. These adapters often rely on single-layer hidden representations (e.g., the last or penultimate layer) from pre-trained image encoders, limiting their ability to simultaneously integrate information from multiple images.
In end-to-end training, LLM-based models handle multiple image inputs by concatenating them with textual prompts to form a composite input for the LLM. For example, the input format is “<Current Image Placeholder> + <Prior Imgae Placeholder> + <Prompt>”.
Although our model adopts an end-to-end training paradigm, it does not rely on the vanilla approach to handle multiple images. Instead, we designed a novel adapter called the Temporal Alignment Connector (TAC). On one hand, TAC leverages all hidden layer features from the image encoder, setting it apart from the default configurations used in previous works. On the other hand, TAC models temporal information directly rather than simply concatenating two images in the latent space of the LLM.
Radiological Image Representation As radiology lies outside the distribution of most general-domain image models, a radiology-specific pre-trained model is crucial for the RRG task (Pérez-García et al., 2024).
Zhou et al. (2023) introduces a unified framework, Masked Record Modeling (MRM), which combines self-supervision and radiology report supervision to enhance radiograph representation learning. BioViL-T (Bannur et al., 2023) integrates a CNN-Transformer hybrid architecture for multimodal modelling, leveraging temporal structures in medical imaging and reports to perform tasks such as disease progression classification and report generation. BiomedCLIP (Zhang et al., 2024b) is a multimodal biomedical foundational model pre-trained on various biomedical tasks. RAD-DINO (Pérez-García et al., 2024) is a medical image encoder that employs a pure image-based self-supervised learning approach from DINOv2 (Oquab et al., 2024) for continuous pretraining, focusing exclusively on image data to avoid the limitations of text supervision.
In recent parallel works, RAD-DINO has been widely applied to RRG tasks, including MAIRA-2 (Bannur et al., 2024), M4CXR (Park et al., 2024). Furthermore, Pérez-García et al. (2024) highlighted that RAD-DINO outperforms other types of image encoders in RRG tasks. Therefore, our model also adopts RAD-DINO as the image encoder.
A.2 Research Objectives
Temporal Information Temporal change is a crucial factor in understanding disease progression, especially in radiology, where paired images and corresponding radiology reports capture changes over time. This temporal context is often reflected in comparisons between current and prior scans, documenting symptom evolution or newly identified findings.
The relativity of position within the timeline is a key determinant of temporal information. This relative positioning does not alter the established description of well-established facts about symptoms already present in the current image. For instance, if the prior image is recent, descriptions of changes will be minimal, whereas an older prior image may highlight more pronounced changes.
Research Object Our study focuses on the chest X-ray (frontal view), rather than the individual patient. Specifically, we aim to generate a report for the current image while referencing the temporal information provided by a single prior image as auxiliary information.
Our model is primarily designed to handle scenarios with limited temporal information in the RRG task. For instance, in a case where a patient has two scans taken just milliseconds apart, the current and prior images would be nearly identical, as no pathological changes would manifest within such a short interval. This extreme scenario demonstrates how the model handles clinical practice under limited temporal information. In such cases, the correct diagnosis for this minimal interval would be that the patient’s condition is “stable.” In our design, the dummy prior image addresses scenarios where a true prior image is absent.
However, patients often undergo multiple prior scans, which may include temporal information from different orientations. This more complex temporal information lies beyond the scope of our current study. We will provide a detailed discussion of this scenario in Appx. A.3.
A.3 Limitations and Future Work
While our work represents a step forward in leveraging temporal information for radiology report generation, it also has several limitations that warrant further exploration.
Handling Multiple Prior Scans Our current model is designed to process a single prior scan alongside the current scan. While this approach aligns with standard clinical workflows, which typically prioritise the most recent prior study for comparisons, it overlooks scenarios where multiple prior scans could offer a richer temporal perspective. For instance, analysing a sequence of images spanning an extended period could provide deeper insights into gradual disease progression. Future efforts should focus on extending our framework to incorporate multiple prior scans efficiently, enabling a more nuanced understanding of temporal patterns in clinical data.
Temporal Information Beyond Image Comparisons At present, our model captures temporal information through paired image comparisons and their corresponding textual reports. However, clinical assessments often draw upon a broader context, including historical notes, laboratory results, and other longitudinal patient data. Expanding our approach to integrate these diverse temporal data sources could facilitate a more holistic understanding of disease trajectories and patient history, significantly enhancing clinical applicability.
Challenges with Sparse Temporal Data In cases where prior scans are unavailable or only minimally informative (e.g., scans taken within a very short time interval), our use of a “dummy prior image” addresses the lack of temporal data. However, the model’s ability to interpret and generate meaningful outputs under these constraints may still be limited. Future work could explore methods for imputing or synthesising temporal context to address such cases more effectively.
Computational Complexity Our use of temporal alignment mechanisms and multi-layer feature integration introduces increased computational demands. This complexity poses challenges for real-time deployment, particularly in resource-constrained clinical environments.
Generalisability Across Modalities and Datasets This study focuses exclusively on frontal-view chest X-rays and utilises the MIMIC-CXR dataset (Johnson et al., 2019b). However, our approach has not been evaluated on other imaging modalities (e.g., CT or MRI) or datasets such as CheXpert (Irvin et al., 2019b) or PadChest (Bustos et al., 2020).
Based on the identified limitations, we propose several directions for future research and development:
Develop frameworks for integrating multiple prior scans with dynamic temporal reasoning.
Extend the model to handle multi-modal imaging and textual data for comprehensive diagnostics.
Investigating the integration of temporal data from diverse clinical sources, including electronic health records (EHRs)
Exploring lightweight model architectures for faster inference while maintaining high performance.
These enhancements will not only address the current limitations but also extend the potential of temporal-aware multimodal models in radiology and beyond.
Appendix B Appendix B
B.1 Training Configuration
We train the model using a standard auto-regressive language modelling loss (cross-entropy). For this study, we utilise Meditron-7b (Chen et al., 2023b) as the LLM, with a total batch size of 16 throughout the training process. The training is conducted on a computational infrastructure equipped with A6000 GPU (48GB of memory) and configured with DeepSpeed ZeRO-3 optimization (Rajbhandari et al., 2020) and BF16 precision enabled. A cosine learning rate scheduler is employed, starting with a warm-up phase of . In the first stage of training, we run for 1 epoch with a learning rate of . In the second stage, the model is trained for 3 epochs at the same learning rate. The LoRA (Hu et al., 2021) parameters are set to and . The final checkpoint for all runs is selected based on the observation of the minimum loss on the evaluation dataset throughout the training process.
B.2 Detailed Description of Datasets
MIMIC-CXR. (Johnson et al., 2019b) This is a large, publicly accessible dataset comprising 377,110 DICOM images across 227,835 studies, each accompanied by a radiology report (Johnson et al., 2019b). For images, we use the commonly available JPEG files from MIMIC-CXR-JPG (Johnson et al., 2019a), rather than the original DICOM files, and we preprocess the dataset to exclude non-AP/PA scans. For each report, we extract the Findings, Impression, Indication, History, Comparison, and Technique sections using rule-based heuristics supported by the official MIMIC code repository (Johnson et al., 2018).
For the Findings section generation task, we discard studies where the Findings could not be extracted, while allowing for missing content in other clinical sections. The same approach is applied to the Impression section generation task. In all our experiments, we adhere to the official MIMIC-CXR dataset split. Additionally, we retrieve prior images by following the chronological order of studies as indicated by the official labels, selecting the closest prior study as the reference image. It is important to note that, to prevent data leakage between the train, validation, and test sets, prior images are retrieved only from within the same split.
Medical-Diff-VQA. (Hu et al., 2023) This dataset is a derivative of the MIMIC-CXR dataset, focused on identifying differences between pairs of main and reference images. The data split adheres to the original labelling, ensuring no data leakage occurs. In total, this dataset comprises 700,703 question-answer pairs derived from 164,324 main-reference image pairs. The questions are divided into seven categories: abnormality, location, type, view, presence, and difference. Each pair consists of a main (current) image and a reference (prior) image, both taken from different studies of the same patient. The reference image is always selected from an earlier visit, with the main image representing the later visit. Of the seven question types, the first six types focus on the main image, while the “difference” questions involve both images.
MIMIC-Ext-MIMIC-CXR-VQA. (Bae et al., 2023) This dataset is an extension of the MIMIC-CXR dataset, specifically designed for VQA tasks within the medical domain, specifically focused on CXRs. It includes questions generated from 48 unique templates covering seven content types: presence, anatomy, attribute, abnormality, size, plane, and gender. Each template was developed with the guidance of board-certified medical experts to ensure clinical relevance, addressing both standard medical VQA content and more complex logical scenarios. In total, the dataset consists of 377,391 unique entries. Since annotations are based on single images, the current image serves as a dummy prior image for all entries in our experiment.
For this study, we carefully selected datasets that provide complete reports and temporal information (i.e., prior images) to ensure alignment with our research objectives for the RRG task. After thoroughly evaluating other datasets, we found them unsuitable for the following reasons:
CheXpert. (Irvin et al., 2019b) This dataset includes annotated scans with label-specific annotations rather than full medical reports. While useful for training image encoders or annotation models, it is not appropriate for the RRG task, which requires complete diagnostic reports.
PadChest. (Bustos et al., 2020) Although it includes reports and corresponding prior images, its reports are in Spanish, placing cross-language training beyond the scope of our model.
IU-Xray. (Demner-Fushman et al., 2016) This dataset lacks patient-level metadata and prior study information, which is critical for our focus on temporal information in chest X-rays.
Chest ImaGenome Dataset. (Wu et al., 2021a)(a derivative of MIMIC-CXR) This dataset does not follow the official MIMIC-CXR split, risking data leakage across train, validation, and test sets.
B.3 Evaluation Metrics Details
Lexical Metrics We employed common natural language generation metrics to quantify the overlap between generated and reference reports. Specifically, ROUGE-L (Lin, 2004) measures the length of the longest common subsequence between the generated and reference reports. BLEU- (Papineni et al., 2002) calculates n-gram precision and applies a brevity penalty to discourage overly short predictions. METEOR (Banerjee & Lavie, 2005), computes the weighted harmonic mean of unigram precision and recall, with an additional penalty for fragmenting consecutive word sequences. Finally, we report BERTScore (Zhang et al., 2020a), which leverages pre-trained contextual embeddings from BERT (Devlin et al., 2019) to match words in candidate and reference sentences based on cosine similarity. We used default parameters for all of these evaluation metrics.
For radiology-specific metrics, we used as many of the same evaluation scores as possible from previous studies (Tu et al., 2023; Hyland et al., 2024; Bannur et al., 2024; Chaves et al., 2024), including the following:
RadGraph-based metrics The RadGraph model (Jain et al., 2021) is designed to parse radiology reports into structured graphs. These graphs consist of clinical entities, which include references to anatomy and observations, as well as the relationships between these entities. This structured representation enables a more detailed and systematic analysis of radiology reports, facilitating downstream tasks such as information extraction, report generation, and clinical decision support.
These include RadGraph-F1 (Jain et al., 2021), which computes the overlap in entities and relations separately and then reports their average. And a variant of it, RG (Delbrouck et al., 2022b), which matches entities based on their text, type, and whether they have at least one relation777RG is implemented as F1RadGraph with reward=partial by the radgraph package..
CheXpert F1 This set of metrics utilizes the CheXbert automatic labeler (Smit et al., 2020) to extract “present”, “absent”, or “uncertain” labels for each of the 14 CheXpert pathologies (Irvin et al., 2019a) from the generated reports and their corresponding references. In line with prior work, we report CheXpert-F1 for all 14 classes, as well as for the 5 most common findings in real-world CXR reports, referring to these as ‘[Macro/Micro]-F1-[5/14]’.
CheXbert vector similarity We also employ CheXbert vector similarity (Yu et al., 2022), which calculates the cosine similarity between the embeddings of the generated and reference reports after processing them through the CheXbert model (Smit et al., 2020).
RadCliQ In addition, we utilise RadCliQ (Radiology Report Clinical Quality), a composite metric that combines RadGraph-F1 and BLEU scores in a linear regression model to estimate the number of errors that radiologists are likely to detect in a report. To maintain consistency with previous research, we use version of it.
Both the CheXbert vector similarity, RadCliQ, and RadGraph-F1 metrics are calculated using the code released by Yu et al. (2022).
Temporal Entity F1 The , which is specifically designed to detect temporal entities. This score is neither a traditional lexical metric nor a adiology-specific metrics. Here is an example to help you understand better, as shown in Table 6.
| Ground Truth | Candidate | ROUGE-L | RadGraph-F1 | |
| Compare with prior scan, pleural effusion has worsened. | The pleural effusion has progressively worsened since previous scan. | 0.47 | 0.86 | 1.0 |
| The pleural effusion is noted again on the current scan. | 0.22 | 0.80 | 0.0 |
From these results, it is clear that the differences in lexical (ROUGE-L) and clinical (RadGraph-F1) metrics between the two candidates are relatively smaller compared to the score. This highlights that Temporal Entity F1 effectively captures and evaluates the quality of temporal information in radiology reports, distinguishing it more accurately than other standard metrics in the context of temporal information descriptions.
B.4 Prompt Example
Here, we select examples from the MIMIC-CXR (Johnson et al., 2019b) dataset as instances and further synthesise them using GPT-4 (OpenAI et al., 2024) to avoid any ethical concerns, as shown in Table 7. Following Hyland et al. (2024)’s work, we adopt a rule-based method to extract relevant sections from the report of the current image, utilising a fixed system prompt and a dynamic clinical prompt for each current scan. We utilised four clinical instructions from the original report: {Indication}, {History}, {Comparison}, and {Technique}. Notably, unlike MAIRA-2 (Bannur et al., 2024), we do not utilise the report of the prior image.
| Original Radiology Report |
| EXAMINATION: Chest (Portable AP) |
| INDICATION: Dyspnea and cough, right-sided back pain. |
| HISTORY: Intubation with pulmonary edema. |
| COMPARISON: Chest radiographs on ___ and CT chest without contrast on ___. |
| TECHNIQUE: Portable upright chest radiograph. |
| FINDINGS: In comparison with the prior study, there are diffuse bilateral pulmonary opacifications, more prominent on the right. These findings could indicate severe pulmonary edema, but superimposed pneumonia or developing ARDS cannot be excluded. Monitoring and support devices are appropriately positioned. |
| Prompt Content |
| [System prompt] { |
| The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes. |
| Image Representation Placeholder |
| [Clinical prompt] { |
| Provide a detailed description of the findings in the radiology image. Following clinical context: |
| Indication: Dyspnea and cough, right-sided back pain. |
| History: Intubation with pulmonary edema. |
| Comparison: Chest radiographs on ___ and CT chest without contrast on ___. |
| Technique: Portable upright chest radiograph. |
Appendix C Appendix C
C.1 Impact of the Temporal Alignment Connector under General-Domain Pre-trained Models
| Metric | Libra-b | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
| Lexical: | |||||
| ROUGE-L | 27.26 | 26.80 (-1.69%) | 26.57 (-2.53%) | 27.00 (-0.95%) | 24.58 (-9.83%) |
| BLEU-1 | 34.94 | 33.61 (-3.81%) | 33.68 (-3.61%) | 34.47 (-1.35%) | 31.40 (-10.13%) |
| BLEU-4 | 11.74 | 10.97 (-6.56%) | 10.94 (-6.81%) | 11.57 (-1.45%) | 8.89 (-24.28%) |
| METEOR | 35.37 | 34.50 (-2.46%) | 34.30 (-3.03%) | 34.93 (-1.24%) | 32.41 (-8.37%) |
| BERTScore | 55.51 | 54.97 (-0.97%) | 54.75 (-1.37%) | 55.26 (-0.45%) | 53.07 (-4.40%) |
| F1 | 24.77 | 23.54 (-4.97%) | 24.00 (-3.11%) | 24.68 (-0.36%) | 22.52 (-9.08%) |
| Clinical: | |||||
| RadGraph-F1 | 21.67 | 21.06 (-2.81%) | 20.73 (-4.34%) | 21.35 (-1.48%) | 19.77 (-8.77%) |
| RG | 26.28 | 25.45 (-3.16%) | 25.14 (-4.34%) | 25.84 (-1.67%) | 23.74 (-9.67%) |
| RadCliQ () | 3.17 | 3.20 (-0.95%) | 3.22 (-1.58%) | 3.18 (-0.32%) | 3.27 (-3.15%) |
| CheXbert vector | 39.58 | 38.74 (-2.12%) | 38.37 (-3.06%) | 39.49 (-0.23%) | 37.56 (-5.10%) |
| CheXpert-F1: | |||||
| Micro-F1-14 | 49.06 | 47.68 (-2.81%) | 47.57 (-3.04%) | 48.40 (-1.35%) | 46.57 (-5.08%) |
| Macro-F1-14 | 33.07 | 31.60 (-4.45%) | 31.78 (-3.90%) | 31.78 (-3.90%) | 31.27 (-5.44%) |
| Micro-F1-5 | 54.55 | 52.14 (-4.42%) | 52.94 (-2.95%) | 53.10 (-2.66%) | 50.72 (-7.02%) |
| Macro-F1-5 | 47.24 | 44.28 (-6.27%) | 44.39 (-6.04%) | 44.68 (-5.42%) | 43.48 (-7.96%) |
Domain-specific pre-trained models (i.e., RAD-DINO (Pérez-García et al., 2024) and Meditron (Chen et al., 2023b)) inherently incorporate domain-specific knowledge, such as phrasing conventions, pronoun usage, and even temporal information embedded in the training corpus.
To eliminate the influence of such factors, we used a general-domain image encoder (DINOv2 (Oquab et al., 2024)) and a LLM (Vicuna-7B-v1.5 (Chiang et al., 2023)), allowing the structural enhancements of TAC to be observed more directly. We adopted the same setup as the second ablation study in Section 3. We first conducted a baseline experiment, referred to as Libra-b, by fine-tuning only the adapter for the Findings generation task. As shown in Table 8, we then conducted ablation studies by sequentially removing different components from the model, including the TFM, LFE, Prior Image Prefix Bias (PIPB) in TFM, and the entire TAC. Removing TFM restricts the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023), but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions. It is important to note that without the TFM module, the model cannot process true prior images or dummy prior images, and is limited to only the current image as input. Without LFE, the model follows the LLaVA setup, using the penultimate layer of the image encoder to process single or paired images.
The ablation results are consistent with those observed when using domain-specific pre-trained models, as presented in Section 3. Compared to Libra-b, removing any TAC submodule results in a decline across all metrics. Removing the TFM caused a notable drop in the F1 score (4%), emphasising its role in capturing temporal information. The removal of the LFE significantly decreased RadGraph-related scores, highlighting its importance for extracting detailed image features. Removing the PIPB impacted clinical scores more than lexical scores. Finally, removing the TAC resulted in significant declines across all metrics, underscoring its critical role in integrating image details and temporal information.
C.2 Impact of the Temporal Alignment Connector After the Second Stage
| Metric | Libra-2 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
| Lexical: | |||||
| ROUGE-L | 35.31 | 35.16 (-0.42%) | 35.09 (-0.64%) | 35.23 (-0.23%) | 34.41 (-2.55%) |
| BLEU-1 | 49.92 | 49.44 (-0.97%) | 49.47 (-0.90%) | 49.75 (-0.34%) | 48.61 (-2.63%) |
| BLEU-4 | 23.05 | 22.67 (-1.66%) | 22.65 (-1.75%) | 22.97 (-0.35%) | 21.51 (-6.70%) |
| METEOR | 47.99 | 47.69 (-0.62%) | 47.62 (-0.77%) | 47.84 (-0.31%) | 46.95 (-2.16%) |
| BERTScore | 61.28 | 61.13 (-0.24%) | 61.07 (-0.34%) | 61.21 (-0.12%) | 60.60 (-1.12%) |
| F1 | 33.52 | 33.10 (-1.27%) | 33.25 (-0.79%) | 33.49 (-0.09%) | 32.73 (-2.36%) |
| Clinical: | |||||
| RadGraph-F1 | 30.77 | 30.55 (-0.72%) | 30.43 (-1.10%) | 30.65 (-0.40%) | 30.07 (-2.27%) |
| RG | 35.44 | 35.16 (-0.79%) | 35.05 (-1.10%) | 35.29 (-0.42%) | 34.55 (-2.51%) |
| RadCliQ () | 2.83 | 2.84 (-0.35%) | 2.84 (-0.35%) | 2.85 (-0.71%) | 2.85 (-0.71%) |
| CheXbert vector | 45.32 | 45.08 (-0.53%) | 44.97 (-0.77%) | 45.27 (-0.11%) | 44.73 (-1.30%) |
| CheXpert-F1: | |||||
| Micro-F1-14 | 54.11 | 53.73 (-0.70%) | 53.70 (-0.76%) | 54.00 (-0.20%) | 53.41 (-1.30%) |
| Macro-F1-14 | 37.16 | 36.74 (-1.13%) | 36.78 (-1.02%) | 36.79 (-1.00%) | 36.64 (-1.40%) |
| Micro-F1-5 | 58.76 | 58.10 (-1.12%) | 58.32 (-0.75%) | 58.36 (-0.68%) | 57.65 (-1.89%) |
| Macro-F1-5 | 51.99 | 51.16 (-1.60%) | 51.19 (-1.54%) | 51.27 (-1.38%) | 50.87 (-2.15%) |
To further evaluate the impact of the Temporal Alignment Connector (TAC) on Libra’s performance, we followed the setup of the second ablation study in Section 3. After the first stage of alignment, the model underwent a second stage of fine-tuning. This stage was designed to optimise the model’s performance on the Findings section generation task by leveraging the aligned visual and textual features learned during the initial stage.
In this phase, we applied Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM (Meditron (Chen et al., 2023b)), while keeping the visual encoder (RAD-DINO (Pérez-García et al., 2024)) and TAC weights frozen. The baseline for this experiment is Libra-2 (in Table 9), which is derived from Libra-1 (in Table 4) after undergoing LoRA fine-tuning.
We conducted ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC. The conclusions remain consistent with those in Section 3, demonstrating that removing any TAC submodule results in a decline across all metrics compared to Libra-2. This indicates that the performance improvements brought by TAC are stable and unaffected by changes in training stages. It further confirms that TAC has effectively embedded the capability to process temporal information within the model.
C.3 Robustness Assessment of the Temporal Alignment Connector
| Metric | Libra-3 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
| Lexical: | |||||
| ROUGE-L | 35.58 | 35.53 (-0.14%) | 35.51 (-0.21%) | 35.55 (-0.08%) | 35.28 (-0.86%) |
| BLEU-1 | 49.54 | 49.38 (-0.32%) | 49.39 (-0.30%) | 49.48 (-0.11%) | 49.10 (-0.88%) |
| BLEU-4 | 23.61 | 23.48 (-0.55%) | 23.47 (-0.58%) | 23.58 (-0.12%) | 23.07 (-2.28%) |
| METEOR | 47.61 | 47.51 (-0.21%) | 47.49 (-0.26%) | 47.56 (-0.10%) | 47.26 (-0.73%) |
| BERTScore | 61.54 | 61.49 (-0.08%) | 61.47 (-0.11%) | 61.52 (-0.04%) | 61.31 (-0.37%) |
| F1 | 33.51 | 33.37 (-0.42%) | 33.42 (-0.27%) | 33.50 (-0.03%) | 33.24 (-0.79%) |
| Clinical: | |||||
| RadGraph-F1 | 29.82 | 29.75 (-0.24%) | 29.71 (-0.37%) | 29.78 (-0.13%) | 29.59 (-0.76%) |
| RG | 35.60 | 35.51 (-0.26%) | 35.47 (-0.37%) | 35.55 (-0.14%) | 35.30 (-0.84%) |
| RadCliQ () | 2.91 | 2.92 (-0.34%) | 2.92 (-0.34%) | 2.91 ( – ) | 2.93 (-0.68%) |
| CheXbert vector | 44.77 | 44.69 (-0.18%) | 44.65 (-0.26%) | 44.75 (-0.04%) | 44.57 (-0.45%) |
| CheXpert-F1: | |||||
| Micro-F1-14 | 52.45 | 52.33 (-0.23%) | 52.32 (-0.25%) | 52.41 (-0.08%) | 52.22 (-0.44%) |
| Macro-F1-14 | 30.77 | 30.65 (-0.38%) | 30.66 (-0.34%) | 30.67 (-0.33%) | 30.63 (-0.47%) |
| Micro-F1-5 | 54.42 | 54.22 (-0.38%) | 54.28 (-0.25%) | 54.30 (-0.23%) | 54.08 (-0.63%) |
| Macro-F1-5 | 44.58 | 44.34 (-0.54%) | 44.35 (-0.52%) | 44.37 (-0.46%) | 44.26 (-0.72%) |
In this experiment, we introduced an additional round of LoRA fine-tuning to induce overtraining and assess the robustness of the Temporal Alignment Connector (TAC). Following the setup in Appx. C.2, after integrating the first LoRA weights, we reinitialised a new set of LoRA adapters for the LLM and conducted further training under the same second-stage fine-tuning configuration for one epoch. The baseline for this experiment is Libra-3 (as shown in Table 10), which is derived from Libra-2 (illustrated in Table 9) following this additional fine-tuning step.
The results reveal that, compared to Libra-2, Libra-3 demonstrates minimal changes in lexical scores, while clinical scores decline due to overfitting caused by the additional fine-tuning. Notably, the CheXpert (Smit et al., 2020) (Macro-F1-[5/14]) scores exhibit the most influential reduction.
However, the ablation studies reaffirm that the performance improvements introduced by TAC are robust and remain unaffected by variations in training strategies. This robustness is attributed to TAC’s ability to effectively capture and preserve temporal image representations during the first training phase, which are subsequently preserved throughout further fine-tuning.
These findings highlight that TAC reliably embeds the capacity for temporal information processing into the model, making it a stable and essential component for handling temporal data in RRG tasks.
C.4 Impact of Radiology-Specific Pre-trained Models
| Metric | Libra-1 | w/o RadDINO | w/o Meditron | w/o RadDINO+Meditron |
| Lexical: | ||||
| ROUGE-L | 27.56 | 27.66 (0.36%) | 27.29 (-0.98%) | 27.26 (-1.09%) |
| BLEU-1 | 34.84 | 35.32 (1.38%) | 34.91 (0.20%) | 34.94 (0.29%) |
| BLEU-4 | 11.51 | 12.56 (9.12%) | 11.61 (0.87%) | 11.74 (2.00%) |
| METEOR | 35.50 | 35.65 (0.42%) | 35.53 (0.08%) | 35.37 (-0.37%) |
| BERTScore | 55.87 | 55.89 (0.04%) | 55.58 (-0.52%) | 55.51 (-0.64%) |
| F1 | 26.63 | 25.53 (-4.13%) | 24.78 (-6.95%) | 24.77 (-6.98%) |
| Clinical: | ||||
| RadGraph-F1 | 22.52 | 22.11 (-1.82%) | 23.13 (2.71%) | 21.67 (-3.77%) |
| RG | 27.32 | 26.72 (-2.20%) | 27.53 (0.77%) | 26.28 (-3.81%) |
| RadCliQ () | 3.10 | 3.13 (-0.97%) | 3.08 (0.65%) | 3.17 (-2.26%) |
| CheXbert vector | 42.02 | 40.78 (-2.95%) | 41.94 (-0.19%) | 39.49 (-6.02%) |
| CheXpert-F1: | ||||
| Micro-F1-14 | 52.84 | 51.55 (-2.44%) | 51.45 (-2.63%) | 49.06 (-7.15%) |
| Macro-F1-14 | 36.87 | 34.58 (-6.21%) | 37.20 (0.90%) | 33.07 (-10.31%) |
| Micro-F1-5 | 56.63 | 55.00 (-2.88%) | 55.39 (-2.19%) | 54.55 (-3.67%) |
| Macro-F1-5 | 49.33 | 47.26 (-4.20%) | 47.62 (-3.47%) | 47.24 (-4.24%) |
Aligning radiology images with textual information is a key challenge in RRG tasks. To demonstrate the benefits of using radiology-specific pre-trained models for more accurate feature representation and improved MLLM performance, we initialised a Libra model with RadDINO, the TAC, and Meditron-7b, conducting the first stage of training, denoted as Libra-1 (This is consistent with the baseline setup of the second ablation study in Section 3). Then we replaced the image encoder and LLM with their general-domain counterparts, DINOv2 and Vicuna-7B-v1.5, respectively. Finally, we replaced both components, which is also referred to as Libra-b.
As shown in Table 11, substituting radiology-specific pre-trained models with general-domain models resulted in a notable decline in clinical scores, while the impact on lexical scores was minimal. Notably, replacing the radiology-specific image encoder caused a more pronounced decline in clinical metrics compared to replacing the language model. This suggests that accurate medical image representation provides greater benefits in RRG tasks, indicating the importance of incorporating domain-specific knowledge into pre-trained models to enhance Libra’s performance.
C.5 Incremental Study
In this section, we conducted an incremental study of our model, mirroring the logical framework we used to construct Libra. We started with an architecture similar to LLaVA, comprising a pre-trained CLIP image encoder, a randomly initialised four-layer MLP as the adapter, and Vicuna-7B-v1.5 as the LLM. We train the adapter on the Findings section generation task to establish an initial baseline.
Gradual enhancements are then introduced, as shown in the Table 12. First, the image encoder is replaced with DINOv2. Building upon this, we integrate the LFE, the prefix module of the TAC connector. Subsequently, we add the TFM, the suffix module, thus completing the full TAC connector. Following this, the image encoder and LLM are progressively replaced with biomedical domain models, RAD-DINO and Meditron, respectively. The dataset used for pretraining in the first stage is then extended. Finally, the model is fine-tuned for downstream tasks to obtain Libra.
| Metric | Stage 1: Temporal Feature Alignment | Stage 2 | ||||||
| ∗Initial | DINO | LFE | TFM | RAD-DINO | Meditron | Dataset | Libra | |
| Lexical: | ||||||||
| ROUGE-L | 23.77 | 24.58 | 26.57 | 27.26 | 27.29 | 27.56 | 27.27 | 36.66 |
| BLEU-1 | 31.48 | 31.40 | 33.68 | 34.94 | 34.91 | 34.84 | 41.24 | 51.25 |
| BLEU-4 | 8.41 | 8.89 | 10.94 | 11.74 | 11.61 | 11.51 | 13.59 | 24.54 |
| METEOR | 32.1 | 32.41 | 34.3 | 35.37 | 35.53 | 35.50 | 39.44 | 48.90 |
| BERTScore | 52.76 | 53.07 | 54.75 | 55.51 | 55.58 | 55.87 | 56.00 | 62.50 |
| F1 | 21.60 | 22.52 | 24.00 | 24.77 | 24.78 | 26.63 | 24.80 | 35.34 |
| Clinical: | ||||||||
| RadGraph-F1 | 18.58 | 19.70 | 20.73 | 21.67 | 23.13 | 22.52 | 20.45 | 32.87 |
| RG | 23.05 | 23.74 | 25.14 | 26.28 | 27.53 | 27.32 | 25.19 | 37.57 |
| RadCliQ() | 3.35 | 3.26 | 3.22 | 3.17 | 3.08 | 3.10 | 3.31 | 2.72 |
| CheXbert vector | 35.59 | 37.94 | 38.37 | 39.49 | 41.94 | 42.02 | 35.33 | 46.85 |
| CheXpert-F1: | ||||||||
| Micro-F1-14 | 44.75 | 46.57 | 47.57 | 49.06 | 51.45 | 52.48 | 43.63 | 55.87 |
| Macro-F1-14 | 25.13 | 31.27 | 31.07 | 33.07 | 37.20 | 36.87 | 25.68 | 40.38 |
| Micro-F1-5 | 45.97 | 50.72 | 52.94 | 54.55 | 55.39 | 56.63 | 49.75 | 60.07 |
| Macro-F1-5 | 36.55 | 43.48 | 44.39 | 47.24 | 47.62 | 49.33 | 40.40 | 53.75 |
We observed that with each step of improvement, the overall performance consistently exceeded the previous state, demonstrating the essential role of each component in our optimisation scheme. Notably, in the alignment stage, the model showed the greatest overall improvement after the addition of the TFM, highlighting its capability to capture temporal information, which significantly enhanced performance on the RRA task.
However, after the data expansion optimisation in the first stage, the scores on lexical metrics improved, while clinical metrics exhibited a slight decline. This can be attributed to the introduction of the VQA task, which led the model to perceive more fine-grained grounded information. This is distinct from the report generation task, which focuses on describing the holistic level. Furthermore, this shift also affects the F1 score, as temporal entities are often associated with the presence of specific symptoms. Ultimately, these dropped scores are improved again through the second stage of fine-tuning.
Evaluation of Libra’s Temporal Awareness. Another approach to investigating the model’s ability to capture temporal information is to evaluate it separately within the test split based on the presence or absence of true prior images, as shown in Table 13.
With the addition of the TFM, the model exhibited temporal awareness. It is worth noting that, for the first time, the F1 score of samples with prior images surpassed those without, and this trend persisted through subsequent optimisations. This indicates that the structural enhancements have resulted in a sustained improvement in the model’s temporal perception capabilities. An effective example is in Section 5.
Appendix D Appendix D
D.1 Discussion on Performance with Radiology Foundation Models
| Metric | MedVersa | MAIRA-2 | M4CXR | Libra (%) |
| Lexical: | ||||
| ROUGE-L | – | 38.4 | 28.5 | 36.7 (-4.4%) |
| BLEU-1 | – | 46.5 | 33.9 | 51.3 (10.3%) |
| BLEU-4 | 17.8 | 23.4 | 10.3 | 24.5 (4.7%) |
| METEOR | – | 42.0 | – | 48.9 (16.4%) |
| BERTScore | 49.7 | – | – | 62.5 (25.8%) |
| Clinical: | ||||
| RadGraph-F1 | 28.0 | 34.6 | 21.8 | 32.9 (-4.9%) |
| RG | – | 39.7 | – | 37.6 (-5.3%) |
| RadCliQ() | 2.7 | 2.6 | – | 2.7 (-3.8%) |
| CheXbert vector | 46.4 | 50.6 | – | 46.9 (-7.3%) |
| CheXpert-F1: | ||||
| Micro-F1-14 | – | 58.5 | 60.6 | 55.9 (-7.8%) |
| Macro-F1-14 | – | 42.7 | 40.0 | 40.4 (-5.4%) |
| Micro-F1-5 | – | 58.9 | 61.8 | 60.1 (-2.8%) |
| Macro-F1-5 | – | 51.5 | 49.5 | 53.8 (4.5%) |
As shown in Table 14, these models belong to the category of radiology foundation models. MedVersa (Zhou et al., 2024) and M4CXR (Park et al., 2024) support multiple tasks, including medical report generation, visual grounding, and visual question answering. In contrast, MAIRA-2 (Bannur et al., 2024) specialises in grounded radiology report generation, a novel task requiring image-level localisation for each identifiable finding or symptom. It is worth noting that the test sets differ slightly. All of these models utilise additional radiology information, such as lateral views, prior study reports of CXRs, or both.
Despite these considerations, Libra achieves the highest scores on most lexical metrics, including BLEU-, METEOR, and BERTScore, while trailing slightly behind MAIRA-2 on ROUGE-L. In clinical metrics, Libra predominantly ranks second, just behind the best-performing model. For clinical metrics, Libra consistently ranks second, just behind the top-performing model. In metrics that evaluate medical entities and their relationships, such as RadGraph-F1, RG, and RadCliQ, Libra also ranks second. Similarly, Libra comes second in the CheXbert vector embedding score. However, in the CheXpert metrics, Libra ranks first in Macro-F1 for the 5-class subset, with only a slight dip in the Micro-F1 score for the 14-class subset.
Incorporating lateral images and prior study reports could enhance clinical scores. Additionally, strategies like chain-of-thought reasoning and grounded report generation further improve performance in RRG tasks. Looking ahead, we plan to develop model architectures that can automatically adapt to multiple tasks and diverse scenarios, enabling more efficient handling of additional radiological information.
D.2 Discussion on Performance with non-LLM-based Models
| Model | Lexical Metrics | Clinical Metrics | |||||||
| B-1 | B-2 | B-3 | B-4 | MTR | R-L | P | R | F1 | |
| R2Gen | 35.3 | 21.8 | 14.5 | 10.3 | 14.2 | 27.0 | 33.3 | 27.3 | 27.6 |
| R2GenCMN | 35.3 | 21.8 | 14.8 | 10.6 | 14.2 | 27.8 | 34.4 | 27.5 | 27.8 |
| M2TR | 37.8 | 23.2 | 15.4 | 10.7 | 14.5 | 27.2 | 24.0 | 42.8 | 30.8 |
| KnowMAT | 36.3 | 22.8 | 15.6 | 11.5 | – | 28.4 | 45.8 | 34.8 | 37.1 |
| CMM-RL | 38.1 | 23.2 | 15.5 | 10.9 | 15.1 | 28.7 | 34.2 | 29.4 | 29.2 |
| CMCA | 36.0 | 22.7 | 15.6 | 11.7 | 14.8 | 28.7 | 44.4 | 29.7 | 35.6 |
| KiUT | 39.3 | 24.3 | 15.9 | 11.3 | 16.0 | 28.5 | 37.1 | 31.8 | 32.1 |
| DCL | – | – | – | 10.9 | 15.0 | 28.4 | 47.1 | 35.2 | 37.3 |
| METrans | 25.0 | 16.9 | 12.4 | 15.2 | – | 29.1 | 36.4 | 30.9 | 31.1 |
| ORGAN | 38.6 | 25.6 | 17.2 | 12.3 | 16.2 | 29.3 | 41.6 | 41.8 | 38.5 |
| COMG | 36.3 | 23.5 | 16.7 | 12.4 | 12.8 | 29.0 | – | – | – |
| MedM2G | 41.2 | 26.9 | 17.9 | 14.2 | – | 30.9 | – | – | – |
| BioViL-T | – | – | – | 9.2 | – | 29.6 | – | – | 17.5 |
| TiBiX | 32.4 | 23.4 | 18.5 | 15.7 | 16.2 | 33.1 | 30.0 | 22.4 | 25.0 |
| RECAP | 42.9 | 26.7 | 17.7 | 12.5 | 16.8 | 28.8 | 38.9 | 44.3 | 39.3 |
| Libra | 51.3 | 38.0 | 30.0 | 24.5 | 48.9 | 36.7 | 59.7 | 52.5 | 55.9 |
To facilitate comparison with non-LLM-based models, we selected evaluation metrics commonly used in these studies. These include BLEU- (Papineni et al., 2002), METEOR (MTR) (Banerjee & Lavie, 2005), and ROUGE-L (R-L) (Lin, 2004). For clinical metrics, we adopted the CheXbert (Irvin et al., 2019a), reporting Precision (P), Recall (R), and F1.
Baseline For performance evaluation, we compare our model with the following baselines: R2Gen (Chen et al., 2021), R2GenCMN (Chen et al., 2021), M2TR (Nooralahzadeh et al., 2021), KnowMAT (Yang et al., 2022), CMM-RL (Qin & Song, 2022), CMCA (Song et al., 2022), KiUT (Huang et al., 2023), DCL (Li et al., 2023b), METrans (Wang et al., 2023b), ORGAN (Hou et al., 2023b), COMG (Gu et al., 2023), MedM2G (Zhan et al., 2024), BioViL-T (Bannur et al., 2023), TiBiX (Sanjeev et al., 2024), RECAP (Hou et al., 2023a).
As the results demonstrate, our model, like other LLM-based models, generally outperforms non-LLM-based models. This is attributed to advancements in LLMs and visual instruction tuning (Liu et al., 2023), enabling multimodal large language models (MLLMs) to excel in the RRG task.
Appendix E Appendix E
E.1 Temporal Alignment Heatmap Analysis and Feature Representation
Current image
RAD-DINO layer 1
RAD-DINO layer 12
LFE
TAC
The heatmap in Figure 4 corresponds to the example in Figure 2, where no true prior image was used as a reference. It highlights the clear differences in feature representations across layers of the RAD-DINO image encoder. The shallow layers primarily capture the overall lung structure, while the deeper layers focus on specific disease locations.
After being processed by the Layerwise Feature Extractor (LFE), the image feature representations show that larger symptom regions are assigned higher weights. This indicates that the image representation achieves a more fine-grained level of detail. Following the Temporal Alignment Connector (TAC), Libra integrates the weighted dummy prior image, resulting in a more uniform feature distribution. This reflects temporal information, indicating no significant changes compared to the prior study. This process helps the model provide a smoother image feature representation for subsequent processing by the text decoder (i.e., the LLM).
Current image
Current LFE
TAC-1
Prior LFE
Prior image
Swap order
Current image (Prior)
Current LFE (Prior)
TAC-2
Prior LFE (Current)
Prior image (Current)
The heatmap in Figure 5 corresponds to the example in Figure 3, where a true prior image is provided. As observed, after processing through the LFE, the model captures higher-granularity feature representations in the symptom areas. After being processed by the TAC, the intermediate representation not only integrates these fine-grained features but also incorporates the difference information between the two images, reflecting the temporal information provided by the reference image, as shown in TAC-1 (Figure 5).
When the image order is swapped, treating the true prior image as the current image, the intermediate output from the LFE remains unchanged. Nevertheless, a comparison of the TAC-2 and TAC-1 outputs reveals noticeable differences in the lung feature representations. This demonstrates that the model’s temporal perception is directional and that the TAC module successfully integrates temporal information from different time points. At the same time, it confirms that the LFE focuses solely on extracting image feature representations without encoding temporal aspects.
This result is expected because the residual connections in the TAC ensure that the current image remains the main modality, while the prior image serves as the auxiliary modality. When the order is swapped, the prior image becomes the main modality for subsequent processing. Thus, swapping the image order also changes the content of the generated report, as mentioned in Section 5. The primary difference lies in the reversal of the temporal state of symptoms, such as changing from “improving” to “worsening”.

### v2

@@
 Abstract
-Radiology report generation (RRG) is a challenging task, as it requires a thorough understanding of medical images, integration of multiple temporal inputs, and accurate report generation.
-Effective interpretation of medical images, such as chest X-rays (CXRs), demands sophisticated visual-language reasoning to map visual findings to structured reports.
-Recent studies have shown that multimodal large language models (MLLMs) can acquire multimodal capabilities by aligning with pre-trained vision encoders.
-However, current approaches predominantly focus on single-image analysis or utilise rule-based symbolic processing to handle multiple images, thereby overlooking the essential temporal information derived from comparing current images with prior ones.
-To overcome this critical limitation, we introduce Libra, a temporal-aware MLLM tailored for CXR report generation using temporal images.
-Libra integrates a radiology-specific image encoder with a MLLM and utilises a novel Temporal Alignment Connector to capture and synthesise temporal information of images across different time points with unprecedented precision.
-Extensive experiments show that Libra achieves new state-of-the-art performance among the same parameter scale MLLMs for RRG tasks on the MIMIC-CXR.
-Specifically, Libra improves the RadCliQ metric by 12.9% and makes substantial gains across all lexical metrics compared to previous models111Our code is open-sourced at https://github.com/X-iZhang/Libra..
+Radiology report generation (RRG) requires advanced medical image analysis, effective temporal reasoning, and accurate text generation.
+While multimodal large language models (MLLMs) align with pre-trained vision encoders to enhance visual-language understanding, most existing methods rely on single-image analysis or rule-based heuristics to process multiple images, failing to fully leverage temporal information in multi-modal medical datasets.
+In this paper, we introduce Libra, a temporal-aware MLLM tailored for chest X-ray report generation.
+Libra combines a radiology-specific image encoder with a novel Temporal Alignment Connector (TAC), designed to accurately capture and integrate temporal differences between paired current and prior images.
+Extensive experiments on the MIMIC-CXR dataset demonstrate that Libra establishes a new state-of-the-art benchmark among similarly scaled MLLMs, setting new standards in both clinical relevance and lexical accuracy.
+Libra: Leveraging Temporal Images for Biomedical Radiology Analysis
+Xi Zhang1,2, Zaiqiao Meng1,2, Jake Lever1,2 & Edmond S. L. Ho
+1Information Retrieval Group, 2AI4BioMed Lab
+School of Computing Science
+University of Glasgow
+{X.Zhang.6}@research.gla.ac.uk
+{Zaiqiao.Meng,Jake.Lever,Shu-Lim.Ho}@glasgow.ac.uk
+https://x-izhang.github.io/Libra
 1 Introduction
-Radiology reports are a key component of biomedical radiology analysis, designed to enhance clarity and efficiency in medical communication.
-These structured reports are typically divided into distinct sections, such as Findings, Impression, Indication, Technique, Comparison, and History (Ganeshan et al., 2018).
-Serving as the primary medium for radiologists to convey their findings and conclusions from imaging studies like chest X-rays (CXRs), radiology reports play a crucial role in guiding diagnostic and therapeutic decisions for various diseases (Najjar, 2023).
-However, generating these reports manually is a complex and time-consuming task.
-Automating radiology report generation (RRG) presents a valuable opportunity to boost radiologist productivity, improve communication, and alleviate burnout (Zhang et al., 2020b).
-Despite its potential, RRG is a challenging task due to the intricate nature of medical imaging and the need for precise and accurate documentation.
-Although the advancements in paradigm models like LLaVA (Liu et al., 2023) and InstructBLIP (Dai et al., 2023), they still exhibit a significant performance gap when applied to RRG tasks.
-In particular, their effectiveness diminishes in biomedical contexts due to the substantial differences between biomedical and general image-text pairs (Tu et al., 2023; Saab et al., 2024).
-Consequently, these models often perform at a surface level, akin to a layperson-level understanding, when dealing with specialised medical imaging tasks.
-Recent advancements have explored continued pre-training of general-purpose foundation models for medical tasks, but these models still fall short in addressing the complexities of medical image analysis (Li et al., 2023a; Chaves et al., 2024; Park et al., 2024).
-Harnessing the full potential of multimodal large language models (MLLMs) for understanding and reasoning over multimodal data in the biomedical domain poses a significant challenge, primarily due to the granularity and specificity required in medical image analysis (Wang et al., 2023a).
-Preliminary attempts to apply MLLMs to RRG have largely concentrated on generating reports from a single image, often neglecting the crucial temporal relationships between images taken at different time points (Zhang et al., 2024b).
-The MIMIC-CXR Database (Johnson et al., 2019a) reveals that 67% of patients underwent at least two studies at different time intervals, underscoring the necessity of incorporating temporal context in clinical practice.
-This highlights a critical gap in existing models, which are not equipped to handle the dynamic progression of medical conditions reflected in sequential imaging studies.
-In standard clinical practice, radiologists rely heavily on comparing current imaging results with previous studies to assess temporal changes.
-However, many models inadvertently introduce spurious references to prior examinations that do not exist during inference, leading to inaccurate reports.
-Recent advancements222Detailed related work is discussed in Appx.
-A.1, and our research objectives are explained in Appx.
-A.2., such as MedVersa (Zhou et al., 2024) and MAIRA-2 (Bannur et al., 2024), can process multiple images simultaneously.
-These models insert visual tokens from different images at specific points within the textual input, hinging on the LLM to interpret reference information.
-Additionally, LLaVA-Rad (Chaves et al., 2024) utilised GPT-4V (OpenAI et al., 2024) for radiology report extraction and augmentation, eliminating hallucinations associated with prior image references in the dataset.
-Despite these efforts, existing models still lack integrated mechanisms for perceiving temporal information within their architectures, which restricts the model’s ability to achieve the temporal awareness required for comprehensive reporting.
-Meanwhile, existing MLLMs typically utilise visual embeddings from the last or penultimate layer of the image encoder (Chen et al., 2023a; Zhang et al., 2024a), which can only capture the global characteristics of the image.
-However, RRG tasks require not only identifying findings but also capturing subtle details such as severity, extent, and the progression of findings (Sloan et al., 2024).
-Such fine-grained details are difficult to represent using embeddings from any single layer of the image encoder (Jiang et al., 2024).
-Therefore, for the high-granularity demands of RRG tasks, relying solely on representation from a single layer limits the quality of report generation.
-To tackle these limitations, we aim to enhance the temporal awareness of MLLMs for RRG task by tackling two main challenges:
-It is non-trivial to design effective structures in MLLMs to handle prior study citations across various time points in the RRG task.
-The scarcity of effective connectors in MLLMs capable of handling the high-granularity requirements of downstream tasks.
-To overcome these challenges, we propose Libra (Leveraging Temporal Images for Biomedical Radiology Analysis), a novel framework designed to incorporate temporal change information into the RRG task.
-Libra employs a pre-trained visual transformer, RAD-DINO (Pérez-García et al., 2024), as the image encoder to generate robust image features, which are then refined using a new adapter specifically designed for the temporal awareness, before being fed into the medical large language model (LLM), Meditron (Chen et al., 2023b).
-Through a two-stage training strategy, Libra demonstrates the potential of MLLMs for specialised radiology applications.
-Our modular approach integrates state-of-the-art open-source pre-trained models for image and text while using a temporal-aware adapter to align these modalities with the text embedding space.
-Extensive experiments show that Libra substantially outperforms existing MLLMs in RRG tasks, demonstrating superior cross-modal understanding and reasoning capabilities.
-Specifically, our contributions include:
-We present Libra, the temporal-aware MLLM capable of capturing temporal changes and overcoming the challenge of handling prior study citations, setting a new state-of-the-art performance in RRG tasks on the MIMIC-CXR dataset among MLLMs of the same parameter scale.
-We designed the Temporal Alignment Connector (TAC) with two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image feature embedding from the encoder, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies.
-This enhances Libra’s ability to capture and utilise temporal information in the RRG task.
-2 Libra: Being Aware of Temporal Changes in Chest X-rays Report Generation
+Radiology reports are critical for biomedical radiology analysis, offering structured summaries of imaging studies such as chest X-rays (CXRs).
+Commonly divided into sections like Findings, Impression, Indication, Technique, Comparison, and History (Ganeshan et al., 2018), these reports guide diagnostic and therapeutic decisions (Najjar, 2023).
+However, manually generating such reports is both complex and time-consuming.
+Automating radiology report generation (RRG) holds great promise for alleviating radiologist burnout, increasing efficiency, and improving communication (Zhang et al., 2020b).
+Despite this, the intricate nature of medical imaging demands precise and detailed documentation, making RRG a challenging task.
+Recent advances in Multimodal Large Language Models (MLLMs), such as LLaVA (Liu et al., 2023) and InstructBLIP (Dai et al., 2023), have demonstrated potential in vision-language tasks.
+However, their performance diminishes in specialised biomedical contexts due to the significant domain shift between general-purpose and medical image-text data (Tu et al., 2023).
+These models often lack fine-grained detail for medical imaging tasks, resulting in surface-level understanding akin to layperson interpretations.
+While continued pre-training on medical datasets and domain-specific fine-tuning (e.g., LLaVA-Med (Li et al., 2023a), MedBLIP (Chen et al., 2023b)) improve performance, they still cannot fully capture the complexities of medical image analysis (Xiao et al., 2024).
+One critical gap in the current MLLM-based approaches is their limited ability to incorporate temporal context, which is pivotal in clinical practice.
+Radiologists routinely compare current imaging results with prior studies to identify temporal changes, a process crucial for understanding disease progression and guiding treatment decisions.
+Indeed, the MIMIC-CXR database (Johnson et al., 2019b) reveals that 67% of patients underwent at least two studies at different time intervals, underscoring the necessity of temporal reasoning.
+However, most MLLMs designed for RRG tasks focus on single-image analysis, neglecting this temporal dimension (Zhang et al., 2024c).
+As illustrated in Figure 1, MAIRA-1 (Hyland et al., 2024) introduces hallucinated prior references in single-image analysis and misinterprets temporal changes when integrating prior studies.
+Although recent models111Detailed related work is discussed in Appx.
+A, and our research objectives are explained in Appx.
+B.
+like MedVersa (Zhou et al., 2024) and MAIRA-2 (Bannur et al., 2024) have introduced multi-image processing, they do not explicitly model or extract temporal differences.
+Instead, they rely on inserting visual tokens from different studies at specific points within textual inputs and delegate the reasoning task to the LLM.
+Similarly, Banerjee et al.
+(2024) and Chaves et al.
+(2024) leverage GPT-4V (OpenAI et al., 2024) to eliminate hallucinated references to prior studies in the dataset but lacks dedicated mechanisms for modelling temporal progression.
+Additionally, existing MLLMs often rely on embeddings from the last or penultimate layer of the image encoder (Chen et al., 2023a; Zhang et al., 2024a), primarily capturing global features.
+However, RRG tasks require fine-grained details222E.g., severity and temporal progression of findings.
+(Sloan et al., 2024), which a single-layer embedding often cannot fully represent (Jiang et al., 2024).
+To tackle these limitations, we enhance MLLM temporal awareness for RRG tasks by addressing two main challenges:
+Designing robust MLLM architectures that seamlessly handle prior study references in RRG.
+The scarcity of effective feature alignment projectors in MLLMs capable of handling the high-granularity requirements of downstream tasks.
+To overcome these gaps, we propose Libra (Leveraging Temporal Images for Biomedical Radiology Analysis), a novel temporal-aware framework tailored for RRG tasks.
+Libra employs a pre-trained visual transformer encoder, RAD-DINO (Pérez-García et al., 2024), to generate robust image features, which are then refined using a new projector crafted for the temporal awareness, before being fed into the medical large language model (LLM), Meditron (Chen et al., 2023c).
+Through a two-stage training strategy, Libra aligns temporal visual features with the text embedding space, improving temporal coherence in RRG.
+Our modular approach integrates state-of-the-art open-source pre-trained models for medical image and text processing while introducing a dedicated temporal-aware adapter to align visual and textual modalities within the embedding space.
+This paper makes the following contributions:
+Libra, a temporal-aware MLLM designed to model temporal references and mitigate temporal hallucinations in RRG tasks.
+Temporal Alignment Connector (TAC), comprising the Layerwise Feature Extractor (LFE) and Temporal Fusion Module (TFM), which extracts high-granularity image features from multiple encoder layers and integrates temporal references from the prior study when available.
+Extensive evaluation on the MIMIC-CXR dataset, achieving state-of-the-art results on average among similarly scaled MLLMs, with case analysis illustrating Libra’s architectural benefits.
+2 Libra
 2.1 Model Architecture
-Libra model follows the common architecture of MLLMs, such as LLaVA, which includes an image encoder, a text decoder (i.e., an LLM pre-trained based on text-only data) and an adapter module for mapping visual representations into text space.
-In this work, we utilise a frozen biomedical image encoder, RAD-DINO (Pérez-García et al., 2024), a visual transformer extensively pre-trained on medical scans using the DINOv2 image-only self-supervised learning approach (Oquab et al., 2024).
-The LLM is deployed by Meditron-7B (Chen et al., 2023b), which builds on Llama-2 (Touvron et al., 2023) and further pre-trained on a specialised medical corpus.
-However, for the adapter module, we specifically designed a Temporal Alignment Connector (TAC) to capture and integrate temporal information across two images from different time points, as shown in Figure 1.
-The TAC bridges the image encoder and LLM.
-It consists of two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image representations, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies.
-This design enables Libra to effectively manage temporal data and enhances its ability to generate accurate and coherent radiology reports.
-2.2 Temporal Alignment Connector (TAC): New Adapter Design
-To address the challenges of integrating temporal information and aligning high-granularity image features in RRG tasks, we developed the TAC module.
-It bridges the gap between visual features extracted from multiple temporal snapshots of a patient’s CXRs and the language model.
-The TAC consists of two sub-modules: the LFE, which is to extract the layerwise representation of images, and the TFM, which is to capture the temporal-aware representation between two images.
-2.2.1 Layerwise Feature Extractor (LFE)
-To leverage abundant feature information in a pre-trained image encoder, we extract all image patch token features from each hidden layer for a given input image.
-Specifically, RAD-DINO has 12 hidden layers and processes 518 × 518 images into 14 × 14 patches, generating 1,369 patch token sequences per hidden layer for each image.
-We do not use the token.
-Following this, we obtain patch embeddings of the same dimension from each layer, denoted as , where is the number of patch tokens and is the embedding dimension of the image encoder.
-Then, we combine them as , where is the number of hidden layers.
-Drawing from VGG (Simonyan & Zisserman, 2015), we employ a progressive compression strategy to reduce the layer dimension, ensuring that the LFE captures the most relevant features.
-Initially, we utilise Squeeze-and-Excitation (SE) Networks (Hu et al., 2019), which construct informative features by integrating both spatial and channel-wise information within local receptive fields at each layer.
-The SE block is applied to obtain calibrated feature representations, using GELU (Hendrycks & Gimpel, 2023) as the activation function.
-Next, we employ a specialised pointwise convolution module to align the feature spaces across different layers, using a depthwise 2D convolution with filters and stride of , without bias.
-The dimensions are gradually reduced to better align with the text in the LLM.
+Our Libra model follows the standard architecture of MLLMs, such as LLaVA (Liu et al., 2023), comprising an image encoder, a text decoder and a connector module to map visual features into the text embedding space.
+Figure 2 shows the overall architecture of Libra.
+Specifically, we utilise a frozen biomedical image encoder, i.e.
+RAD-DINO (Pérez-García et al., 2024), a visual transformer extensively pre-trained on medical scans using the DINOv2 image-only self-supervised learning approach (Oquab et al., 2024).
+The text encoder is deployed by Meditron-7B (Chen et al., 2023c), which builds on Llama-2 and is further pre-trained on specialised medical corpora.
+To effectively connect the image encoder and LLM, we design a novel Temporal Alignment Connector (TAC) tailored to capture and integrate temporal information from paired images taken at different time points.
+Meanwhile, when no prior image is available, we employ a dummy prior image, which is simply a copy of the current image, to mitigate spurious references to nonexistent scans, as shown in Figure 2 (bottom).
+This design enables Libra to effectively manage temporal data (e.g., stable, improved, worsening) and enhances its ability to generate accurate and coherent radiology reports.
+2.2 Temporal Alignment Connector
+To address the challenges of integrating temporal information and aligning high-granularity visual features for RRG tasks, TAC bridges the image encoder and the LLM.
+It processes visual features from two temporal snapshots to produce a unified representation sensitive to temporal changes.
+As shown in Figure 2 (right), TAC includes two key components: the Layerwise Feature Extractor, which extracts high-granularity image representations, and the Temporal Fusion Module, which integrates temporal references from the prior study.
+2.2.1 Layerwise Feature Extractor
+To leverage abundant image feature representations encoded by a pre-trained image encoder, we extract image patch token features of all the hidden layers for a given pair of input images.
+By default, the RAD-DINO image encoder (Pérez-García et al., 2024) has 12 hidden layers and processes 518 × 518 input images into 14 × 14 patches, generating 1,369 patch token sequences per hidden layer.
+Rather than relying on a single global feature token (e.g., the token), we collect same-dimensional patch embeddings from each layer per image, denoted as , where is the number of patch tokens and is the embedding dimension of the image encoder.
+Then, these embeddings are concatenated across all layers as , where is the number of hidden layers.
+Drawing from the progressive compression strategy in VGG (Simonyan and Zisserman, 2015), our Layerwise Feature Extractor (LFE) reduces dimensionality across layers while preserving critical information.
+First, we utilise Squeeze-and-Excitation (SE) Networks (Hu et al., 2019), which construct informative features by integrating both spatial and channel-wise information within local receptive fields at each layer.
+The SE block is applied to obtain calibrated feature representations, using GELU (Hendrycks and Gimpel, 2023) as the activation function.
+Next, we employ a specialised pointwise convolution module to align the feature spaces across different layers, using a depthwise 2D convolution with filters and stride of 1, without bias.
 The compressed features are represented as , where is the original layer number and is the layer number after compression.
-Following the size-reduction pattern of convolutional layers in VGG, the image representations are compressed according to 333Since RAD-DINO has 12 hidden layers, the prime factorisation chain provides the factors as ..
-Through three stages of progressive compression, we obtain the final representation of image patch features.
-We define the function to project image features into the same dimension:
-where , . The LFE independently processes each image to produce a unified representation for temporal alignment, as (a) in Figure 1.
-2.2.2 Temporal Fusion Module (TFM)
-The fusion module is inspired by the transformer decoder.
-It utilises prior images as auxiliary information to weight the current image, generating a refined, temporal-aware representation suitable for the LLM’s latent space.
-The TFM uses an attention mechanism to dynamically learn relationships between image pairs, enabling the model to adapt to various temporal changes.
-Prior Image Prefix Bias: The dataset contains samples with and without a prior image.
-The patch token features of the current and previous images processed by the LFE are denoted as and .
-If no true prior image exists, we set , effectively using the current image as a dummy prior.
-To differentiate this case, we introduce a trainable bias .
-Following the attention mechanism’s scaling techniques for adjusting hidden space degrees of freedom with a chi-square distribution (Vaswani, 2017; Wang et al., 2018a), we apply an exponent of to the cosine similarity, where is the hidden dimension of the LLM, to obtain a weighted bias addition to :
-This nonlinear scaling amplifies higher similarity values, effectively modulating the influence of prior image features.
-When no true prior image is available, the high similarity score ensures that the effect of the dummy prior is adequately represented.
-This adjustment prevents samples with a dummy prior image from undergoing redundant rounds of parallel multi-head self-attention during subsequent propagation through the transformer blocks, as shown in Figure 1.
-Transformer Block: Our transformer block includes a multi-head cross-attention sub-layer, multi-head self-attention sub-layers, and two multi-layer perceptron (MLP) sub-layers.
-As shown in Figure 1 (b).
-The paired are processed with layer normalization and residual connections:
+Following the size-reduction pattern of convolutional layers in VGG, the image features are compressed according to 333Since RAD-DINO has hidden layers, the prime factorisation chain provides the factors as ..
+Through three stages of progressive compression, we obtain the final patch-level representation:
+For simplicity, we use to denote the above three stages of compression, which project a given input image into its feature representation of the fixed dimension, :
+By progressively refining each image’s representations through multiple stages, the LFE generates a unified and compact feature set suitable for temporal alignment.
+This design ensures that both high-granularity and global context are retained, as illustrated in (a) of Figure 2.
+2.2.2 Temporal Fusion Module
+The Temporal Fusion Module (TFM) is inspired by the transformer decoder and is designed to integrate temporal information by leveraging prior images as auxiliary context.
+It takes as input a paired set of compressed features from both the current and prior images, denoted as and , respectively, which are obtained after processing through the LFE.
+The temporal fusion process is defined as:
+where TFM learns to weigh the current image using prior image features, refining the representation to enhance temporal awareness.
+The resulting feature sequence, , serves as the input to the LLM, where is the number of patch tokens and is the hidden dimension of the LLM.
+This process encapsulates the temporal evolution of the patient’s condition, allowing the language model to generate accurate and contextually aware radiology reports.
+Prior Image Prefix Bias
+The dataset contains samples with and without a prior image.
+When a prior image is not available, we set .
+However, this “dummy prior image” is indistinguishable from a true prior in raw features.
+To differentiate it, we add a trainable bias, as .
+Following the attention scaling techniques for adjusting hidden space degrees of freedom with a chi-square distribution (Vaswani, 2017), a nonlinear scaling function amplifies higher similarity values.
+The cosine similarity between the current and prior images is scaled with an exponent of , where is the hidden dimension of the LLM:
+This nonlinear scaling emphasises high similarity values, modulating the influence of prior image features.
+When no prior image is available, the high similarity score ensures that the effect of the dummy prior is adequately represented.
+This adjustment prevents samples with a dummy prior image from undergoing redundant rounds of parallel multi-head self-attention during subsequent propagation through the transformer blocks, in Figure 2.
+Transformer Block
+The Transformer Block in TFM follows the standard Transformer design but is optimized for handling temporal image pairs.
+It consists of multi-head self-attention (), multi-head cross-attention (), and two multi-layer perceptron () sub-layers.
+As illustrated in (b) of Figure 2.
+The paired are processed with layer normalization () and residual connections:
 where is a simple neural network composed of two fully connected layers with GELU as the activation function.
 After that, the features are processed through , a straightforward neural network consisting of four fully connected layers with the same activation function, but with hidden dimensions matching those of the LLM.
-We define our fusion module function as .
-Subsequently, becomes the input sequence that LLM can comprehend, where is the number of patch tokens and is the hidden dimension of LLM.
-The final output is a refined feature representation that encapsulates the temporal evolution of the patient’s condition, making it suitable for the language model to generate accurate and contextually aware radiology reports.
 2.3 Prompt Design
-To enhance Libra’s ability to perceive temporal changes and integrate medical information in the RRG task, we designed a comprehensive prompting strategy comprising a system prompt and detailed clinical prompts, as shown in Figure 1.
-The system prompt is crafted to enable the LLM to recognise temporal variations: “The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes.” Following this, additional sections of the report, such as {Indication}, {History}, {Comparison}, and {Technique}, are incorporated into the clinical instructions (e.g., in Appx.
-B.4).
-The clinical prompt to Libra is: “Provide a detailed description of the findings in the radiology image.
-Following clinical context: .” If no clinical instructions are available, the prompt defaults to: “Provide a detailed description of the findings in the radiology image.” After tokenizing and embedding the prompt and the answer, we insert the image patch tokens at the specified location, typically between the system prompt and the clinical prompts.
+To enhance Libra’s ability to perceive temporal changes and integrate medical information in RRG, we design a structured prompting strategy, consisting of a system prompt and a clinical prompt, as shown in Figure 2 (left).
+The system prompt enables the LLM to recognise temporal variations, while standard report sections (Indication History, Comparison, Technique) are integrated into the clinical prompt (see Appx.
+C for a detailed example).
+The full prompt is: “Provide a detailed description of the findings in the radiology image.
+Following clinical context:.” There are datasets, e.g.
+MIMIC-CXR (Johnson et al., 2019b), where the report sections are unavailable.
+For these datasets, we set the prompt as follows: “Provide a detailed description of the findings in the radiology image.” After tokenising and embedding prompts, the refined image features () are inserted between the system prompt and clinical prompts.
 2.4 Temporal-aware Training
-In this study, we focus on frontal-view images, either posterior-anterior (PA) or anterior-posterior (AP), and the Findings sections of radiology reports, as they contain the most direct clinical observations.
-We adopt a two-stage training strategy inspired by recent advancements in MLLM fine-tuning techniques (McKinzie et al., 2024).
-This proposed approach aims to progressively teach the model essential skills, including visual feature alignment and temporal information extraction.
-Libra’s training process involves two stages: temporal feature alignment and downstream task fine-tuning.
-In the first stage, the visual encoder and LLM weights are frozen, and the TAC is trained with Findings and Impression generation, along with CXR-related VQA tasks to extract high-quality image representations and capture temporal changes.
-In the second stage, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM on the Findings section generation task, keeping the visual encoder and connector weights frozen.
-Unlike traditional full fine-tuning, LoRA achieves comparable performance with significantly lower training costs.
-The detailed training configuration, including learning rate schedules and model parameters, is provided in Appx.
-B.1.
+Libra focuses on frontal-view images, either posterior-anterior (PA) or anterior-posterior (AP), and targets the Findings sections of RRG, as these contain the most direct clinical observations.
+It employs a two-stage training strategy, inspired by recent MLLM fine-tuning techniques (McKinzie et al., 2024), to progressively learn visual feature alignment and temporal information extraction.
+Temporal Feature Alignment
+In the first stage, the visual encoder and LLM weights are frozen, while the TAC is trained.
+This stage focuses on Findings and Impression generation from paired images and performing CXR-related visual question answering (VQA) tasks to extract high-quality image representations and capture temporal changes.
+Downstream Task Fine-tuning
+In the second stage, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the LLM on the Findings generation task, while keeping the visual encoder and TAC weights frozen.
+LoRA achieves performance comparable to full fine-tuning at a substantially lower computational cost.
+The detailed training configuration, including learning rate schedules and hyperparameters, is provided in Appx.
+D.
+| Metric | LLaVA-Med | CheXagent | GPT-4V | Med-PaLM | LLaVA-Rad | MAIRA-1 | Libra (%) |
+| Lexical: | |||||||
+| ROUGE-L | 27.6 | 21.5 | 13.2 | 27.5 | 30.6 | 28.9 | 36.2 (+18.3%) |
+| BLEU-1 | 35.4 | 16.9 | 16.4 | 32.3 | 38.1 | 39.2 | 51.2 (+30.6%) |
+| BLEU-4 | 14.9 | 4.7 | 17.8 | 11.5 | 15.4 | 14.2 | 24.3 (+36.5%) |
+| METEOR | 35.3 | – | – | – | – | 33.3 | 48.7 (+38.0%) |
+| Clinical: | |||||||
+| RadGraph-F1 | 19.1 | – | – | 26.7 | – | 24.3 | 32.4 (+21.3%) |
+| RG | 23.8 | 20.5 | 13.2 | – | 29.4 | 29.6 | 36.9 (+25.0%) |
+| RadCliQ() | 3.30 | – | – | – | – | 3.10 | 2.76 (+11.0%) |
+| CheXbert vector | 36.9 | – | – | – | – | 44.0 | 46.3 (+5.2%) |
+| CheXpert-F1: | |||||||
+| Micro-F1-14 | 42.7 | 39.3 | 35.5 | 53.6 | 57.3 | 55.7 | 55.3 (-3.4%) |
+| Macro-F1-14 | 26.9 | 24.7 | 20.4 | 39.8 | 39.5 | 38.6 | 40.2 (+1.1%) |
+| Micro-F1-5 | 43.9 | 41.2 | 25.8 | 57.9 | 57.4 | 56.0 | 58.9 (+1.8%) |
+| Macro-F1-5 | 36.3 | 34.5 | 19.6 | 51.6 | 47.7 | 47.7 | 52.6 (+2.0%) |
 3 Experiments
 3.1 Task and Dataset
-Task Description We focus on generating the Findings section of radiology reports for a set of frontal CXRs to ensure fair comparison with previous work.
-The Findings section contains descriptions by radiologists of both normal and abnormal findings.
-Typically, these CXRs are accompanied by additional sections such as Indication and Technique, which provide the rationale for the study, including clinical history or specific requests from the referring physician.
-While not directly related to diagnostic interpretation, these sections serve as routine records and assist the model in better understanding temporal changes between images across different periods.
-Therefore, we incorporate clinical instructions about the current image as prompts to guide Libra to complete the RRG task.
+Task Description
+We focus on generating the Findings section of radiology reports for frontal CXRs, ensuring a fair comparison with prior work.
+The Findings section provides radiologists’ observations, encompassing both normal and abnormal findings.
+While additional sections like Indication and Technique primarily serve as routine records (e.g., clinical history or specific physician requests), they also assist the model in understanding temporal changes across images.
+Hence, we incorporate clinical instructions about the current image as prompts to guide Libra to complete the RRG task.
 The most common CXR is frontal views, either PA or AP.
-While other views, such as lateral images, are occasionally used for diagnostic purposes, they mainly serve as supplementary tools to aid in interpreting anatomical structures more comprehensively (Islam et al., 2023).
-Therefore, the frontal view remains the standard perspective for clinical interpretation of CXRs, and our study is consistent with prior work in RRG tasks, including studies by Chaves et al.
+Although lateral views are occasionally used to supplement anatomical assessments (Islam et al., 2023), they are excluded in this study to maintain consistency with previous research on RRG tasks, such as Chaves et al.
 (2024) and Hyland et al.
 (2024).
-For both the current and prior images, we exclusively use a single frontal view.
-Dataset Description To train Libra, we utilise the MIMIC-CXR dataset (Johnson et al., 2019b) and its derivative datasets, Medical-Diff-VQA (Hu et al., 2023) and MIMIC-Ext-MIMIC-CXR-VQA (Bae et al., 2023), as shown in the Table 1.
-All datasets are split according to the official labels to ensure no data leakage.
-Detailed descriptions of datasets can be found in the Appx.
-B.2.
-| Data source | Task Type | # Samples | % Has Prior | ||||
+Both current and prior images in our experiments exclusively utilise single frontal views.
+Dataset Description
+Libra is trained and evaluated using the MIMIC-CXR dataset (Johnson et al., 2019b) and its derivative datasets, including Medical-Diff-VQA (Hu et al., 2023) and MIMIC-Ext-MIMIC-CXR-VQA (Bae et al., 2023).
+All datasets are split according to the official labels to prevent data leakage.
+Detailed dataset descriptions and preprocessing steps are in Appx.
+E.
+Following the dataset scaling law utilised in multi-stage MLLM fine-tuning methods (Zhu et al., 2023), we adopt a two-stage training strategy, as noted in Sec.
+2.4.
+The first stage trains TAC on 1.2M CXR-image text pairs from MIMIC-CXR and its derivatives, including Findings, Impression, and VQA tasks, enabling it to learn CXR token distributions and image-text relationships.
+The second stage fine-tunes the model on downstream tasks, refining the LLM to align high-granularity CXR features with the Findings section of reports.
+Beyond Findings section generation, the first stage incorporates Impression section and VQA tasks.
+The Impression section, which summarises diagnoses and proposes further investigations (Babar et al., 2021), facilitates alignment between CXRs and their textual descriptions.
+We use the same system and clinical prompts as for Findings, replacing ‘Findings’ with ‘Impression’.
+For VQA, the system prompts remain unchanged, while clinical prompts are adapted to address medical-specific questions, guiding caption generation.
+These VQA tasks refine the MLLM’s biomedical vocabulary usage and strengthen image-text alignment.
+3.2 Evaluation Metrics
+We evaluate the generated reports using lexical and radiology-specific metrics, adhering to established protocols.
+Lexical metrics include ROUGE-L (Lin, 2004), BLEU- (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005) and BERT (Devlin et al., 2019).
+Radiology-specific metrics include RadGraph-F1 (Jain et al., 2021), RG (Delbrouck et al., 2022a), F1-CheXpert (Irvin et al., 2019), CheXbert vector similarity (Smit et al., 2020a), and RadCliQ (Yu et al., 2022) version .
+These clinical metrics typically emphasise the accuracy of medical findings, prioritising the detection of clinically relevant entities.
+However, they do not evaluate the model’s ability to capture temporal information.
+Therefore, we introduce the temporal entity F1 score () to assess this aspect.
+In particular, the temporal entity F1 score specifically measures the accuracy of entities related to progression over time described in the report 444Full metric descriptions, including , are in Appx.
+F..
+Temporal Entity F1
+Building on the work of Bannur et al.
+(2023), we set a reward list comprising common radiology-related keywords indicative of temporal changes.
+Temporal entities are then extracted from both the ground truth and the generated reports without applying stemming or lemmatization, preserving the precision of temporal descriptions.
+After extraction, we compute precision and recall , which are subsequently used to calculate the , defined as the harmonic mean of precision and recall (Van Rijsbergen, 1974), also known as the score.
+where is a small value, set to a default of , to prevent division by zero (it is also added to the numerator for special cases where no temporal entities are present in the ground truth).
+3.3 Baselines
+While the MIMIC-CXR dataset provides an “official” test split, strict comparisons with prior studies are challenging due to differences in inclusion criteria and pre-processing steps.
+For instance, Yu et al.
+(2022) and Jeong et al.
+(2023) included only one image per study, resulting in a test set of 1,597 samples, while Tanida et al.
+(2023) followed the Chest ImaGenome split (Wu et al., 2021).
+Such variations in test set distributions can significantly impact the reported results (Park et al., 2024).
+To ensure fairness 555The test set includes 2,461 frontal-view samples., we use a widely adopted test set focused on frontal-view CXRs, aligned with previous studies such as MAIRA-1 (Hyland et al., 2024) and LLaVA-Rad (Chaves et al., 2024).
+Recent concurrent work, such as M4CXR (Park et al., 2024), employs multi-turn chain-of-thought prompting (Wei et al., 2023) for report generation, which differs from our task setup.
+Additionally, we do not compare with MAIRA-2 (Bannur et al., 2024), a model designed for grounded radiology report generation incorporating lateral views and prior study reports for each subject within the input prompt.
+Bannur et al.
+(2024) emphasises a positive transfer between this distinct task setup and standard RRG, which falls beyond our study’s scope.
+For comparison and discussion of the latest concurrent and non-LLM-based models, see Appx.
+G.
+Considering these factors, we compared our model with state-of-the-art models, including LLaVA-Med (Li et al., 2023a), CheXagent (Chen et al., 2024b), GPT-4V (OpenAI et al., 2024), Med-PaLM (Tu et al., 2023), LLaVA-Rad and MAIRA-1.
+Table 1 presents the results.
+As many of these models are not publicly available, we present their evaluation results as reported in the original sources.
+3.4 Results
+From Table 1, Libra666Libra was tested on single-image inputs without priors for fair comparison with models lacking temporal modelling.
+achieves competitive results across most traditional lexical and clinical metrics, excelling in ROUGE-L, BLEU, METEOR, and RadGraph-based scores.
+It also leads in the radiologist-aligned RadCliQ metric and CheXbert vector similarity.
+In the CheXpert classification, it attains the highest Macro-F1 scores and a competitive Micro-F1.
+Overall, Libra demonstrates robust performance in RRG by effectively leveraging temporal information, with only minor gaps in select clinical metrics.
+These results highlight the effectiveness of its TAC in capturing temporal contexts and generating clinically relevant radiology reports.
+| Metric | Libra-1 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
+| Lexical: | |||||
+| ROUGE-L | 27.56 | 27.33 (-0.85%) | 27.21 (-1.27%) | 27.43 (-0.48%) | 26.17 (-5.04%) |
+| BLEU-1 | 34.84 | 34.17 (-1.92%) | 34.21 (-1.82%) | 34.60 (-0.67%) | 33.03 (-5.20%) |
+| BLEU-4 | 11.51 | 11.13 (-3.33%) | 11.11 (-3.47%) | 11.43 (-0.73%) | 10.02 (-12.98%) |
+| METEOR | 35.50 | 35.06 (-1.24%) | 34.96 (-1.52%) | 35.28 (-0.62%) | 33.98 (-4.28%) |
+| BERTScore | 55.87 | 55.60 (-0.49%) | 55.49 (-0.69%) | 55.74 (-0.23%) | 54.63 (-2.22%) |
+| F1 | 26.63 | 25.96 (-2.51%) | 26.21 (-1.57%) | 26.58 (-0.18%) | 25.39 (-4.65%) |
+| Clinical: | |||||
+| RadGraph-F1 | 22.52 | 22.20 (-1.42%) | 22.03 (-2.19%) | 22.35 (-0.74%) | 21.51 (-4.48%) |
+| RG | 27.32 | 26.89 (-1.59%) | 26.72 (-2.19%) | 27.09 (-0.84%) | 25.97 (-4.96%) |
+| RadCliQ () | 3.10 | 3.12 (-0.65%) | 3.12 (-0.65%) | 3.11 (-0.32%) | 3.15 (-1.61%) |
+| CheXbert vector | 42.02 | 41.57 (-1.07%) | 41.37 (-1.54%) | 41.92 (-0.24%) | 40.93 (-2.59%) |
+| CheXpert-F1: | |||||
+| Micro-F1-14 | 52.48 | 51.74 (-1.42%) | 51.68 (-1.53%) | 52.13 (-0.67%) | 51.13 (-2.57%) |
+| Macro-F1-14 | 36.87 | 36.04 (-2.25%) | 36.12 (-2.03%) | 36.14 (-1.97%) | 35.85 (-2.76%) |
+| Micro-F1-5 | 56.63 | 55.37 (-2.23%) | 55.79 (-1.49%) | 55.87 (-1.34%) | 54.51 (-3.74%) |
+| Macro-F1-5 | 49.33 | 47.76 (-3.18%) | 47.82 (-3.06%) | 47.98 (-2.75%) | 47.22 (-4.28%) |
+4 Ablation Studies
+We conducted ablation studies on Libra’s key components, evaluating module and dataset configurations.
+All experiments were performed on the MIMIC-CXR test split for the Findings generation, with prior images included by default and consistent hyperparameters during training and inference.
+Does the Temporal Alignment Connector improve model performance?
+To evaluate the impact of TAC on Libra’s performance in RRG, we used a model initialised with the RAD-DINO (Pérez-García et al., 2024) image encoder, TAC, and Meditron-7b (Chen et al., 2023c) as the LLM.
+The baseline (Libra-1) was conducted by fine-tuning only the TAC for the Findings generation task.
+As shown in Table 2, we performed ablation studies by progressively removing different TAC components, including the Temporal Fusion Module (TFM), Layerwise Feature Extractor (LFE), Prior Image Prefix Bias (PIPB), and the entire TAC.
+Removing TFM restricted the model to single-image processing, akin to LLaVA (Liu et al., 2023), but with a four-layer MLP for aligning image features with the LLM’s hidden dimensions.
+Without LFE, the model used the penultimate layer of the encoder.
+Removing PIPB excluded the mechanism for differentiating true and dummy prior images.
+Finally, removing the entire TAC left the model reliant solely on the image encoder and LLM.
+The results indicate that removing any TAC submodule leads to performance declines across all metrics compared to Libra-1.
+TFM removal caused a notable drop in the F1 score (2%), highlighting its role in capturing temporal information.
+LFE removal significantly decreased RadGraph-related scores, underscoring its importance in extracting detailed image features.
+PIPB removal impacted clinical metrics more than lexical metrics, indicating its role in enhancing clinical relevance.
+Complete TAC removal led to substantial declines in all metrics, demonstrating its critical role in integrating image details and temporal information.
+The evaluation confirms that TAC plays a vital role in improving Libra’s ability to generate high-quality, temporally aware radiology reports.
+For additional ablation studies exploring TAC’s contributions, including its impact under general-domain and radiology-specific pre-trained models, its performance after the second training stage, and its robustness through extended fine-tuning and diverse conditions, and an analysis of whether incorporating temporal information improves Libra’s performance in RRG tasks, please refer to Appx.
+H.
+Are additional Impression and VQA datasets necessary during the feature alignment?
+To assess the impact of incorporating additional datasets during the first stage of training, we compared a model (Libra-f) trained solely on the Findings data with Libra, which also used Impression and VQA data for feature alignment, as shown in Table 3.
+After the first stage, Libra outperformed Libra-f in lexical metrics but showed a slight decline in clinical scores.
+This decline stems from VQA tasks emphasizing fine-grained, grounded descriptions rather than holistic findings.
+VQA focuses on individual symptoms, whereas Findings integrates multiple normal and abnormal observations, affecting F1 by reducing identified temporal entities.
+In the second stage, fine-tuning on Findings restored balance, further improving performance.
+These results indicate that additional datasets enhance Libra’s RRG ability, while second-stage fine-tuning ensures well-rounded report generation.
+| Metric | Stage: 1 | Stage: 2 | ||
+| Libra-f | Libra | Libra-f | Libra | |
+| Lexical: | ||||
+| ROUGE-L | 27.56 | 27.27▼ | 35.31 | 36.66△ |
+| BLEU-1 | 34.84 | 41.24△ | 49.92 | 51.25△ |
+| BLEU-4 | 11.51 | 13.59△ | 23.05 | 24.54△ |
+| METEOR | 35.50 | 39.44△ | 47.99 | 48.90△ |
+| BERTScore | 55.87 | 56.00△ | 61.28 | 62.50△ |
+| F1 | 26.63 | 24.80▼ | 33.52 | 35.34△ |
+| Clinical: | ||||
+| RadGraph-F1 | 22.52 | 20.45▼ | 30.77 | 32.87△ |
+| RG | 27.32 | 25.19▼ | 35.44 | 37.27△ |
+| RadCliQ () | 3.10 | 3.31▼ | 2.83 | 2.72△ |
+| CheXbert vector | 42.02 | 35.33▼ | 45.32 | 46.85△ |
+| CheXpert-F1: | ||||
+| Micro-F1-14 | 52.48 | 43.63▼ | 54.11 | 55.87△ |
+| Macro-F1-14 | 36.87 | 25.68▼ | 37.16 | 40.38△ |
+| Micro-F1-5 | 56.63 | 49.75▼ | 58.76 | 60.07△ |
+| Macro-F1-5 | 49.33 | 40.40▼ | 51.99 | 53.75△ |
+5 Performance Analysis
+We qualitatively assess Libra’s ability to generate temporally consistent radiology reports.
+Cases without Prior Image
+As shown in Figure 3 (a), Libra produced detailed descriptions beyond the ground truth, identifying “sternal wires” and their type.
+This demonstrates its capability to deliver clinically relevant information without spurious referencing nonexistent prior studies.
+Cases with Prior Image
+In Figure 3 (b), new abnormalities such as pleural effusion and pneumonia appeared in the current image.
+Without a prior image, Libra correctly described the present findings but did not infer disease progression, avoiding spurious references while still suggesting further investigations.
+When the prior image was considered, Libra effectively captured these progressive changes, provided detailed descriptions, and explicitly referenced the comparison.
+This facilitated a clear understanding of temporal changes and more accurate descriptions of disease progression.
+Evaluating Temporal Consistency
+To assess temporal reasoning, we swapped image order, using the prior image as the current image and vice versa.
+The generated report then reflected an improved patient condition, aligning with the reversed input sequence but contradicting the ground truth of the original current image.
+Notably, the report closely resembled the original description of the prior image, as shown at the bottom of Figure 3 (b).
+This indicates that Libra can effectively adapt to both temporal contexts, generating accurate and contextually consistent reports that simulate the conditions of standard clinical practice.
+6 Conclusion
+In this study, we introduced Libra, a temporal-aware multimodal large language model tailored for chest X-ray report generation tasks.
+Libra employs a two-stage training framework, leveraging a radiology-specific image encoder and language model connected via the Temporal Alignment Connector, enabling seamless integration of visual and textual modalities.
+Trained solely on the open-access MIMIC-CXR dataset (Johnson et al., 2019b), Libra demonstrates notable performance gains across key metrics compared to similarly scaled models.
+Through qualitative and quantitative analysis, we showed that Libra effectively utilises temporal relationships between current and prior scans, addressing challenges such as hallucinations in referencing prior studies.
+This highlights Libra’s ability to generate clinically accurate and temporally consistent radiology reports, setting a new paradigm for multimodal medical AI research.
+Future work will focus on expanding Libra’s clinical applicability by incorporating diverse imaging modalities and enhancing temporal reasoning capabilities, and extending it in an agentic way.
+Limitations
+Despite Libra’s ability to model temporal paired images for radiology report generation (RRG), certain limitations remain.
+First, Libra relies on single prior images for temporal modelling, whereas clinical practice often involves multiple prior scans with varied intervals and angles.
+Extending the model to handle multiple temporally sequenced images remains an open challenge.
+Second, our study is based on a single-source dataset with inherent biases in patient demographics and imaging protocols, which may limit generalizability across broader clinical settings.
+Lastly, while Libra is designed for CXR-based RRG, its applicability to other imaging modalities (e.g., CT, MRI) and integration with structured medical knowledge remains unexplored.
+For a detailed discussion of these limitations and future directions, see Appx.
+J.
+Ethics Statement
+This work presents Libra, a model designed to enhance radiology report generation by integrating temporal and visual information.
+While Libra has the potential to improve clinical workflows, reduce radiologist workload, and enhance diagnostic consistency, its deployment must be approached with caution to ensure ethical and responsible use.
+Our research exclusively utilises the publicly available and “de-identified” MIMIC-CXR dataset (Johnson et al., 2019b), in accordance with its official documentation, ensuring adherence to ethical and privacy standards under CITI Data or Specimens Only Research certification.
+By relying solely on open datasets, we prioritise transparency and reproducibility, aligning with best practices in ethical AI research.
+This work is intended to support, not replace, medical professionals, ensuring it serves as a complementary tool within clinical practice.
+While the societal implications are largely positive, further validation across diverse patient populations and healthcare systems is necessary to address potential biases inherent in the dataset.
+Additionally, it is crucial to mitigate the risks of over-reliance on AI systems, which could inadvertently undermine human oversight or exacerbate healthcare disparities.
+Future efforts will aim to extend the model’s capabilities to encompass multiple imaging modalities and broader datasets, ensuring greater generalisability, fairness, and adaptability across diverse clinical settings.
+Appendix Contents
+[…]
+Appendix A Related Work
+A.1 Radiology Report Generation
+Radiology report generation (RRG) aims to address the long-tail distribution of observations in chest X-rays (CXRs) and produce fine-grained descriptions of clinical findings, making it a key objective in automated medical imaging analysis (Wang et al., 2018).
+Early RRG systems relied on recurrent neural networks (RNNs) (Liu et al., 2019), which have since been largely replaced by transformer-based architectures (Miura et al., 2021b; Chen et al., 2022), including large language models (LLMs) such as PaLM (Chowdhery et al., 2022) and Vicuna-7B (Chiang et al., 2023).
+These models excel at language generation, offering substantial improvements in fluency and factual accuracy.
+To further enhance clinical accuracy, some methods incorporate reinforcement learning (RL) to optimise for task-specific rewards, such as capturing “clinically relevant” features (Liu et al., 2019; Irvin et al., 2019) or maintaining logical consistency (Miura et al., 2021a; Delbrouck et al., 2022a).
+However, these approaches often rely on external tools like CheXbert (Smit et al., 2020a) or RadGraph (Jain et al., 2021), adding complexity to the optimisation process.
+Recent advancements in LLMs have shown that plain auto-regressive language modelling can achieve strong performance in RRG tasks.
+However, RL-based objectives and task-specific optimisations remain complementary, offering additional opportunities for improvement.
+Research on leveraging temporal information in RRG tasks can be broadly categorised into LLM-based and non-LLM-based methods, each presenting distinct advantages and challenges.
+A.2 LLM-based Model
+LLM-based models have achieved significant success in the RRG task, primarily due to advancements in visual instruction tuning (Liu et al., 2023).
+Structurally, these models (Li et al., 2023a; Chaves et al., 2024; Hyland et al., 2024; Zhou et al., 2024; Park et al., 2024) typically consist of an image encoder and an adapter that connects the encoder’s outputs to the LLM.
+Most existing adapters use single-layer hidden representations (e.g., the last or penultimate layer) from pre-trained image encoders, limiting their ability to integrate features from multiple images effectively.
+In end-to-end training, LLM-based models handle multiple image inputs by concatenating them with textual prompts, forming a composite input to the LLM.
+For instance, the input format is often structured as “<Current Image Placeholder> + <Prior Image Placeholder> + <Prompt>”.
+However, this approach provides limited guidance on the relationship between the images within the prompt.
+Ding et al.
+(2024) proposed the High-Resolution Instruction-Aware Adapter (HiA) to refine image-text representations, improving the model’s ability to follow textual prompts with multiple images.
+While this enhances instruction adherence, it does not explicitly model relationships between paired images.
+In contrast to this vanilla approach, Libra explicitly models temporal relationships in paired images through its Temporal Alignment Connector (TAC).
+Instead of simply concatenating images in the LLM’s latent space, TAC leverages all hidden-layer features from the image encoder to provide richer feature representations.
+By directly modelling temporal dynamics, Libra enables more precise and context-aware radiology report generation.
+A.3 non-LLM-based Model
+Non-LLM-based models typically employ transformer encoder-decoder architectures or their variants, which often require separate training for individual modules.
+These approaches handle “single-” and “double-” image inputs by symbolically differentiating tasks and employing distinct architectures tailored for each input type.
+Additionally, they frequently incorporate extra information such as prior reports, symptom labels, and knowledge graphs.
+For instance, Serra et al.
+(2023) uses symbolic alignment in its Longitudinal Projection Module along with a separately trained BERT-based (Devlin et al., 2019) text generator.
+RECAP (Hou et al., 2023a) implements a two-stage training process: classification tasks followed by report generation, leveraging a transformer encoder-decoder with symbolic task differentiation.
+TiBiX (Sanjeev et al., 2024) incorporates causal attention layers and learnable padding tokens to handle cases without prior images, while BioViL-T (Bannur et al., 2023) is a self-supervised vision-language training framework that features a CNN–Transformer hybrid multi-image encoder trained jointly with a BERT-based text model.
+On one hand, the difference in model parameter sizes, and on the other, as LLM-based models generally outperform other types of models (i.e.
+non-LLM-based) in the RRG task, papers on non-LLM-based models or those using small language models (SLMs) typically do not compare their methods with LLM-based approaches.
+Nonetheless, we conducted comparisons and discussions to reaffirm this observation, as detailed in Appx.
+G.2.
+A.4 Radiological Image Representation
+Radiology-specific pre-trained image encoder models are essential for RRG tasks due to the unique characteristics of radiological images, which fall outside the distribution of general-domain image models (Pérez-García et al., 2024).
+Several notable advancements have been made in this domain.
+Zhou et al.
+(2023) proposed Masked Record Modeling (MRM), a unified framework combining self-supervision with radiology report supervision to enhance radiograph representation learning.
+Similarly, BioViL-T (Bannur et al., 2023) employs a CNN-Transformer hybrid architecture to model multimodal relationships and leverage temporal structures for tasks such as disease progression classification and report generation.
+In addition, BiomedCLIP (Zhang et al., 2024c) is a multimodal biomedical foundational model pre-trained across diverse biomedical tasks.
+RAD-DINO (Pérez-García et al., 2024) is a medical image encoder that employs a pure image-based self-supervised learning approach from DINOv2 (Oquab et al., 2024) for continuous pretraining, focusing exclusively on image data to avoid the limitations of text supervision.
+Recent works have extensively applied RAD-DINO to RRG tasks, including MAIRA-2 (Bannur et al., 2024) and M4CXR (Park et al., 2024).
+Notably, Pérez-García et al.
+(2024) demonstrated that RAD-DINO outperforms other image encoders in RRG tasks.
+Building on this evidence, our model incorporates RAD-DINO as its image encoder to ensure high-quality radiological image representations, providing a robust foundation for downstream RRG tasks.
+Appendix B Research Objectives
+B.1 Temporal Information
+Temporal changes are critical for understanding disease progression.
+In radiology, paired images and their corresponding reports document subtle evolutions of symptoms over time.
+This temporal information is often captured by comparing current scans with prior ones to highlight symptom evolution or newly identified findings.
+The relative positioning of scans within the timeline determines the extent of temporal information.
+Therefore, the relative timing between scans is key: when the prior scan is recent, reported changes tend to be minimal; conversely, an older prior scan reveals more pronounced differences.
+Importantly, while temporal context enriches the diagnostic narrative, it does not alter the factual observations present in the current scan—it merely provides additional layers of interpretative insight.
+B.2 Research Aims
+This study aims to enhance radiology report generation (RRG) by effectively incorporating temporal information into the modelling process.
+In clinical practice, chest X-ray (CXR) analysis often depends on comparing the current scan with the prior image to capture disease progression and evolution.
+Our primary objective is to leverage these temporal cues to generate more accurate, context-aware radiological reports that faithfully reflect both stable conditions and clinically significant changes.
+Unlike previous LLM-based models (discussed in Appx.
+A.2), which depend on the LLM to infer temporal information solely from text, our approach explicitly models temporal relationships at the architectural level.
+Inspired by the principle “structure determines function” (Fischer, 1894; Pauling et al., 1951; Watson and Crick, 1953), we introduce the Temporal Alignment Connector (TAC), a dedicated module designed to capture temporal dynamics.
+Details are provided in Sec.
+2.2.
+B.3 Research Scope
+This study focuses on frontal chest X-rays, treating each examination per image while incorporating a single prior image as an auxiliary input when available.
+Rather than modelling patient-level longitudinal history, our goal is to generate a report for the current image while leveraging temporal information from one preceding scan.
+To ensure fairness in benchmarking, Libra was evaluated on single-image inputs without priors (see Table 1).
+Yet, temporal information remains implicitly present through several factors:
+Explicit temporal states (e.g., “stable” or “unstable”) are frequently described in radiology reports.
+Latent temporal progression exists in the dataset, as prior studies influence diagnostic phrasing.
+The absence of a prior image itself constitutes a temporal scenario, representing an extreme case where the patient’s condition is assumed stable due to a lack of comparative reference.
+Our model can effectively handle scenarios with limited temporal information in the RRG task.
+For instance, in a case where a patient has two scans taken just milliseconds apart, the current and prior images would be nearly identical, as no pathological changes would manifest within such a short interval.
+This extreme scenario demonstrates how the model handles clinical practice under limited temporal information.
+In such cases, the correct diagnosis for this minimal interval would be that the patient’s condition is “stable”; our model should then generate a report reflecting this stability.
+When no prior image is available, we employ a dummy prior image (a copy of the current image) to maintain input consistency and mitigate spurious references to nonexistent priors.
+However, in clinical practice, patients often undergo multiple prior scans, sometimes from different orientations, providing a more complex temporal context.
+This lies beyond the scope of our current study, and a detailed discussion of such scenarios is provided in Appx.
+J.
+Appendix C Prompt Example
+| Original Radiology Report |
+| EXAMINATION: Chest (Portable AP) |
+| INDICATION: Dyspnea and cough, right-sided back pain. |
+| HISTORY: Intubation with pulmonary edema. |
+| COMPARISON: Chest radiographs on ___ and CT chest without contrast on ___. |
+| TECHNIQUE: Portable upright chest radiograph. |
+| FINDINGS: In comparison with the prior study, there are diffuse bilateral pulmonary opacifications, more prominent on the right.
+These findings could indicate severe pulmonary edema, but superimposed pneumonia or developing ARDS cannot be excluded.
+Monitoring and support devices are appropriately positioned.
+|
+| Prompt Content |
+| [System prompt]: { |
+| The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes. |
+| Image Representation Placeholder |
+| [Clinical prompt]: { |
+| Provide a detailed description of the findings in the radiology image. Following clinical context: |
+| Indication: Dyspnea and cough, right-sided back pain. |
+| History: Intubation with pulmonary edema. |
+| Comparison: Chest radiographs on ___ and CT chest without contrast on ___. |
+| Technique: Portable upright chest radiograph. |
+We selected examples from the MIMIC-CXR (Johnson et al., 2019b) dataset and synthesised them using GPT-4 (OpenAI et al., 2024) to ensure ethical compliance, as illustrated in Table 4.
+Following the rule-based approach by Hyland et al.
+(2024), we extracted key sections from the report of the current image.
+Each example combines a fixed system prompt with a dynamic clinical prompt tailored to the current scan.
+We utilised four clinical instructions from the original report: {Indication}, {History}, {Comparison}, and {Technique}.
+In contrast, MAIRA-2 (Bannur et al., 2024), which incorporates prior image reports, our approach focuses exclusively on the current image’s context, maintaining a clear distinction from prior study information of the report.
+Appendix D Training Configuration
+Libra is trained using a standard auto-regressive language modelling loss (cross-entropy).
+For this study, we employ Meditron-7b (Chen et al., 2023c) as the LLM, with a total batch size of 16 throughout the training process.
+The training is conducted on a computational infrastructure equipped with A6000 GPU (48GB of memory) and using DeepSpeed optimization (Rajbhandari et al., 2020) with ZeRO-2 for stage 1 and ZeRO-3 for stage 2, and BF16 precision is enabled.
+A cosine learning rate scheduler is employed, starting with a warm-up phase of .
+In the first stage of training, we run for 1 epoch (385 hours) with a learning rate of .
+In the second stage, the model is trained for 3 epochs (213 hours) at the same learning rate.
+The LoRA (Hu et al., 2021) parameters are set to and .
+The final checkpoint for all runs is selected based on the observation of the minimum loss on the evaluation dataset throughout the training process.
+Note: Prior work, especially in the medical domain, typically employs full model fine-tuning for RRG tasks.
+However, due to hardware constraints, we can only adopt a lightweight training technique for parameter-efficient adaptation.
+As a result, our approach may underperform full model fine-tuning strategies in the second stage, despite maintaining computational efficiency.
+Appendix E Datasets Description
+| Dataset | Task Type | # Samples | % Has Prior | ||||
 | Train (%) | Valid (%) | Test (%) | Train | Valid | Test | ||
@@
 | Total | Multi-type | 1 213 097 (100%) | 146 594 (100%) | 88 669 (100%) | 64.73 | 49.09 | 83.67 |
-Inspired by the multi-stage training techniques commonly employed in recent MLLMs fine-tuning methods (Zhu et al., 2023), we adopt a two-stage training strategy to infuse multimodal capabilities into the model.
-In the first stage, we leverage a large-scale dataset to train the adapter, primarily focusing on learning the distribution of CXR token embeddings and the general relationships between CXR and text tokens across the MIMIC dataset.
-The second stage fine-tuning the model on downstream tasks, where the text decoder (i.e., the LLM) is refined to focus on aligning high-granularity information between CXRs and the Findings section of radiology reports.
-In addition to Findings section generation, the temporal feature alignment stage also uses the Impression section and VQA tasks.
-The Impression section summarises diagnoses and proposes further investigations (Babar et al., 2021), aiding the alignment between CXRs and their textual descriptions.
-We use the same system and clinical prompts as for Findings, replacing ‘Findings’ with ‘Impression’.
-For VQA, system prompts remain the same, while clinical prompts adapt to medical-specific questions, guiding the model to generate corresponding captions.
-These fine-grained VQA tasks expand the MLLM’s biomedical vocabulary and improve the alignment of image-text annotations.
-3.2 Evaluation Metrics
-We evaluate the generated reports using lexical and radiology-specific metrics, following established protocols..
-For lexical metrics, we use ROUGE-L (Lin, 2004), BLEU- (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005) and BERT (Devlin et al., 2019).
-For radiology-specific metrics, we use RadGraph-F1 (Jain et al., 2021), RG (Delbrouck et al., 2022a), F1-CheXpert (Irvin et al., 2019a), CheXbert vector similarity (Yu et al., 2022), as well as RadCliQ version Yu et al.
-(2022).
-Additionally, we introduce a novel temporal entity F1 score (called ) to assess the model’s ability to capture temporal information.
-Clinical metrics typically emphasise the accuracy of medical findings, prioritising the detection of clinically relevant entities.
-In contrast, the temporal entity score specifically measures the accuracy of entities related to progression over time described in the report.
-Detailed descriptions of all metrics and an illustrative analysis are provided in Appx.
-B.3.
-Temporal Entity F1 Following the work of Bannur et al.
-(2023), we set a reward list that includes common radiology-related keywords associated with temporal changes.
-Temporal entities are then extracted from both the ground truth and the generated reports .
-Importantly, no stemming or lemmatization is applied during token processing to preserve the precise description of temporal changes.
-After extraction, we compute precision and recall , which are subsequently used to calculate the temporal entity score, defined as the harmonic mean of precision and recall (Van Rijsbergen, 1974), also known as the F1 score.
-where is a small value, set to a default of , to prevent division by zero (it is also added to the numerator for special cases where no temporal entities are present in the ground truth).
-The coefficient controls the balance between precision and recall, with a default value of .
-3.3 Main Result
-Although MIMIC-CXR provides an “official” test split, strict comparisons with prior studies remain challenging due to differences in test set inclusion criteria and pre-processing steps.
-For instance, Yu et al.
-(2022) and Jeong et al.
-(2023) included only one image per study, resulting in a test set of 1,597 samples, while Tanida et al.
-(2023) followed the Chest ImaGenome split (Wu et al., 2021b).
-Such variations in test set distributions can significantly impact the reported results (Park et al., 2024).
-To ensure fairness, we use a widely adopted test set focused on frontal-view CXRs, aligned with previous studies such as MAIRA-1 (Hyland et al., 2024) and LLaVA-Rad (Chaves et al., 2024) 444The test set consists of 2,461 frontal-view image-report samples..
-Additionally, the latest concurrent work M4CXR (Park et al., 2024) employs multi-turn chain-of-thought (Wei et al., 2023) prompting to generate reports, which differs from our task setup.
-Furthermore, we do not compare with the recent work MAIRA-2, as it is a MLLM specifically designed for grounded radiology report generation task, which incorporates lateral views and prior study reports for each subject within the input prompt.
-Bannur et al.
-(2024) emphasises a positive transfer between this task and general report generation, which is also beyond the scope of our study 555For comparisons with the latest concurrent and non-LLM-based models can be found in Appx.
-D..
-Taking these considerations into account, we compared our model with state-of-the-art models, including LLaVA-Med (Li et al., 2023a), CheXagent (Chen et al., 2024), GPT-4V (OpenAI et al., 2024), Med-PaLM (Tu et al., 2023), LLaVA-Rad and MAIRA-1.
-The results are shown in Table 2.
-Since many of these models are not publicly available, we present their evaluation results as reported in the original sources.
-| Metric | LLaVA-Med | CheXagent | GPT-4V | Med-PaLM | LLaVA-Rad | MAIRA-1 | Libra(%) |
-| Lexical: | |||||||
-| ROUGE-L | 27.6 | 21.5 | 13.2 | 27.5 | 30.6 | 28.9 | 36.7 (19.9%) |
-| BLEU-1 | 35.4 | 16.9 | 16.4 | 32.3 | 38.1 | 39.2 | 51.3 (30.9%) |
-| BLEU-4 | 14.9 | 4.7 | 17.8 | 11.5 | 15.4 | 14.2 | 24.5 (37.6%) |
-| METEOR | 35.3 | – | – | – | – | 33.3 | 48.9 (38.5%) |
-| Clinical: | |||||||
-| RadGraph-F1 | 19.1 | – | – | 26.7 | – | 24.3 | 32.9 (23.2%) |
-| RG | 23.8 | 20.5 | 13.2 | – | 29.4 | 29.6 | 37.6 (27.0%) |
-| RadCliQ() | 3.3 | – | – | – | – | 3.1 | 2.7 (12.9%) |
-| CheXbert vector | 36.9 | – | – | – | – | 44.0 | 46.9 (6.59%) |
-| CheXpert-F1: | |||||||
-| Micro-F1-14 | 42.7 | 39.3 | 35.5 | 53.6 | 57.3 | 55.7 | 55.9 (-2.4%) |
-| Macro-F1-14 | 26.9 | 24.7 | 20.4 | 39.8 | 39.5 | 38.6 | 40.4 (1.5%) |
-| Micro-F1-5 | 43.9 | 41.2 | 25.8 | 57.9 | 57.4 | 56.0 | 60.1 (3.8%) |
-| Macro-F1-5 | 36.3 | 34.5 | 19.6 | 51.6 | 47.7 | 47.7 | 53.8 (4.3%) |
-From Table 2 we can see that, for all traditional lexical metrics, Libra substantially outperforms the baseline models, achieving scores of 36.7 for ROUGE-L, 51.3 and 24.5 for BLEU-1 and BLEU-4 respectively, and 48.9 for METEOR.
-In clinical metrics, Libra also showed outstanding performance, substantially improving in RadGraph-F1 and its variant RG, with scores of 32.9 and 37.6, respectively.
-Notably, Libra achieved the highest score (i.e.
-2.7) on the radiologist-aligned RadCliQ metric.
-For CheXbert-derived metrics, it also secured the top position with a vector embedding matrix score of 46.9.
-In the CheXpert 5-class subset, Libra excelled with Micro-F1 and Macro-F1 scores of 60.1 and 53.8, respectively.
-In the CheXpert 14-class subset, Libra achieved the highest Macro-F1 score of 40.4, while its Micro-F1 score of 55.9 was only slightly below the leading model.
-In summary, Libra demonstrated outstanding performance across both sets of metrics, with only minor gaps in one specific measure.
-The remarkable capabilities of Libra can be attributed to its innovative Temporal Alignment Connector which effectively captures temporal changes in the dataset, enabling the generation of highly accurate and clinically relevant reports.
-4 Ablation Study
-To thoroughly assess the impact of various optimisation components, we conducted a series of ablation experiments, evaluating different modules and dataset expansions within Libra.
-All ablation models were tested on the MIMIC-CXR official test split for the Findings section generation task, using identical hyperparameter settings during training and inference.
-Q1: Does incorporating temporal information bring positive effects to Libra in RRG tasks?
-| Metric | Libra | |
-| w/ true | w/ dummy(%) | |
-| Lexical: | ||
-| ROUGE-L | 36.66 | 36.17 (-1.34%) |
-| BLEU-1 | 51.25 | 51.20 (-0.10%) |
-| BLEU-4 | 24.54 | 24.33 (-0.86%) |
-| METEOR | 48.90 | 48.69 (-0.43%) |
-| BERTScore | 62.50 | 61.94 (-0.90%) |
-| F1 | 35.34 | 32.72 (-7.41%) |
-| Clinical: | ||
-| RadGraph-F1 | 32.87 | 32.42 (-1.37%) |
-| RG | 37.57 | 36.92 (-1.73%) |
-| RadCliQ() | 2.72 | 2.76 (-1.47%) |
-| CheXbert vector | 46.85 | 46.31 (-1.15%) |
-| CheXpert-F1: | ||
-| Micro-F1-14 | 55.87 | 55.25 (-1.11%) |
-| Macro-F1-14 | 40.38 | 40.15 (-0.57%) |
-| Micro-F1-5 | 60.07 | 58.93 (-1.90%) |
-| Macro-F1-5 | 53.75 | 52.61 (-2.12%) |
-Temporal information is embedded in paired images and referenced in the corresponding radiology reports, reflecting changes over time through references to previous symptoms and descriptions of their progression.
-As indicated in Table 1, 86% of the test data includes true prior images.
-To assess whether Libra can effectively perceive and utilise temporal information during inference, we conducted separate evaluations using either true prior images or only dummy prior images.
-When true prior images were unavailable, the model treated the current image as a dummy prior image.
-As shown in Table 3, Libra demonstrated substantial enhancements across all metrics when true prior images were used as references.
-Specifically, the improvement in clinical scores exceeds that of lexical scores, indicating that temporal information is crucial for generating high-quality medical reports, beyond merely improving linguistic fluency.
-Notably, the F1 score shows the most significant impact, with a difference of 7.41%.
-This highlights Libra’s capability to effectively leverage temporal changes provided by true prior images, thereby enhancing the quality of the generated Findings section and overall performance in the RRG task.
-Q2: Does the Temporal Alignment Connector effectively improve model performance?
-| Metric | Libra-1 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
-| Lexical: | |||||
-| ROUGE-L | 27.56 | 27.33 (-0.85%) | 27.21 (-1.27%) | 27.43 (-0.48%) | 26.17 (-5.04%) |
-| BLEU-1 | 34.84 | 34.17 (-1.92%) | 34.21 (-1.82%) | 34.60 (-0.67%) | 33.03 (-5.20%) |
-| BLEU-4 | 11.51 | 11.13 (-3.33%) | 11.11 (-3.47%) | 11.43 (-0.73%) | 10.02 (-12.98%) |
-| METEOR | 35.50 | 35.06 (-1.24%) | 34.96 (-1.52%) | 35.28 (-0.62%) | 33.98 (-4.28%) |
-| BERTScore | 55.87 | 55.60 (-0.49%) | 55.49 (-0.69%) | 55.74 (-0.23%) | 54.63 (-2.22%) |
-| F1 | 26.63 | 25.96 (-2.51%) | 26.21 (-1.57%) | 26.58 (-0.18%) | 25.39 (-4.65%) |
-| Clinical: | |||||
-| RadGraph-F1 | 22.52 | 22.20 (-1.42%) | 22.03 (-2.19%) | 22.35 (-0.74%) | 21.51 (-4.48%) |
-| RG | 27.32 | 26.89 (-1.59%) | 26.72 (-2.19%) | 27.09 (-0.84%) | 25.97 (-4.96%) |
-| RadCliQ () | 3.10 | 3.12 (-0.65%) | 3.12 (-0.65%) | 3.11 (-0.32%) | 3.15 (-1.61%) |
-| CheXbert vector | 42.02 | 41.57 (-1.07%) | 41.37 (-1.54%) | 41.92 (-0.24%) | 40.93 (-2.59%) |
-| CheXpert-F1: | |||||
-| Micro-F1-14 | 52.48 | 51.74 (-1.42%) | 51.68 (-1.53%) | 52.13 (-0.67%) | 51.13 (-2.57%) |
-| Macro-F1-14 | 36.87 | 36.04 (-2.25%) | 36.12 (-2.03%) | 36.14 (-1.97%) | 35.85 (-2.76%) |
-| Micro-F1-5 | 56.63 | 55.37 (-2.23%) | 55.79 (-1.49%) | 55.87 (-1.34%) | 54.51 (-3.74%) |
-| Macro-F1-5 | 49.33 | 47.76 (-3.18%) | 47.82 (-3.06%) | 47.98 (-2.75%) | 47.22 (-4.28%) |
-To evaluate the impact of TAC on Libra’s performance in the RRG task, we initialised a model with the RAD-DINO (Pérez-García et al., 2024) image encoder, TAC, and Meditron (Chen et al., 2023b) as the LLM.
-A baseline experiment (Libra-1) was conducted by fine-tuning only the TAC for the Findings generation task.
-As shown in Table 4, we performed ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC.
-Removing TFM restricted the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023) but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions.
-Without LFE, the model follows the LLaVA setup, utilising the penultimate layer of the image encoder.
-The results indicate that removing any TAC submodule leads to a decline in all metrics compared to Libra-1.
-Removing TFM caused a notable drop in the F1 score (2%), emphasising its role in capturing temporal information.
-Removing LFE especially decreases RadGraph-related scores, highlighting its importance in extracting detailed image features.
-The removal of PIPB impacted clinical scores more than lexical scores.
-Finally, removing the entire TAC results in substantial declines across all metrics, underscoring its critical role in integrating image details and temporal information666For more ablation studies, including the impact of radiology-specific pre-trained models, see Appx.
-C..
-Q3: Are additional Impression and VQA datasets needed during the feature alignment?
-| Metric | Stage: 1 | Stage: 2 | ||
-| Libra | Libra-f | Libra | Libra-f | |
-| Lexical: | ||||
-| ROUGE-L | 27.27 | 27.56△ | 36.66 | 35.31▼ |
-| BLEU-1 | 41.24 | 34.84▼ | 51.25 | 49.92▼ |
-| BLEU-4 | 13.59 | 11.51▼ | 24.54 | 23.05▼ |
-| METEOR | 39.44 | 35.50▼ | 48.90 | 47.99▼ |
-| BERTScore | 56.00 | 55.87▼ | 62.50 | 61.28▼ |
-| F1 | 24.80 | 26.63△ | 35.34 | 33.52▼ |
-| Clinical: | ||||
-| RadGraph-F1 | 20.45 | 22.52△ | 32.87 | 30.77▼ |
-| RG | 25.19 | 27.32△ | 37.27 | 35.44▼ |
-| RadCliQ () | 3.31 | 3.10△ | 2.72 | 2.83▼ |
-| CheXbert vector | 35.33 | 42.02△ | 46.85 | 45.32▼ |
-| CheXpert-F1: | ||||
-| Micro-F1-14 | 43.63 | 52.48△ | 55.87 | 54.11▼ |
-| Macro-F1-14 | 25.68 | 36.87△ | 40.38 | 37.16▼ |
-| Micro-F1-5 | 49.75 | 56.63△ | 60.07 | 58.76▼ |
-| Macro-F1-5 | 40.40 | 49.33△ | 53.75 | 51.99▼ |
-To evaluate the impact of dataset expansion during the first stage on Libra’s performance, we trained a model (Libra-f) using only the Findings data in the first stage, while Libra incorporated additional Impression and VQA data for alignment, in Table 5.
-After the first stage, Libra showed superior performance in lexical metrics compared to Libra-f, but a slight decline in clinical metrics.
-This can be attributed to the inclusion of VQA tasks, which encouraged the model to focus on more fine-grained and grounded information.
-While VQA emphasises the detailed description of the single symptom, the Findings section generation task requires a comprehensive overview of the CXR, encompassing multiple normal and abnormal findings.
-This shift in focus also impacted the F1 score, as each finding typically involves its own temporal changes.
-Additionally, the reduction in the number of generated findings entities led to a decrease in the number of identified temporal entities.
-In the second stage, fine-tuned on the Findings dataset, Libra achieved the best results across all metrics, indicating that incorporating additional datasets in the first stage enhances Libra’s understanding and reasoning of CXRs.
-5 Performance Analysis
-We used CXRs from the official test split to generate Findings sections and perform a qualitative analysis of Libra.
-Figure 2 illustrates a case where no true prior image was available for reference.
-For example, while the ground truth report only mentioned the presence of “sternal wires,” Libra not only identified their presence but also provided details on the specific type.
-This demonstrates Libra can deliver more precise clinical information by not only recognizing conditions but also distinguishing the specific types of surgical interventions.
-Original
-report
-The lungs are well expanded and clear. The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal. No pleural effusion or pneumothorax is present. Sternal wires are intact.
-Libra
-The lungs are well expanded and clear.
-The cardiomediastinal silhouette, hilar contours, and pleural surfaces are normal.
-No pleural effusion or pneumothorax is present.
-Median sternotomy wires are intact.
-For the samples with true prior images, Libra accurately describes the presence and changes in radiological conditions.
-As shown in the Figure 3, the current image revealed additional findings of pleural effusion and pneumonia compared to the prior image.
-The model effectively detected these progressive changes, provided detailed descriptions, and suggested potential causes for further investigation.
-Notably, it explicitly mentioned the comparative information relative to the prior image, offering a clear understanding of the temporal changes.
-Current image
-Prior image
-Original report (current image)
-In comparison
-with the study of , there is again evidence of a superior mediastinal mass consistent with the known thyroid enlargement that is better seen on the recent CT scan.
-Continued bibasilar opacification most likely represents atelectasis, though in the appropriate clinical setting, superimposed pneumonia would have to be considered.
-The cardiac silhouette is within normal limits, though there may be some mild elevation of pulmonary venous pressure.
-Small pleural effusions are probably present bilaterally.
-Libra
-As compared to
-radiograph, cardiomediastinal contours are stable.
-New patchy and linear opacities have developed at the lung bases, right greater than left, and may reflect patchy atelectasis, aspiration, or developing infectious pneumonia.
-Small pleural effusions are present bilaterally.
-w/ dummy prior image
-Cardiac silhouette is upper limits of normal in size and accompanied by pulmonary vascular congestion and minimal interstitial edema.
-Patchy bibasilar opacities are present, and could reflect patchy atelectasis, aspiration, or developing infectious pneumonia.
-Followup radiographs may be helpful in this regard.
-Small bilateral pleural effusions are also demonstrated.
-Swap order
-Cardiomediastinal contours are normal.
-Previously present bibasilar opacities have nearly completely resolved with minimal residual linear opacification remaining at the right lung base.
-No new areas of consolidation are identified, and there are no pleural effusions.
-Original report (prior image)
-The heart is normal in size.
-The mediastinal and hilar contours appear stable.
-The aortic arch is calcified.
-There is no definite pleural effusion or pneumothorax.
-Streaky opacity in the right costophrenic angle suggests minor atelectasis, which has decreased.
-There is similar mild spinal curvature.
-To further assess the model’s robustness, we conducted a more detailed analysis of this case.
-First, we generated a report using only the dummy prior image as a reference.
-While Libra successfully detected abnormal findings, it failed to specify changes relative to the previous image, describing only the current state without indicating disease progression.
-Moreover, it provided less clarity in distinguishing the distribution and severity of abnormalities.
-This suggests that, without a true prior image for comparison, the model struggles to accurately describe the progression of findings.
-Evaluating Temporal Consistency in Libra’s Report Generation To investigate the model’s ability to capture temporal changes, we swapped the order of the images, using the true prior image as the current image and vice versa.
-Interestingly, the generated report indicated an improvement in the patient’s condition, with most abnormalities resolved and no new findings detected, which is contrary to the ground truth of the original current image report.
-Moreover, the report generated from the swapped images closely resembled the original report of the true prior image, as shown at the bottom of Figure 3, which described a stable or improving condition for the patient.
-This striking observation demonstrates that Libra can effectively utilise both current and prior studies to generate accurate reports for assessing changes, simulating the conditions of standard clinical practice.
-6 Conclusion
-In this study, we presented Libra, a temporal-aware multimodal large language model specifically designed for chest X-ray report generation tasks.
-Trained exclusively on the open-access MIMIC-CXR dataset (Johnson et al., 2019b), Libra utilises a two-stage training framework that leverages a radiology-specific pre-trained image encoder and language model, connected through a specially designed Temporal Alignment Connector to bridge visual and textual modalities.
-By identifying observations from current CXR and referencing prior scan, Libra generates radiology reports with substantial improvements across all metrics compared to existing models of the same parameter scale.
-Qualitative analysis in the case study highlights Libra’s ability to effectively utilise the temporal relationships between current and prior images, addressing the challenge of hallucinations related to prior study references in RRG tasks.
-Libra sets a new paradigm for MLLMs in multimodal medical AI research.
-In the future, we aim to further enhance Libra’s clinical applicability and accuracy, paving the way for the development of general-purpose AI models in the medical domain.
-For a detailed discussion of Libra’s limitations and future work, please refer to Appx.
-A.3.
-Appendix A Appendix A
-A.1 Related Work
-Radiology Report Generation Several efforts have been made to develop models tailored for radiology report generation (RRG), driven by the need to address the long-tail distribution of observations in chest X-rays and provide fine-grained descriptions of findings.
-This task has gained prominence as an important objective for the designed system (Wang et al., 2018b).
-Such systems inherently require a language generation component.
-Initially, recurrent neural networks were used (Liu et al., 2019), but these have recently been replaced by Transformer architectures (Miura et al., 2021b; Chen et al., 2022), including large language models (LLMs) such as PaLM (Chowdhery et al., 2022) and Vicuna-7B (Chiang et al., 2023).
-To ensure clinical accuracy in generated text, some studies have moved beyond standard language modelling losses, instead employing reinforcement learning (RL) to optimise for rewards that prioritise “clinically relevant” features, such as the inclusion of specific findings (Liu et al., 2019; Irvin et al., 2019b) or maintaining logical consistency (Miura et al., 2021a; Delbrouck et al., 2022a).
-However, these methods often depend on external models like CheXbert (Smit et al., 2020) or RadGraph (Jain et al., 2021) to extract clinical entities, adding complexity to the optimisation process.
-With the advancement of LLMs, an increasing number of studies have demonstrated that plain auto-regressive language modelling can achieve substantial results.
-However, we acknowledge that the benefits of more complex training objectives or RL-based approaches may be complementary.
-Consequently, research on leveraging temporal information in RRG tasks can be broadly categorised into two types: LLM-based methods and non-LLM-based methods.
-non-LLM-based Model These architectures are typically transformer encoder-decoder models and their variants.
-They are often composed of multiple modules that require separate training.
-Meanwhile, they handle single and double image inputs by symbolically differentiating tasks and employing two distinct neural network architectures: one designed for single-image inputs and another for double-image inputs.
-Serra et al.
-(2023) utilises symbolic alignment in its Longitudinal Projection Module and employs a separately trained BERT-based (Devlin et al., 2019) text generator.
-RECAP (Hou et al., 2023a) implements a transformer encoder-decoder with symbolic task differentiation and two-stage training, beginning with classification tasks followed by report generation.
-TiBiX (Sanjeev et al., 2024) adopts a transformer model with causal attention layers and incorporates learnable padding tokens to handle cases without prior images.
-BioViL-T (Bannur et al., 2023) is a self-supervised vision-language training framework that features a CNN–Transformer hybrid multi-image encoder trained jointly with a BERT-based text model.
-On one hand, the difference in model parameter sizes, and on the other, as LLM-based models generally outperform other types of models in the RRG task, papers on non-LLM-based models or those using small language models typically do not compare their methods with LLM-based approaches.
-At the same time, we conducted comparisons and discussions to reaffirm this point, as detailed in Appx.
-D.2.
-LLM-based Model Current LLM-based models have achieved notable success in the RRG task, largely benefiting from visual instruction tuning (Liu et al., 2023).
-Structurally, these models (Li et al., 2023a; Chaves et al., 2024; Hyland et al., 2024; Zhou et al., 2024; Park et al., 2024) typically consist of an image encoder and an adapter that bridges the connection to the LLM.
-These adapters often rely on single-layer hidden representations (e.g., the last or penultimate layer) from pre-trained image encoders, limiting their ability to simultaneously integrate information from multiple images.
-In end-to-end training, LLM-based models handle multiple image inputs by concatenating them with textual prompts to form a composite input for the LLM.
-For example, the input format is “<Current Image Placeholder> + <Prior Imgae Placeholder> + <Prompt>”.
-Although our model adopts an end-to-end training paradigm, it does not rely on the vanilla approach to handle multiple images.
-Instead, we designed a novel adapter called the Temporal Alignment Connector (TAC).
-On one hand, TAC leverages all hidden layer features from the image encoder, setting it apart from the default configurations used in previous works.
-On the other hand, TAC models temporal information directly rather than simply concatenating two images in the latent space of the LLM.
-Radiological Image Representation As radiology lies outside the distribution of most general-domain image models, a radiology-specific pre-trained model is crucial for the RRG task (Pérez-García et al., 2024).
-Zhou et al.
-(2023) introduces a unified framework, Masked Record Modeling (MRM), which combines self-supervision and radiology report supervision to enhance radiograph representation learning.
-BioViL-T (Bannur et al., 2023) integrates a CNN-Transformer hybrid architecture for multimodal modelling, leveraging temporal structures in medical imaging and reports to perform tasks such as disease progression classification and report generation.
-BiomedCLIP (Zhang et al., 2024b) is a multimodal biomedical foundational model pre-trained on various biomedical tasks.
-RAD-DINO (Pérez-García et al., 2024) is a medical image encoder that employs a pure image-based self-supervised learning approach from DINOv2 (Oquab et al., 2024) for continuous pretraining, focusing exclusively on image data to avoid the limitations of text supervision.
-In recent parallel works, RAD-DINO has been widely applied to RRG tasks, including MAIRA-2 (Bannur et al., 2024), M4CXR (Park et al., 2024).
-Furthermore, Pérez-García et al.
-(2024) highlighted that RAD-DINO outperforms other types of image encoders in RRG tasks.
-Therefore, our model also adopts RAD-DINO as the image encoder.
-A.2 Research Objectives
-Temporal Information Temporal change is a crucial factor in understanding disease progression, especially in radiology, where paired images and corresponding radiology reports capture changes over time.
-This temporal context is often reflected in comparisons between current and prior scans, documenting symptom evolution or newly identified findings.
-The relativity of position within the timeline is a key determinant of temporal information.
-This relative positioning does not alter the established description of well-established facts about symptoms already present in the current image.
-For instance, if the prior image is recent, descriptions of changes will be minimal, whereas an older prior image may highlight more pronounced changes.
-Research Object Our study focuses on the chest X-ray (frontal view), rather than the individual patient.
-Specifically, we aim to generate a report for the current image while referencing the temporal information provided by a single prior image as auxiliary information.
-Our model is primarily designed to handle scenarios with limited temporal information in the RRG task.
-For instance, in a case where a patient has two scans taken just milliseconds apart, the current and prior images would be nearly identical, as no pathological changes would manifest within such a short interval.
-This extreme scenario demonstrates how the model handles clinical practice under limited temporal information.
-In such cases, the correct diagnosis for this minimal interval would be that the patient’s condition is “stable.” In our design, the dummy prior image addresses scenarios where a true prior image is absent.
-However, patients often undergo multiple prior scans, which may include temporal information from different orientations.
-This more complex temporal information lies beyond the scope of our current study.
-We will provide a detailed discussion of this scenario in Appx.
-A.3.
-A.3 Limitations and Future Work
-While our work represents a step forward in leveraging temporal information for radiology report generation, it also has several limitations that warrant further exploration.
-Handling Multiple Prior Scans Our current model is designed to process a single prior scan alongside the current scan.
-While this approach aligns with standard clinical workflows, which typically prioritise the most recent prior study for comparisons, it overlooks scenarios where multiple prior scans could offer a richer temporal perspective.
-For instance, analysing a sequence of images spanning an extended period could provide deeper insights into gradual disease progression.
-Future efforts should focus on extending our framework to incorporate multiple prior scans efficiently, enabling a more nuanced understanding of temporal patterns in clinical data.
-Temporal Information Beyond Image Comparisons At present, our model captures temporal information through paired image comparisons and their corresponding textual reports.
-However, clinical assessments often draw upon a broader context, including historical notes, laboratory results, and other longitudinal patient data.
-Expanding our approach to integrate these diverse temporal data sources could facilitate a more holistic understanding of disease trajectories and patient history, significantly enhancing clinical applicability.
-Challenges with Sparse Temporal Data In cases where prior scans are unavailable or only minimally informative (e.g., scans taken within a very short time interval), our use of a “dummy prior image” addresses the lack of temporal data.
-However, the model’s ability to interpret and generate meaningful outputs under these constraints may still be limited.
-Future work could explore methods for imputing or synthesising temporal context to address such cases more effectively.
-Computational Complexity Our use of temporal alignment mechanisms and multi-layer feature integration introduces increased computational demands.
-This complexity poses challenges for real-time deployment, particularly in resource-constrained clinical environments.
-Generalisability Across Modalities and Datasets This study focuses exclusively on frontal-view chest X-rays and utilises the MIMIC-CXR dataset (Johnson et al., 2019b).
-However, our approach has not been evaluated on other imaging modalities (e.g., CT or MRI) or datasets such as CheXpert (Irvin et al., 2019b) or PadChest (Bustos et al., 2020).
-Based on the identified limitations, we propose several directions for future research and development:
-Develop frameworks for integrating multiple prior scans with dynamic temporal reasoning.
-Extend the model to handle multi-modal imaging and textual data for comprehensive diagnostics.
-Investigating the integration of temporal data from diverse clinical sources, including electronic health records (EHRs)
-Exploring lightweight model architectures for faster inference while maintaining high performance.
-These enhancements will not only address the current limitations but also extend the potential of temporal-aware multimodal models in radiology and beyond.
-Appendix B Appendix B
-B.1 Training Configuration
-We train the model using a standard auto-regressive language modelling loss (cross-entropy).
-For this study, we utilise Meditron-7b (Chen et al., 2023b) as the LLM, with a total batch size of 16 throughout the training process.
-The training is conducted on a computational infrastructure equipped with A6000 GPU (48GB of memory) and configured with DeepSpeed ZeRO-3 optimization (Rajbhandari et al., 2020) and BF16 precision enabled.
-A cosine learning rate scheduler is employed, starting with a warm-up phase of .
-In the first stage of training, we run for 1 epoch with a learning rate of .
-In the second stage, the model is trained for 3 epochs at the same learning rate.
-The LoRA (Hu et al., 2021) parameters are set to and .
-The final checkpoint for all runs is selected based on the observation of the minimum loss on the evaluation dataset throughout the training process.
-B.2 Detailed Description of Datasets
-MIMIC-CXR.
+MIMIC-CXR
 (Johnson et al., 2019b) This is a large, publicly accessible dataset comprising 377,110 DICOM images across 227,835 studies, each accompanied by a radiology report (Johnson et al., 2019b).
@@
 For each report, we extract the Findings, Impression, Indication, History, Comparison, and Technique sections using rule-based heuristics supported by the official MIMIC code repository (Johnson et al., 2018).
-For the Findings section generation task, we discard studies where the Findings could not be extracted, while allowing for missing content in other clinical sections.
+For the Findings section generation task, studies without extractable Findings are discarded, while other missing sections are permitted.
 The same approach is applied to the Impression section generation task.
 In all our experiments, we adhere to the official MIMIC-CXR dataset split.
-Additionally, we retrieve prior images by following the chronological order of studies as indicated by the official labels, selecting the closest prior study as the reference image.
+Meanwhile, we retrieve prior images by following the chronological order of studies as indicated by the official labels, selecting the closest prior study as the reference image.
 It is important to note that, to prevent data leakage between the train, validation, and test sets, prior images are retrieved only from within the same split.
-Medical-Diff-VQA.
+Medical-Diff-VQA
 (Hu et al., 2023) This dataset is a derivative of the MIMIC-CXR dataset, focused on identifying differences between pairs of main and reference images.
@@
 In total, this dataset comprises 700,703 question-answer pairs derived from 164,324 main-reference image pairs.
-The questions are divided into seven categories: abnormality, location, type, view, presence, and difference.
+As shown in Table 5, the questions are divided into seven categories: abnormality, location, type, view, presence, and difference.
 Each pair consists of a main (current) image and a reference (prior) image, both taken from different studies of the same patient.
@@
 Of the seven question types, the first six types focus on the main image, while the “difference” questions involve both images.
-MIMIC-Ext-MIMIC-CXR-VQA.
-(Bae et al., 2023) This dataset is an extension of the MIMIC-CXR dataset, specifically designed for VQA tasks within the medical domain, specifically focused on CXRs.
-It includes questions generated from 48 unique templates covering seven content types: presence, anatomy, attribute, abnormality, size, plane, and gender.
+MIMIC-Ext-MIMIC-CXR-VQA
+(Bae et al., 2023) This dataset extends MIMIC-CXR for VQA tasks tailored to CXRs.
+It includes questions generated from 48 unique templates covering seven content types: presence, anatomy, attribute, abnormality, size, plane, and gender, as shown in Table 5.
 Each template was developed with the guidance of board-certified medical experts to ensure clinical relevance, addressing both standard medical VQA content and more complex logical scenarios.
@@
 Since annotations are based on single images, the current image serves as a dummy prior image for all entries in our experiment.
-For this study, we carefully selected datasets that provide complete reports and temporal information (i.e., prior images) to ensure alignment with our research objectives for the RRG task.
+For this study, we carefully selected datasets that provide complete reports and temporal information (i.e., prior images) to ensure alignment with our research objectives (see Appx.
+B) for the RRG task.
 After thoroughly evaluating other datasets, we found them unsuitable for the following reasons:
-CheXpert.
-(Irvin et al., 2019b) This dataset includes annotated scans with label-specific annotations rather than full medical reports.
+CheXpert
+(Irvin et al., 2019) This dataset includes annotated scans with label-specific annotations rather than full medical reports.
 While useful for training image encoders or annotation models, it is not appropriate for the RRG task, which requires complete diagnostic reports.
-PadChest. (Bustos et al., 2020) Although it includes reports and corresponding prior images, its reports are in Spanish, placing cross-language training beyond the scope of our model.
-IU-Xray. (Demner-Fushman et al., 2016) This dataset lacks patient-level metadata and prior study information, which is critical for our focus on temporal information in chest X-rays.
-Chest ImaGenome Dataset. (Wu et al., 2021a)(a derivative of MIMIC-CXR) This dataset does not follow the official MIMIC-CXR split, risking data leakage across train, validation, and test sets.
-B.3 Evaluation Metrics Details
-Lexical Metrics We employed common natural language generation metrics to quantify the overlap between generated and reference reports.
+PadChest
+(Bustos et al., 2020) Although it includes reports and corresponding prior images, its reports are in Spanish, placing cross-language training beyond the scope of our model.
+IU-Xray
+(Demner-Fushman et al., 2016) This dataset lacks patient-level metadata and prior study information, which is critical for our focus on temporal information in chest X-rays.
+Chest ImaGenome Dataset
+(Wu et al., 2021) Although derived from MIMIC-CXR (Johnson et al., 2019b), it does not follow the official split, raising concerns about potential data leakage between training, validation, and test sets.
+Meanwhile, the following two datasets were processed using GPT-4 (OpenAI et al., 2024) to eliminate hallucinated references to prior exams.
+While this prevents erroneous comparisons, it also removes essential temporal information originally present in the reports, potentially affecting tasks that rely on temporal reasoning.
+LLaVA-Rad MIMIC-CXR Dataset
+(Chaves et al., 2024) This dataset was refined using GPT-4 (OpenAI et al., 2024) through a structured text-cleaning pipeline.
+The process involved: (1) correcting typographical errors and split words, (2) removing redundant or repeated phrases to improve clarity, (3) eliminating explicit temporal references (e.g., “Compared to the prior study, no significant interval change was noted”) to ensure the report focuses exclusively on the current image, and (4) restructuring content into standardised sections, including Indication, Findings, and Impression.
+ReXPref-Prior Dataset
+(Banerjee et al., 2024) A modified version of MIMIC-CXR (Johnson et al., 2019b) in which GPT-4 (OpenAI et al., 2024) systematically removes all references to prior exams from both the Findings and Impression sections.
+While this adjustment prevents spurious prior-study references, it also eliminates crucial temporal context, limiting its suitability for applications requiring longitudinal assessment of disease progression.
+Appendix F Evaluation Metrics
+F.1 Lexical Metrics
+We employed standard natural language generation metrics to quantify the overlap between generated and reference reports.
 Specifically, ROUGE-L (Lin, 2004) measures the length of the longest common subsequence between the generated and reference reports.
 BLEU- (Papineni et al., 2002) calculates n-gram precision and applies a brevity penalty to discourage overly short predictions.
-METEOR (Banerjee & Lavie, 2005), computes the weighted harmonic mean of unigram precision and recall, with an additional penalty for fragmenting consecutive word sequences.
+METEOR (Banerjee and Lavie, 2005), computes the weighted harmonic mean of unigram precision and recall, with an additional penalty for fragmenting consecutive word sequences.
 Finally, we report BERTScore (Zhang et al., 2020a), which leverages pre-trained contextual embeddings from BERT (Devlin et al., 2019) to match words in candidate and reference sentences based on cosine similarity.
 We used default parameters for all of these evaluation metrics.
+F.2 Clinical Metrics
 For radiology-specific metrics, we used as many of the same evaluation scores as possible from previous studies (Tu et al., 2023; Hyland et al., 2024; Bannur et al., 2024; Chaves et al., 2024), including the following:
-RadGraph-based metrics The RadGraph model (Jain et al., 2021) is designed to parse radiology reports into structured graphs.
+RadGraph-based metrics
+RadGraph model (Jain et al., 2021) is designed to parse radiology reports into structured graphs.
 These graphs consist of clinical entities, which include references to anatomy and observations, as well as the relationships between these entities.
@@
 And a variant of it, RG (Delbrouck et al., 2022b), which matches entities based on their text, type, and whether they have at least one relation777RG is implemented as F1RadGraph with reward=partial by the radgraph package..
-CheXpert F1 This set of metrics utilizes the CheXbert automatic labeler (Smit et al., 2020) to extract “present”, “absent”, or “uncertain” labels for each of the 14 CheXpert pathologies (Irvin et al., 2019a) from the generated reports and their corresponding references.
-In line with prior work, we report CheXpert-F1 for all 14 classes, as well as for the 5 most common findings in real-world CXR reports, referring to these as ‘[Macro/Micro]-F1-[5/14]’.
-CheXbert vector similarity We also employ CheXbert vector similarity (Yu et al., 2022), which calculates the cosine similarity between the embeddings of the generated and reference reports after processing them through the CheXbert model (Smit et al., 2020).
-RadCliQ In addition, we utilise RadCliQ (Radiology Report Clinical Quality), a composite metric that combines RadGraph-F1 and BLEU scores in a linear regression model to estimate the number of errors that radiologists are likely to detect in a report.
+CheXpert F1
+This set of metrics utilizes the CheXbert automatic labeler (Smit et al., 2020a) to extract “present”, “absent”, or “uncertain” labels for each of the 14 CheXpert pathologies (Irvin et al., 2019) from the generated reports and their corresponding references.
+In line with prior work, we report CheXpert-F1 for all 14 classes, as well as for the 5 most common findings in real-world CXR reports, referring to these as “[Macro/Micro]-F1-[5/14]”.
+CheXbert vector similarity
+RadCliQ
+In addition, we utilise RadCliQ (Radiology Report Clinical Quality) (Yu et al., 2022), a composite metric that combines RadGraph-F1 and BLEU scores in a linear regression model to estimate the number of errors that radiologists are likely to detect in a report.
 To maintain consistency with previous research, we use version of it.
 Both the CheXbert vector similarity, RadCliQ, and RadGraph-F1 metrics are calculated using the code released by Yu et al. (2022).
-Temporal Entity F1 The , which is specifically designed to detect temporal entities.
-This score is neither a traditional lexical metric nor a adiology-specific metrics.
-Here is an example to help you understand better, as shown in Table 6.
+F.3 Temporal Entity F1
+We introduced , a metric specifically designed to detect temporal entities reflecting changes over time.
+Unlike traditional lexical or radiology-specific metrics, evaluates the quality of temporal information in radiology reports.
 | Ground Truth | Candidate | ROUGE-L | RadGraph-F1 | |
@@
 | The pleural effusion is noted again on the current scan. | 0.22 | 0.80 | 0.0 |
-From these results, it is clear that the differences in lexical (ROUGE-L) and clinical (RadGraph-F1) metrics between the two candidates are relatively smaller compared to the score.
-This highlights that Temporal Entity F1 effectively captures and evaluates the quality of temporal information in radiology reports, distinguishing it more accurately than other standard metrics in the context of temporal information descriptions.
-B.4 Prompt Example
-Here, we select examples from the MIMIC-CXR (Johnson et al., 2019b) dataset as instances and further synthesise them using GPT-4 (OpenAI et al., 2024) to avoid any ethical concerns, as shown in Table 7.
-Following Hyland et al.
-(2024)’s work, we adopt a rule-based method to extract relevant sections from the report of the current image, utilising a fixed system prompt and a dynamic clinical prompt for each current scan.
-We utilised four clinical instructions from the original report: {Indication}, {History}, {Comparison}, and {Technique}.
-Notably, unlike MAIRA-2 (Bannur et al., 2024), we do not utilise the report of the prior image.
-| Original Radiology Report |
-| EXAMINATION: Chest (Portable AP) |
-| INDICATION: Dyspnea and cough, right-sided back pain. |
-| HISTORY: Intubation with pulmonary edema. |
-| COMPARISON: Chest radiographs on ___ and CT chest without contrast on ___. |
-| TECHNIQUE: Portable upright chest radiograph. |
-| FINDINGS: In comparison with the prior study, there are diffuse bilateral pulmonary opacifications, more prominent on the right.
-These findings could indicate severe pulmonary edema, but superimposed pneumonia or developing ARDS cannot be excluded.
-Monitoring and support devices are appropriately positioned.
-|
-| Prompt Content |
-| [System prompt] { |
-| The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes. |
-| Image Representation Placeholder |
-| [Clinical prompt] { |
-| Provide a detailed description of the findings in the radiology image. Following clinical context: |
-| Indication: Dyspnea and cough, right-sided back pain. |
-| History: Intubation with pulmonary edema. |
-| Comparison: Chest radiographs on ___ and CT chest without contrast on ___. |
-| Technique: Portable upright chest radiograph. |
-Appendix C Appendix C
-C.1 Impact of the Temporal Alignment Connector under General-Domain Pre-trained Models
+As shown in Table 6, the differences in lexical (ROUGE-L (Lin, 2004)) and clinical (RadGraph-F1 (Jain et al., 2021)) metrics between the two candidates are relatively smaller compared to the score.
+This demonstrates that Temporal Entity F1 effectively captures and evaluates the quality of temporal information in radiology reports, distinguishing it more accurately than other standard metrics in the context of temporal information descriptions.
+Appendix G Analysis of Concurrent Work and Non-LLM-based Models
+G.1 Discussion on Performance with Radiology Foundation Models
+| Metric | RaDialog | MedVersa | MAIRA-2 | M4CXR | Libra (%) |
+| Lexical: | |||||
+| ROUGE-L | 31.6 | – | 38.4 | 28.5 | 36.7 (-4.4%) |
+| BLEU-1 | 39.2 | – | 46.5 | 33.9 | 51.3 (10.3%) |
+| BLEU-4 | 14.8 | 17.8 | 23.4 | 10.3 | 24.5 (4.7%) |
+| METEOR | – | – | 42.0 | – | 48.9 (16.4%) |
+| BERTScore | – | 49.7 | – | – | 62.5 (25.8%) |
+| Clinical: | |||||
+| RadGraph-F1 | – | 28.0 | 34.6 | 21.8 | 32.9 (-4.9%) |
+| RG | – | – | 39.7 | – | 37.6 (-5.3%) |
+| RadCliQ() | – | 2.7 | 2.6 | – | 2.7 (-3.8%) |
+| CheXbert vector | – | 46.4 | 50.6 | – | 46.9 (-7.3%) |
+| CheXpert-F1: | |||||
+| Micro-F1-14 | 39.2 | – | 58.5 | 60.6 | 55.9 (-7.8%) |
+| Macro-F1-14 | – | – | 42.7 | 40.0 | 40.4 (-5.4%) |
+| Micro-F1-5 | – | – | 58.9 | 61.8 | 60.1 (-2.8%) |
+| Macro-F1-5 | – | – | 51.5 | 49.5 | 53.8 (4.5%) |
+As shown in Table 7, these models belong to the category of radiology foundation models.
+DaDialog (Pellegrini et al., 2023) is a conversational MLLM designed for a broad range of dialogue-based medical assistance tasks.
+To enhance structured findings extraction, it employs the publicly available CheXbert model (Smit et al., 2020b) to extract symptom labels from scans, facilitating a structured representation of findings.
+MedVersa (Zhou et al., 2024) and M4CXR (Park et al., 2024) support a diverse set of tasks, including medical report generation, visual grounding, and visual question answering.
+These models aim to provide general-purpose multimodal medical assistance by leveraging vision-language pre-training strategies.
+MAIRA-2 (Bannur et al., 2024) specialises in grounded radiology report generation, which differs from traditional report generation tasks by requiring explicit image-level localization of findings and symptoms.
+Grounded radiology reporting, as defined by Bannur et al.
+(2024), structures the report as a list of sentences, where each sentence: (1) is linked to zero or more spatial image annotations, and (2) describes at most a single finding from an image.
+To support this task, MAIRA-2 introduces a custom dataset, explicitly designed to provide structured annotations aligning textual descriptions with spatial regions of interest in radiological images.
+This approach contrasts with conventional RRG models that generate unstructured free-text reports.
+It is worth noting that the inference sets differ slightly across these models.
+Additionally, all these models leverage supplementary radiology information, such as lateral view scans, prior study reports, or both (as detailed in Appx.
+A.2), to enhance their performance in radiology-related tasks.
+Despite these considerations, Libra achieves the highest scores on most lexical metrics, including BLEU-, METEOR, and BERTScore, while trailing slightly behind MAIRA-2 on ROUGE-L.
+In clinical metrics, Libra predominantly ranks second, just behind the best-performing model.
+For clinical metrics, Libra consistently ranks second, just behind the top-performing model.
+In metrics that evaluate medical entities and their relationships, such as RadGraph-F1, RG, and RadCliQ, Libra also ranks second.
+Similarly, Libra comes second in the CheXbert vector embedding score.
+However, in the CheXpert metrics, Libra ranks first in Macro-F1 for the 5-class subset, with only a slight dip in the Micro-F1 score for the 14-class subset.
+Incorporating lateral images and prior study reports could enhance clinical scores.
+Additionally, strategies like chain-of-thought reasoning and grounded report generation further improve performance in RRG tasks.
+Looking ahead, we plan to develop model architectures that can automatically adapt to multiple tasks and diverse scenarios, enabling more efficient handling of additional radiological information.
+G.2 Discussion on Performance with non-LLM-based Models
+| Model | Lexical Metrics | Clinical Metrics | |||||||
+| B-1 | B-2 | B-3 | B-4 | MTR | R-L | P | R | F1 | |
+| ST | 29.9 | 18.4 | 12.1 | 8.4 | 12.4 | 26.3 | 24.9 | 20.3 | 20.4 |
+| ATT2IN | 32.5 | 20.3 | 13.6 | 9.6 | 13.4 | 27.6 | 32.3 | 23.9 | 20.4 |
+| ADAATT | 29.9 | 18.5 | 12.4 | 8.8 | 11.8 | 26.6 | 26.8 | 18.6 | 18.1 |
+| TopDown | 31.7 | 19.5 | 13.0 | 9.2 | 12.8 | 26.7 | 32.0 | 23.1 | 23.8 |
+| R2Gen | 35.3 | 21.8 | 14.5 | 10.3 | 14.2 | 27.0 | 33.3 | 27.3 | 27.6 |
+| R2GenCMN | 35.3 | 21.8 | 14.8 | 10.6 | 14.2 | 27.8 | 34.4 | 27.5 | 27.8 |
+| XPRONET | 34.4 | 21.5 | 14.6 | 10.5 | 13.8 | 27.9 | – | – | – |
+| CMCL | 34.4 | 21.7 | 14.0 | 9.7 | 13.3 | 28.1 | – | – | – |
+| PPKED | 36.0 | 22.4 | 14.9 | 10.6 | 14.9 | 28.4 | – | – | – |
+| AlignTransformer | 37.8 | 23.5 | 15.6 | 11.2 | 15.8 | 28.3 | – | – | – |
+| CA | 35.0 | 21.9 | 15.2 | 10.9 | 15.1 | 28.3 | 35.2 | 29.8 | 30.3 |
+| LKBMA | 38.6 | 23.7 | 15.7 | 11.1 | – | 27.4 | 42.0 | 33.9 | 35.2 |
+| M2TR | 37.8 | 23.2 | 15.4 | 10.7 | 14.5 | 27.2 | 24.0 | 42.8 | 30.8 |
+| KnowMAT | 36.3 | 22.8 | 15.6 | 11.5 | – | 28.4 | 45.8 | 34.8 | 37.1 |
+| RAMT | 36.2 | 22.9 | 15.7 | 11.3 | 15.3 | 28.4 | 38.0 | 34.2 | 33.5 |
+| CMM-RL | 38.1 | 23.2 | 15.5 | 10.9 | 15.1 | 28.7 | 34.2 | 29.4 | 29.2 |
+| CMCA | 36.0 | 22.7 | 15.6 | 11.7 | 14.8 | 28.7 | 44.4 | 29.7 | 35.6 |
+| KiUT | 39.3 | 24.3 | 15.9 | 11.3 | 16.0 | 28.5 | 37.1 | 31.8 | 32.1 |
+| DCL | – | – | – | 10.9 | 15.0 | 28.4 | 47.1 | 35.2 | 37.3 |
+| MMTN | 37.9 | 23.8 | 15.9 | 11.6 | 16.1 | 28.3 | – | – | – |
+| METrans | 25.0 | 16.9 | 12.4 | 15.2 | – | 29.1 | 36.4 | 30.9 | 31.1 |
+| ORGAN | 38.6 | 25.6 | 17.2 | 12.3 | 16.2 | 29.3 | 41.6 | 41.8 | 38.5 |
+| COMG | 36.3 | 23.5 | 16.7 | 12.4 | 12.8 | 29.0 | – | – | – |
+| MedM2G | 41.2 | 26.9 | 17.9 | 14.2 | – | 30.9 | – | – | – |
+| CvT2DistilGPT2 | 39.2 | 24.5 | 16.9 | 12.4 | 15.3 | 28.5 | 35.9 | 41.2 | 38.4 |
+| RGRG | 37.3 | 24.9 | 17.5 | 12.6 | 16.8 | 26.4 | 46.1 | 47.5 | 44.7 |
+| BioViL-T | – | – | – | 9.2 | – | 29.6 | – | – | 17.5 |
+| VLCI | 40.0 | 24.5 | 16.5 | 11.9 | 15.0 | 28.0 | 48.9 | 34.0 | 40.1 |
+| TiBiX | 32.4 | 23.4 | 18.5 | 15.7 | 16.2 | 33.1 | 30.0 | 22.4 | 25.0 |
+| RECAP | 42.9 | 26.7 | 17.7 | 12.5 | 16.8 | 28.8 | 38.9 | 44.3 | 39.3 |
+| MS-TF | 43.6 | 27.5 | 18.4 | 12.9 | 17.7 | 30.5 | – | – | 41.1 |
+| Libra | 51.3 | 38.0 | 30.0 | 24.5 | 48.9 | 36.7 | 59.7 | 52.5 | 55.9 |
+To facilitate comparison with non-LLM-based models, we selected evaluation metrics commonly used in these studies.
+These include BLEU- (Papineni et al., 2002), METEOR (MTR) (Banerjee and Lavie, 2005), and ROUGE-L (R-L) (Lin, 2004).
+For clinical metrics, we adopted the CheXbert (Irvin et al., 2019), reporting Precision (P), Recall (R), and F1.
+Baseline
+For performance evaluation, we compare our model with the following baselines: ST (Vinyals et al., 2015), ATT2IN (Rennie et al., 2017), ADAATT (Lu et al., 2017), TopDown (Anderson et al., 2018), R2Gen (Chen et al., 2020), R2GenCMN (Chen et al., 2021), M2TR (Nooralahzadeh et al., 2021), CMCL (Liu et al., 2021a),PPKED (Liu et al., 2021b), AlignTransformer (You et al., 2021), CA (Liu et al., 2021c), LKBMA (Yang et al., 2022b), KnowMAT (Yang et al., 2022a), XPRONET (Wang et al., 2022), CMM-RL (Qin and Song, 2022), RAMT (Zhang et al., 2024b), CMCA (Song et al., 2022), KiUT (Huang et al., 2023), DCL (Li et al., 2023b), MMTN (Cao et al., 2023), METrans (Wang et al., 2023), ORGAN (Hou et al., 2023b), COMG (Gu et al., 2023), BioViL-T (Bannur et al., 2023), RGRG (Tanida et al., 2023), RECAP (Hou et al., 2023a), CvT2DistilGPT2 (Nicolson et al., 2023), VLCI (Chen et al., 2024a), TiBiX (Sanjeev et al., 2024), MedM2G (Zhan et al., 2024), MS-TF (Mei et al., 2024).
+To ensure fairness, Libra also utilizes prior images, aligning with other models that leverage prior images or additional information.
+As demonstrated in Table 8, Libra, similar to other LLM-based models, consistently outperforms non-LLM-based models.
+This advantage is largely attributed to advancements in LLMs and visual instruction tuning (Liu et al., 2023), enabling multimodal large language models (MLLMs) to achieve superior performance in RRG tasks.
+Appendix H Additional Ablation Studies
+H.1 Impact of Temporal Information on Libra in RRG
+| Metric | Libra | |
+| w/o prior | w/ prior (%) | |
+| Lexical: | ||
+| ROUGE-L | 36.17 | 36.66 (+1.35%) |
+| BLEU-1 | 51.20 | 51.25 (+0.10%) |
+| BLEU-4 | 24.33 | 24.54 (+0.86%) |
+| METEOR | 48.69 | 48.90 (+0.43%) |
+| BERTScore | 61.94 | 62.50 (+0.90%) |
+| F1 | 32.72 | 35.34 (+8.00%) |
+| Clinical: | ||
+| RadGraph-F1 | 32.42 | 32.87(+1.39%) |
+| RG | 36.92 | 37.57(+1.76%) |
+| RadCliQ() | 2.76 | 2.72 (+1.45%) |
+| CheXbert vector | 46.31 | 46.85 (+1.17%) |
+| CheXpert-F1: | ||
+| Micro-F1-14 | 55.25 | 55.87 (+1.12%) |
+| Macro-F1-14 | 40.15 | 40.38(+0.57%) |
+| Micro-F1-5 | 58.93 | 60.07(+1.93%) |
+| Macro-F1-5 | 52.61 | 53.75(+2.17%) |
+Temporal information is embedded in paired images and referenced in the corresponding radiology reports, capturing changes over time through references to prior symptoms and their progression, as discussed in Appx.
+B.1.
+As shown in Table 5, 86% of the test data includes prior images, providing a solid foundation for evaluating the impact of temporal information.
+During training, Libra integrates the ability to perceive and utilise temporal information into its architecture.
+To evaluate whether Libra effectively leverage temporal information during inference, we assess its performance using prior images when available as references to determine their impact on the overall capability.
+In Table 9, the inclusion of prior images substantially enhances Libra’s performance across all metrics.
+Notably, clinical scores exhibit greater improvements compared to lexical scores, underscoring the importance of temporal information in generating high-quality medical reports beyond merely improving linguistic fluency.
+The F1 score shows the most substantial improvement, with an increase of 8%, highlighting Libra’s capability to effectively leverage temporal changes provided by prior images.
+These results validate the role of temporal information in enhancing the quality of the generated Findings section and improving Libra’s overall performance in RRG tasks.
+H.2 Impact of the Temporal Alignment Connector under General-Domain Pre-trained Models
 | Metric | Libra-b | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
@@
 | Macro-F1-5 | 47.24 | 44.28 (-6.27%) | 44.39 (-6.04%) | 44.68 (-5.42%) | 43.48 (-7.96%) |
-Domain-specific pre-trained models (i.e., RAD-DINO (Pérez-García et al., 2024) and Meditron (Chen et al., 2023b)) inherently incorporate domain-specific knowledge, such as phrasing conventions, pronoun usage, and even temporal information embedded in the training corpus.
-To eliminate the influence of such factors, we used a general-domain image encoder (DINOv2 (Oquab et al., 2024)) and a LLM (Vicuna-7B-v1.5 (Chiang et al., 2023)), allowing the structural enhancements of TAC to be observed more directly.
-We adopted the same setup as the second ablation study in Section 3.
+Domain-specific pre-trained models (i.e., RAD-DINO (Pérez-García et al., 2024) and Meditron (Chen et al., 2023c)) inherently incorporate domain-specific knowledge, such as phrasing conventions, pronoun usage, and even temporal information embedded in the training corpus.
+To isolate the structural impact of TAC, we used a general-domain image encoder (DINOv2 (Oquab et al., 2024)) and a LLM (Vicuna-7B-v1.5 (Chiang et al., 2023)), allowing the structural enhancements of TAC to be observed more directly.
+We replicated the first ablation setup from Sec.
+4.
 We first conducted a baseline experiment, referred to as Libra-b, by fine-tuning only the adapter for the Findings generation task.
-As shown in Table 8, we then conducted ablation studies by sequentially removing different components from the model, including the TFM, LFE, Prior Image Prefix Bias (PIPB) in TFM, and the entire TAC.
-Removing TFM restricts the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023), but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions.
-It is important to note that without the TFM module, the model cannot process true prior images or dummy prior images, and is limited to only the current image as input.
+As shown in Table 10, we then conducted ablation studies by sequentially removing different components from the model, including the Temporal Fusion Module (TFM), Layerwise Feature Extractor (LFE), Prior Image Prefix Bias (PIPB), and the entire TAC.
+Removing TFM restricts the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023), but with a four-layer MLP to align the image feature with the LLM’s hidden dimensions.
+Notably, without TFM, the model cannot process prior images or dummy prior images, and is limited to only the current image as input.
 Without LFE, the model follows the LLaVA setup, using the penultimate layer of the image encoder to process single or paired images.
-The ablation results are consistent with those observed when using domain-specific pre-trained models, as presented in Section 3.
-Compared to Libra-b, removing any TAC submodule results in a decline across all metrics.
-Removing the TFM caused a notable drop in the F1 score (4%), emphasising its role in capturing temporal information.
-The removal of the LFE significantly decreased RadGraph-related scores, highlighting its importance for extracting detailed image features.
-Removing the PIPB impacted clinical scores more than lexical scores.
-Finally, removing the TAC resulted in significant declines across all metrics, underscoring its critical role in integrating image details and temporal information.
-C.2 Impact of the Temporal Alignment Connector After the Second Stage
+The ablation results are consistent with those observed using domain-specific models, as presented in Table 2.
+Removing any TAC submodule led to declines across all metrics.
+Specifically, removing TFM caused a notable drop in the F1 score (4%), emphasising its role in capturing temporal information.
+The absence of LFE significantly reduced RadGraph-related scores, demonstrating its importance for detailed image feature extraction.
+PIPB removal primarily impacted clinical metrics, while removing the entire TAC resulted in substantial declines across all metrics.
+These findings reaffirm the critical role of TAC in integrating image details and temporal information effectively.
+H.3 Impact of the Temporal Alignment Connector After the Second-Stage Fine-tuning
 | Metric | Libra-2 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
@@
 | Macro-F1-5 | 51.99 | 51.16 (-1.60%) | 51.19 (-1.54%) | 51.27 (-1.38%) | 50.87 (-2.15%) |
-To further evaluate the impact of the Temporal Alignment Connector (TAC) on Libra’s performance, we followed the setup of the second ablation study in Section 3.
+To further evaluate the impact of the Temporal Alignment Connector (TAC) on Libra’s performance, we followed the setup of the first ablation study in Sec.
+4.
 After the first stage of alignment, the model underwent a second stage of fine-tuning.
 This stage was designed to optimise the model’s performance on the Findings section generation task by leveraging the aligned visual and textual features learned during the initial stage.
-In this phase, we applied Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM (Meditron (Chen et al., 2023b)), while keeping the visual encoder (RAD-DINO (Pérez-García et al., 2024)) and TAC weights frozen.
+In this phase, we applied Low-Rank Adaptation (LoRA) Hu et al.
+(2021) to fine-tune the pre-trained LLM (Meditron Chen et al.
+(2023c)), while keeping the visual encoder (RAD-DINO Pérez-García et al.
+(2024)) and TAC weights frozen.
 The baseline for this experiment is Libra-2 (in Table 11), which is derived from Libra-1 (in Table 2) after undergoing LoRA fine-tuning.
 We conducted ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC.
-The conclusions remain consistent with those in Section 3, demonstrating that removing any TAC submodule results in a decline across all metrics compared to Libra-2.
-This indicates that the performance improvements brought by TAC are stable and unaffected by changes in training stages.
-It further confirms that TAC has effectively embedded the capability to process temporal information within the model.
-C.3 Robustness Assessment of the Temporal Alignment Connector
+Results consistently showed declines across all metrics compared to Libra-2, mirroring the trends observed in Sec.
+4.
+This reinforces that the performance improvements brought by TAC are stable and unaffected by changes in training stages.
+It further confirms that TAC has embedded the capability to process temporal information within the model.
+H.4 Robustness Evaluation of the Temporal Alignment Connector
+To evaluate the robustness of the Temporal Alignment Connector (TAC), we introduced an additional round of LoRA fine-tuning to induce overtraining.
+Following the setup in Appx.
+H.3, after integrating the first LoRA weights, a new set of LoRA adapters was reinitialised for the LLM and trained for one epoch under the same second-stage fine-tuning configuration.
+The baseline for this experiment is Libra-3 (as shown in Table 12), which is derived from Libra-2 (illustrated in Table 11) following this additional fine-tuning step.
 | Metric | Libra-3 | w/o TFM | w/o LFE | w/o PIPB | w/o TAC |
@@
 | Macro-F1-5 | 44.58 | 44.34 (-0.54%) | 44.35 (-0.52%) | 44.37 (-0.46%) | 44.26 (-0.72%) |
-In this experiment, we introduced an additional round of LoRA fine-tuning to induce overtraining and assess the robustness of the Temporal Alignment Connector (TAC).
-Following the setup in Appx.
-C.2, after integrating the first LoRA weights, we reinitialised a new set of LoRA adapters for the LLM and conducted further training under the same second-stage fine-tuning configuration for one epoch.
-The baseline for this experiment is Libra-3 (as shown in Table 10), which is derived from Libra-2 (illustrated in Table 9) following this additional fine-tuning step.
-The results reveal that, compared to Libra-2, Libra-3 demonstrates minimal changes in lexical scores, while clinical scores decline due to overfitting caused by the additional fine-tuning.
-Notably, the CheXpert (Smit et al., 2020) (Macro-F1-[5/14]) scores exhibit the most influential reduction.
-However, the ablation studies reaffirm that the performance improvements introduced by TAC are robust and remain unaffected by variations in training strategies.
-This robustness is attributed to TAC’s ability to effectively capture and preserve temporal image representations during the first training phase, which are subsequently preserved throughout further fine-tuning.
-These findings highlight that TAC reliably embeds the capacity for temporal information processing into the model, making it a stable and essential component for handling temporal data in RRG tasks.
-C.4 Impact of Radiology-Specific Pre-trained Models
+The results reveal that, compared to Libra-2, Libra-3 exhibits minimal changes in lexical scores, while clinical scores decline due to overfitting caused by the additional fine-tuning.
+Notably, the CheXpert (Smit et al., 2020a) (Macro-F1-[5/14]) scores exhibit the most influential reduction.
+Despite this decline, ablation studies confirm that TAC’s performance improvements remain robust, unaffected by variations in training strategies.
+This resilience stems from TAC’s ability to capture and retain temporal image representations during the initial training phase, which are preserved through subsequent fine-tuning.
+These findings underscore TAC’s reliability as a critical component for temporal information processing in RRG tasks. It ensures stability even under diverse training conditions.
+H.5 Impact of Radiology-Specific Pre-trained Models on Libra
 | Metric | Libra-1 | w/o RadDINO | w/o Meditron | w/o RadDINO+Meditron |
@@
 Aligning radiology images with textual information is a key challenge in RRG tasks.
-To demonstrate the benefits of using radiology-specific pre-trained models for more accurate feature representation and improved MLLM performance, we initialised a Libra model with RadDINO, the TAC, and Meditron-7b, conducting the first stage of training, denoted as Libra-1 (This is consistent with the baseline setup of the second ablation study in Section 3).
+To demonstrate the benefits of using radiology-specific pre-trained models for more accurate feature representation and improved MLLM performance, we initialised a Libra model with RadDINO, the TAC, and Meditron-7b, conducting the first stage of training, denoted as Libra-1 (This is consistent with the baseline setup of the previous ablation study in Sec.
+4).
 Then we replaced the image encoder and LLM with their general-domain counterparts, DINOv2 and Vicuna-7B-v1.5, respectively.
-Finally, we replaced both components, which is also referred to as Libra-b.
+Finally, we replaced both components, which is also referred to as Libra-b (in Table 10).
 As shown in Table 13, substituting radiology-specific pre-trained models with general-domain models resulted in a notable decline in clinical scores, while the impact on lexical scores was minimal.
@@
 This suggests that accurate medical image representation provides greater benefits in RRG tasks, indicating the importance of incorporating domain-specific knowledge into pre-trained models to enhance Libra’s performance.
-C.5 Incremental Study
-In this section, we conducted an incremental study of our model, mirroring the logical framework we used to construct Libra.
-We started with an architecture similar to LLaVA, comprising a pre-trained CLIP image encoder, a randomly initialised four-layer MLP as the adapter, and Vicuna-7B-v1.5 as the LLM.
-We train the adapter on the Findings section generation task to establish an initial baseline.
-Gradual enhancements are then introduced, as shown in the Table 12.
-First, the image encoder is replaced with DINOv2.
-Building upon this, we integrate the LFE, the prefix module of the TAC connector.
-Subsequently, we add the TFM, the suffix module, thus completing the full TAC connector.
-Following this, the image encoder and LLM are progressively replaced with biomedical domain models, RAD-DINO and Meditron, respectively.
-The dataset used for pretraining in the first stage is then extended.
-Finally, the model is fine-tuned for downstream tasks to obtain Libra.
+H.6 Incremental Component Analysis
 | Metric | Stage 1: Temporal Feature Alignment | Stage 2 | ||||||
@@
 | Macro-F1-5 | 36.55 | 43.48 | 44.39 | 47.24 | 47.62 | 49.33 | 40.40 | 53.75 |
-We observed that with each step of improvement, the overall performance consistently exceeded the previous state, demonstrating the essential role of each component in our optimisation scheme.
-Notably, in the alignment stage, the model showed the greatest overall improvement after the addition of the TFM, highlighting its capability to capture temporal information, which significantly enhanced performance on the RRA task.
-However, after the data expansion optimisation in the first stage, the scores on lexical metrics improved, while clinical metrics exhibited a slight decline.
-This can be attributed to the introduction of the VQA task, which led the model to perceive more fine-grained grounded information.
-This is distinct from the report generation task, which focuses on describing the holistic level.
-Furthermore, this shift also affects the F1 score, as temporal entities are often associated with the presence of specific symptoms.
-Ultimately, these dropped scores are improved again through the second stage of fine-tuning.
-Evaluation of Libra’s Temporal Awareness.
-Another approach to investigating the model’s ability to capture temporal information is to evaluate it separately within the test split based on the presence or absence of true prior images, as shown in Table 13.
+We conducted an incremental study to evaluate the effectiveness of each component in Libra’s architecture.
+Starting with a baseline model similar to LLaVA—comprising a pre-trained CLIP image encoder, a randomly initialised four-layer MLP adapter, and Vicuna-7B-v1.5 as the LLM—we trained the adapter on the Findings section generation task.
+Improvements were introduced incrementally, as summarised in Table 14.
+First, we replaced the image encoder with DINOv2.
+Next, we incorporated the LFE (prefix module of TAC) and subsequently added the TFM (suffix module), completing the TAC connector.
+We then replaced the image encoder and LLM with RAD-DINO and Meditron, respectively.
+The dataset for the first stage was expanded, and final fine-tuning was conducted for downstream tasks to produce Libra.
+With each enhancement, the model’s performance improved, demonstrating the critical role of each component.
+Notably, the addition of the TFM during the alignment stage provided the most significant improvement, showcasing its ability to capture temporal information, which is essential for the RRG task.
+However, data expansion in the first stage led to improved lexical scores but a slight decline in clinical metrics, likely due to the VQA task’s focus on fine-grained grounded information rather than holistic report generation, as mentioned in Sec.
+4.
+This shift also affected the F1 score, as temporal entities are often linked to specific symptoms.
+These declines were subsequently addressed through second-stage fine-tuning, resulting in overall improved performance.
+Evaluation of Libra’s Temporal Awareness
+Another approach to investigating the model’s ability to capture temporal information is to evaluate it separately within the test split based on the presence or absence of prior images, as shown in Table 15.
 With the addition of the TFM, the model exhibited temporal awareness.
@@
 This indicates that the structural enhancements have resulted in a sustained improvement in the model’s temporal perception capabilities.
-An effective example is in Section 5.
-Appendix D Appendix D
-D.1 Discussion on Performance with Radiology Foundation Models
-| Metric | MedVersa | MAIRA-2 | M4CXR | Libra (%) |
-| Lexical: | ||||
-| ROUGE-L | – | 38.4 | 28.5 | 36.7 (-4.4%) |
-| BLEU-1 | – | 46.5 | 33.9 | 51.3 (10.3%) |
-| BLEU-4 | 17.8 | 23.4 | 10.3 | 24.5 (4.7%) |
-| METEOR | – | 42.0 | – | 48.9 (16.4%) |
-| BERTScore | 49.7 | – | – | 62.5 (25.8%) |
-| Clinical: | ||||
-| RadGraph-F1 | 28.0 | 34.6 | 21.8 | 32.9 (-4.9%) |
-| RG | – | 39.7 | – | 37.6 (-5.3%) |
-| RadCliQ() | 2.7 | 2.6 | – | 2.7 (-3.8%) |
-| CheXbert vector | 46.4 | 50.6 | – | 46.9 (-7.3%) |
-| CheXpert-F1: | ||||
-| Micro-F1-14 | – | 58.5 | 60.6 | 55.9 (-7.8%) |
-| Macro-F1-14 | – | 42.7 | 40.0 | 40.4 (-5.4%) |
-| Micro-F1-5 | – | 58.9 | 61.8 | 60.1 (-2.8%) |
-| Macro-F1-5 | – | 51.5 | 49.5 | 53.8 (4.5%) |
-As shown in Table 14, these models belong to the category of radiology foundation models.
-MedVersa (Zhou et al., 2024) and M4CXR (Park et al., 2024) support multiple tasks, including medical report generation, visual grounding, and visual question answering.
-In contrast, MAIRA-2 (Bannur et al., 2024) specialises in grounded radiology report generation, a novel task requiring image-level localisation for each identifiable finding or symptom.
-It is worth noting that the test sets differ slightly.
-All of these models utilise additional radiology information, such as lateral views, prior study reports of CXRs, or both.
-Despite these considerations, Libra achieves the highest scores on most lexical metrics, including BLEU-, METEOR, and BERTScore, while trailing slightly behind MAIRA-2 on ROUGE-L.
-In clinical metrics, Libra predominantly ranks second, just behind the best-performing model.
-For clinical metrics, Libra consistently ranks second, just behind the top-performing model.
-In metrics that evaluate medical entities and their relationships, such as RadGraph-F1, RG, and RadCliQ, Libra also ranks second.
-Similarly, Libra comes second in the CheXbert vector embedding score.
-However, in the CheXpert metrics, Libra ranks first in Macro-F1 for the 5-class subset, with only a slight dip in the Micro-F1 score for the 14-class subset.
-Incorporating lateral images and prior study reports could enhance clinical scores.
-Additionally, strategies like chain-of-thought reasoning and grounded report generation further improve performance in RRG tasks.
-Looking ahead, we plan to develop model architectures that can automatically adapt to multiple tasks and diverse scenarios, enabling more efficient handling of additional radiological information.
-D.2 Discussion on Performance with non-LLM-based Models
-| Model | Lexical Metrics | Clinical Metrics | |||||||
-| B-1 | B-2 | B-3 | B-4 | MTR | R-L | P | R | F1 | |
-| R2Gen | 35.3 | 21.8 | 14.5 | 10.3 | 14.2 | 27.0 | 33.3 | 27.3 | 27.6 |
-| R2GenCMN | 35.3 | 21.8 | 14.8 | 10.6 | 14.2 | 27.8 | 34.4 | 27.5 | 27.8 |
-| M2TR | 37.8 | 23.2 | 15.4 | 10.7 | 14.5 | 27.2 | 24.0 | 42.8 | 30.8 |
-| KnowMAT | 36.3 | 22.8 | 15.6 | 11.5 | – | 28.4 | 45.8 | 34.8 | 37.1 |
-| CMM-RL | 38.1 | 23.2 | 15.5 | 10.9 | 15.1 | 28.7 | 34.2 | 29.4 | 29.2 |
-| CMCA | 36.0 | 22.7 | 15.6 | 11.7 | 14.8 | 28.7 | 44.4 | 29.7 | 35.6 |
-| KiUT | 39.3 | 24.3 | 15.9 | 11.3 | 16.0 | 28.5 | 37.1 | 31.8 | 32.1 |
-| DCL | – | – | – | 10.9 | 15.0 | 28.4 | 47.1 | 35.2 | 37.3 |
-| METrans | 25.0 | 16.9 | 12.4 | 15.2 | – | 29.1 | 36.4 | 30.9 | 31.1 |
-| ORGAN | 38.6 | 25.6 | 17.2 | 12.3 | 16.2 | 29.3 | 41.6 | 41.8 | 38.5 |
-| COMG | 36.3 | 23.5 | 16.7 | 12.4 | 12.8 | 29.0 | – | – | – |
-| MedM2G | 41.2 | 26.9 | 17.9 | 14.2 | – | 30.9 | – | – | – |
-| BioViL-T | – | – | – | 9.2 | – | 29.6 | – | – | 17.5 |
-| TiBiX | 32.4 | 23.4 | 18.5 | 15.7 | 16.2 | 33.1 | 30.0 | 22.4 | 25.0 |
-| RECAP | 42.9 | 26.7 | 17.7 | 12.5 | 16.8 | 28.8 | 38.9 | 44.3 | 39.3 |
-| Libra | 51.3 | 38.0 | 30.0 | 24.5 | 48.9 | 36.7 | 59.7 | 52.5 | 55.9 |
-To facilitate comparison with non-LLM-based models, we selected evaluation metrics commonly used in these studies.
-These include BLEU- (Papineni et al., 2002), METEOR (MTR) (Banerjee & Lavie, 2005), and ROUGE-L (R-L) (Lin, 2004).
-For clinical metrics, we adopted the CheXbert (Irvin et al., 2019a), reporting Precision (P), Recall (R), and F1.
-Baseline For performance evaluation, we compare our model with the following baselines: R2Gen (Chen et al., 2021), R2GenCMN (Chen et al., 2021), M2TR (Nooralahzadeh et al., 2021), KnowMAT (Yang et al., 2022), CMM-RL (Qin & Song, 2022), CMCA (Song et al., 2022), KiUT (Huang et al., 2023), DCL (Li et al., 2023b), METrans (Wang et al., 2023b), ORGAN (Hou et al., 2023b), COMG (Gu et al., 2023), MedM2G (Zhan et al., 2024), BioViL-T (Bannur et al., 2023), TiBiX (Sanjeev et al., 2024), RECAP (Hou et al., 2023a).
-As the results demonstrate, our model, like other LLM-based models, generally outperforms non-LLM-based models.
-This is attributed to advancements in LLMs and visual instruction tuning (Liu et al., 2023), enabling multimodal large language models (MLLMs) to excel in the RRG task.
-Appendix E Appendix E
-E.1 Temporal Alignment Heatmap Analysis and Feature Representation
-Current image
-RAD-DINO layer 1
-RAD-DINO layer 12
-LFE
-TAC
-The heatmap in Figure 4 corresponds to the example in Figure 2, where no true prior image was used as a reference.
-It highlights the clear differences in feature representations across layers of the RAD-DINO image encoder.
-The shallow layers primarily capture the overall lung structure, while the deeper layers focus on specific disease locations.
-After being processed by the Layerwise Feature Extractor (LFE), the image feature representations show that larger symptom regions are assigned higher weights.
-This indicates that the image representation achieves a more fine-grained level of detail.
-Following the Temporal Alignment Connector (TAC), Libra integrates the weighted dummy prior image, resulting in a more uniform feature distribution.
-This reflects temporal information, indicating no significant changes compared to the prior study.
-This process helps the model provide a smoother image feature representation for subsequent processing by the text decoder (i.e., the LLM).
-Current image
-Current LFE
-TAC-1
-Prior LFE
-Prior image
-Swap order
-Current image (Prior)
-Current LFE (Prior)
-TAC-2
-Prior LFE (Current)
-Prior image (Current)
-The heatmap in Figure 5 corresponds to the example in Figure 3, where a true prior image is provided.
-As observed, after processing through the LFE, the model captures higher-granularity feature representations in the symptom areas.
-After being processed by the TAC, the intermediate representation not only integrates these fine-grained features but also incorporates the difference information between the two images, reflecting the temporal information provided by the reference image, as shown in TAC-1 (Figure 5).
-When the image order is swapped, treating the true prior image as the current image, the intermediate output from the LFE remains unchanged.
-Nevertheless, a comparison of the TAC-2 and TAC-1 outputs reveals noticeable differences in the lung feature representations.
-This demonstrates that the model’s temporal perception is directional and that the TAC module successfully integrates temporal information from different time points.
-At the same time, it confirms that the LFE focuses solely on extracting image feature representations without encoding temporal aspects.
-This result is expected because the residual connections in the TAC ensure that the current image remains the main modality, while the prior image serves as the auxiliary modality.
-When the order is swapped, the prior image becomes the main modality for subsequent processing.
-Thus, swapping the image order also changes the content of the generated report, as mentioned in Section 5.
-The primary difference lies in the reversal of the temporal state of symptoms, such as changing from “improving” to “worsening”.
+An effective example is in Sec.
+5.
+Appendix I Heatmap Analysis and Temporal Feature Representation
+The heatmap in Figure 4 corresponds to the example in (a) of Figure 3, where no prior image was used as a reference.
+It illustrates the clear differences in feature representations across layers of the RAD-DINO (Pérez-García et al., 2024) image encoder.
+The shallow layers primarily capture the overall lung structure, while the deeper layers focus on specific disease regions.
+After passing through the Layerwise Feature Extractor (LFE), the image feature representations assign higher weights to larger symptom regions, achieving finer granularity.
+Following the Temporal Alignment Connector (TAC), the model integrates the weighted dummy prior image, producing a uniform feature distribution that reflects temporal information.
+This indicates no significant changes compared to the prior study and facilitates smoother image feature representations for downstream text generation by the LLM.
+The heatmap in Figure 5 corresponds to the example in (b) of Figure 3, where a prior image is provided.
+After processing through the LFE, the model captures fine-grained feature representations in symptom areas.
+When processed by the TAC, these features are integrated with the differences between the two images, effectively reflecting temporal information, as demonstrated in TAC-1 (top) of Figure 5.
+When the image order is swapped, treating the prior image as the current image, the LFE output remains unchanged.
+However, comparing TAC-2 (bottom of Figure 5) and TAC-1 outputs reveals significant differences in lung feature representations.
+This highlights the model’s directional temporal perception and confirms that the TAC module effectively encodes temporal information from different time points, while the LFE focuses solely on image features without temporal encoding.
+This behaviour aligns with the design of the TAC, where residual connections prioritise the current image as the main modality and the prior image as the auxiliary.
+Swapping the image order changes the main modality, altering the temporal state of symptoms in the generated report, such as reversing descriptions from “improving” to “worsening,” as discussed in Sec.
+5.
+Appendix J Extended Discussion on Limitations
+While our work represents a step forward in leveraging temporal information for radiology report generation, it also has several limitations that warrant further exploration.
+Handling Multiple Prior Scans
+Our current model is designed to process a single prior scan alongside the current scan.
+While this approach aligns with standard clinical workflows, which typically prioritise the most recent prior study for comparisons, it overlooks scenarios where multiple prior scans could offer a richer temporal perspective.
+For instance, analysing a sequence of images spanning an extended period could provide deeper insights into gradual disease progression.
+Future efforts should focus on extending our framework to incorporate multiple prior scans efficiently, enabling a more nuanced understanding of temporal patterns in clinical data.
+Temporal Information Beyond Image Comparisons
+Currently, our model captures temporal information through paired image comparisons and corresponding textual reports.
+However, clinical assessments often draw upon a broader context, including historical notes, laboratory results, and other longitudinal patient data.
+Expanding our approach to integrate these diverse temporal data sources could facilitate a more holistic understanding of disease trajectories and patient history, significantly enhancing clinical applicability.
+Sparse Temporal Data Challenges
+In cases where prior scans are unavailable or minimally informative (e.g., taken within a short interval), our “dummy prior image” provides a workaround.
+However, the model’s ability to interpret and generate meaningful outputs under these constraints may still be limited.
+Future research could focus on synthesising or imputing temporal context to enhance performance under these constraints.
+Computational Complexity
+The use of temporal alignment mechanisms and multi-layer feature integration increases computational demands, posing challenges for deployment in resource-constrained environments.
+Future optimisation efforts should focus on reducing computational overhead while maintaining performance.
+Generalisability Across Modalities and Datasets
+Our study is limited to frontal-view chest X-rays and the MIMIC-CXR dataset (Johnson et al., 2019b).
+The applicability of our approach to other imaging modalities (e.g., CT, MRI) and datasets (e.g., CheXpert (Irvin et al., 2019), PadChest (Bustos et al., 2020)) remains unexplored.
+Future studies should assess the model’s generalisability to a broader range of datasets and imaging contexts.
+Based on the identified limitations, we outline the following directions:
+Develop frameworks for integrating multiple prior scans with dynamic temporal reasoning to better capture longitudinal changes.
+Expand the model to incorporate multi-modal imaging and textual data for more comprehensive diagnostic insights.
+Investigate the integration of diverse temporal data sources, such as electronic health records (EHRs), to enhance clinical applicability.
+Exploring lightweight model architectures for faster inference while maintaining high performance.
+These advancements aim to address the current limitations while broadening the applicability of temporal-aware multimodal models in radiology and other clinical domains.